### v2

# Understanding Ethereum Mempool Security under Asymmetric DoS by Symbolic Fuzzing

Abstract
In blockchains, mempool controls transaction flow before consensus, denial of whose service hurts the health and security of blockchain networks. This paper presents mpfuzz, the first mempool fuzzer to find asymmetric DoS bugs by symbolically exploring mempool state space and optimistically estimating the promisingness an intermediate state is in reaching bug oracles. Compared to the baseline blockchain fuzzers, mpfuzz achieves a >100×>100$\times$> 100 × speedup in finding known DETER exploits. Running mpfuzz on six major Ethereum clients leads to the discovering of new mempool vulnerabilities, which exhibit a wide variety of sophisticated patterns including stealthy mempool eviction and mempool locking. Rule-based mitigation schemes are proposed against newly discovered vulnerabilities.
1 Introduction
In Ethereum, a mempool buffers unconfirmed transactions from web3 users before they are included in the next blocks. Mempool provides the essential functionality to bridge the gap between varying rates of submitted transactions and rates of produced blocks, despite what types of transactions it serves, whether public or private transactions. As shown in recent studies [22], denying a mempool service can force the blockchain to produce blocks of low or even zero (Gas) utilization, undermining validators’ incentives and shrinking the blockchain networks in the long run, re-introducing the $51\%$ attacks. Besides, a denied mempool service can drop normal transactions, cutting millions of web3 users off the blockchain and failing the decentralized applications relying on real-time blockchain access.
Problem: Spamming the mempool to deny its service has been studied for long [17, 9, 27, 12]. Early designs by sending spam transactions at high fees burden attackers with high costs and are of limited practicality. What poses a real threat is Asymmetric DeniAl of Mempool Service, coined by ADAMS, in which the mempool service is denied at an asymmetrically low cost. That is, the attack costs in terms of the fees of adversarial transactions are significantly lower than those of normal transactions victimized by the denied mempool. In the existing literature, DETER [22] is the first ADAMS attack which works by sending invalid transactions to directly evict normal transactions in the mempool. MemPurge [29] is a similar mempool attack that finds a way to send overdraft transactions into Geth’s pending-transaction pool and cause eviction there. These known attacks are easy to detect (i.e., following the same direct-eviction pattern), and in fact, the DETER bugs reported in 2021 have been successfully fixed in all major Ethereum clients as of Fall 2023, including Geth, Nethermind, Erigon, and Besu. Given this state of affairs, we pose the following research question: Are there new ADAMS vulnerabilities in the latest Ethereum clients already patched against direct-eviction based attacks?
This work takes an automatic and systematic approach to discovering ADAMS vulnerabilities, unlike the existing DETER bugs that are manually found. Fuzzing mempool implementations is promising but also poses unique challenges: Unlike the consensus implementation that reads only valid confirmed transactions, the mempool residing pre-consensus needs to handle various unconfirmed transactions, imposing a much larger input space for the fuzzer. For instance, a mempool can receive invalid transactions under legitimate causes, [^1] For instance, future transactions can be caused by out-of-order information propagation in Ethereum. and factors such as fees are key in determining transaction admission outcomes. Existing blockchain fuzzers including Fluffy [30], Loki [23] and Tyr [19] all focus on fuzzing consensus implementation and don’t explore the extra transaction space required by mempool fuzzing. As a result, the naive application of a consensus fuzzer to fuzz mempool misses detecting the DETER bugs as evaluated in § 4.1, let alone discover more sophisticated new ADAMS bugs.
Proposed methods: To enable efficient mempool fuzzing, our key observation is that real-world mempool implementations admit transactions based on abstract “symbols”, such as pending and future transactions, instead of concrete value. Sending multiple concrete transactions generated under the same symbol triggers the same mempool behavior and should be avoided. Mutating one transaction under one symbol during fuzzing suffices to cover mempool behavior.
We propose symbolic stateful mempool fuzzing or mpfuzz. We first extract seven symbols from Ethereum semantics to cover future, parent, child, overdraft, latent overdraft, and replacement transactions. Then, mpfuzz generates a new transaction by symbol, sends the input to execute on a target mempool implementation, observes the end state of the mempool, and uses the symbolic state coverage as feedback to guide the next round of mutation. By this means, mpfuzz searches the symbolic mempool state space for the one matching the bug oracles induced from ADAMS. To further speed up the search in a limited time, we prioritize “promising” states by optimistically estimating the costs of unconfirmed transactions whose validity is subject to change in the future.
Found attacks: We run mpfuzz on six Ethereum execution-layer clients including Geth [2], Nethermind [5], Erigon [1], Besu [3], Reth [7], and OpenEthereum [6]. Compared to the baseline fuzzers, mpfuzz can find exploits faster by more than two orders of magnitude. On Ethereum clients of historical versions, mpfuzz can rediscover known DETER bugs. On the latest Ethereum clients where DETER bugs are fixed, mpfuzz can find new ADAMS attacks, described next.
We summarize the newly discovered ADAMS exploits into several patterns: 1) Indirect eviction by valid-turned-invalid transactions. Unlike the direct-eviction pattern used in DETER, the indirect-eviction attack works more stealthily in two steps: The attacker first sends normal-looking transactions to evict victim transactions from the mempool and then sends another set of transactions to turn the admitted normal-looking transactions into invalid ones, bringing down the cost. 2) Locking mempool, which is to occupy the mempool to decline subsequent victim transactions. Mempool locking does not need to evict existing transactions from the mempool as the eviction-based DETER attacks do. 3) Adaptive attack strategies where the attack composes adversarial transactions in an adaptive way to the specific policies and implementations of the mempool under test. Example strategies include composing transactions of multiple patterns in one attack, re-sending evicted transactions to evict “reversible” mempools, locking mempool patched against turning, etc.
We propose rule-based mitigation schemes against some of our discovered bugs and identify design challenges in mitigating others, which is left for future works.
All newly found ADAMS bugs are reported to the Ethereum Foundation with $7$ bugs confirmed and $3$ fixed in recent client releases [4]. Bug reporting is documented on the webpage [13] which also includes demos of mpfuzz on six Ethereum clients. mpfuzz will be open-sourced to facilitate vulnerability discovery in more and future clients.
Contributions: The paper makes contributions as follows:
$\bullet$ New fuzzing problem: This paper is the first to formulate the mempool-fuzzing problem to automatically find asymmetric DoS vulnerabilities as bug oracles. Fuzzing mempools poses new unique challenges that existing blockchain fuzzers don’t address and entails a larger search space including invalid transactions and varying fees.
$\bullet$ Symbolic fuzzing tool: The paper presents the design and implementation of mpfuzz, a symbolic stateful mempool fuzzer. Given a mempool implementation, mpfuzz searches the mempool state space using the feedback of symbolic state coverage and promising-ness to reach bug oracles. With a Python prototype and runs on real Ethereum clients, mpfuzz achieves a >100×>100$\times$> 100 × speedup in finding known DETER exploits compared to baselines.
$\bullet$ New mempool vulnerabilities: mpfuzz discovers new asymmetric-DoS vulnerabilities in six major Ethereum clients in the mainnet. By evaluation under real transaction workloads, all found attacks achieve $84.6-99.6\%$ success rates and low costs such as adversarial transaction fees 100×100$\times$100 × lower than the victim transaction fees.
2 Related Works
Consensus fuzzers: Fluffy [30] is a code-coverage based differential fuzzer to find consensus bugs in Ethereum Virtual Machines (EVM). Specifically, given an EVM input, that is, a transaction sequence, Fluffy sends it to multiple nodes running different Ethereum clients (e.g., Geth and OpenEthereum) for execution. 1) The test oracle in Fluffy is whether the EVM end states across these nodes are different, implying consensus failure. 2) Fluffy mutates ordered transaction sequences at two levels: it reorders/adds/deletes transactions in the sequence, and for each new transaction to generate, it randomly selects values in Gas limits and Ether amounts. For the data field, it employs semantic-aware strategies to add/delete/mutate bytecode instructions of pre-fixed smart contract templates. 3) Fluffy uses code coverage as feedback to guide the mutation.
Loki [23] is a stateful fuzzer to find consensus bugs causing crashes in blockchain network stacks. Specifically, Loki runs as a fuzzer node interacting with a tested blockchain node through sending and receiving network messages. 1) The test oracle in Loki is whether the tested blockchain node crashes. 2) The fuzzer node uses a learned model of blockchain network protocol to generate the next message of complying format, in which the content is mutated by randomly choosing integers and bit-flipping strings. In other words, the mutation in Loki is unaware of application level semantics. For instance, if the message is about propagating an Ethereum transaction, the transaction’s nonce, Ether amount, and Gas price are all randomly chosen; the sender, receiver, and data are bit flipped. 3) Loki records messages sent and received from the test node. It uses them as the state. New states receive positive feedback.
Tyr [19] is a property-based stateful fuzzer that finds consensus bugs causing the violation of safety, liveness, integrity, and other properties in blockchain network stacks. Specifically, in Tyr, a fuzzer node is connected to and executes a consensus protocol with multiple neighbor blockchain nodes. 1) The test oracle Tyr uses is the violation of consensus properties, including safety (e.g., invalid transactions cannot be confirmed), liveness (e.g., all valid transactions will be confirmed), integrity, etc. The checking is realized by matching an observed end state with a “correct” state defined and run by the Tyr fuzzer. 2) Tyr mutates transactions on generic attributes like randomly selecting senders and varying the amount of cryptocurrencies transferred (specifically, with two values, the full sender balance, and balance plus one). When applied to Ethereum, Tyr does not mutate Ethereum-specific attributes, including nonce or Gas price. It generates one type of invalid transaction, that is, the double-spending one, which in account-based blockchains like Ethereum are replacement transactions of the same nonce. 3) Tyr uses state divergence as the feedback, that is, the difference of end states on neighbor nodes receiving consensus messages in the same iteration. In Tyr, the states on a node include both the messages the node sent and received, as well as blocks, confirmed transactions, and other state information.
These existing blockchain fuzzers [19, 23, 30] cannot detect mempool DoS bugs and are different from this work. Briefly, mempool DoS does not trigger system crashes and can not be detected by Loki. Mempool content difference across nodes does not mean insecurity, and the mempool DoS bugs can not be detected by Fluffy (which detects consensus state difference). While Tyr aims to detect liveness and safety violations, their model of invalid transactions is rudimentary. Specifically, they only model double-spending transactions and cannot detect DETER bugs [22] that rely on more advanced invalid transactions, like future transactions and latent overdrafts.
Blockchain DoS: Blockchain DoS security has been examined at different system layers, including P2P networks [20, 24, 15, 28], mining-based consensus [25, 8], and application-level smart contracts [26, 18, 11] and DApp (decentralized application) services [21]. These DoSes are not related to mempools and are orthogonal to this work.
For mempool DoS, a basic attack is by sending spam transactions with high fees to evict victim transactions of normal fees [17, 9, 27, 12], which incur high costs to attackers. DETER [22] is the first asymmetric DoS on Ethereum mempool where the adversary sends invalid transactions of high fees (e.g., future transactions or latent overdrafts) to evict victim transactions of normal fees. Concurrent to this work is MemPurge [29] posted online in June 2023; in it, the attacker reconnects her adversarial future transactions and makes them latent-overdraft transactions. As we tested on the latest Ethereum clients in July 2023, such as Geth $v1.11.4$, the DETER bugs have been fixed. Both DETER and MemPurge are manually discovered.
3 Threat Model
In the threat model, an attacker aims at disabling the mempools on one or multiple nodes and causing the disruption of the transaction propagation service. This can lead to damage, such as delayed or dropped transaction inclusion and low-utilization blocks (even empty ones). The damage can benefit certain attackers in practice, such as competing RPC services, MEV bots, or block validators. Besides, the attacker aims at causing the above damage at asymmetrically low costs, that is, the fees of adversarial transactions included in the blockchain are lower than the fees of victim transactions outside the mempool, as will be presented.
The attacker has the same capability as in the DETER work [22], where the attacker joins an Ethereum network, discovers critical nodes (e.g., backend nodes of a RPC service, a node hosting MEV bot, or target block validators), and sends adversarial transactions.
3.1 Notations
Transactions: In Ethereum, we consider two types of accounts: a benign account of index $i$ denoted by $B_{i}$, and an adversarial account of index $i$ denoted by $A_{i}$. A (concrete) Ethereum transaction $tx$ is denoted by $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{s}&{v}\\ {n}&{f}\end{bmatrix}$}}$, where the sender is $s$, the nonce is $n$, the Gas price is $f$, and the transferred value in Ether is $v$. This work does not consider the “data” field in Ethereum transactions or smart contracts because smart contracts do not affect the validity of the Ethereum transaction. The notations are summarized in Table 1.
Table 1: Notations
Annotation
Meaning
Symbol
Meaning
$m$
Mempool length
$ops$
Tx sequence
$st$
State of residential txs in mempool
$dc$
History of declined txs from mempool
$A$
Adversarial tx sender
$B$
Benign tx sender
$tx\begin{bmatrix}{s}&{v}\\ {n}&{f}\end{bmatrix}$
A transaction from sender $s$, of nonce $n$, transferring Ether value $v$, and with Gas Price $f$
We use each of the following symbols to represent a disjoint group of transactions. Informally, $\mathcal{N}$ is a transaction sent from a benign account, $\mathcal{F}$ a future transaction, $\mathcal{P}$ a parent transaction, $\mathcal{C}$ a child transaction, $\mathcal{O}$ an overdraft transaction, $\mathcal{L}$ a latent overdraft transaction, and $\mathcal{R}$ a replacement transaction. These symbols are formally specified in § 4.
Particularly, we use the notion of latent overdraft from the existing work [22], which indicates a child transaction that by itself does not overdraft but does overdraft when taking into account its parent transactions. Suppose Alice has a balance of $5$ Ether and sends $tx_{1}$ of nonce $1$ spending $2$ Ether and $tx_{2}$ of nonce $2$ spending $4$ Ether. $tx_{2}$ is a latent overdraft.
States: We recognize two mempool-related states in Ethereum, the collection of transactions stored in mempool $st$, and the collection of declined transactions $dc$.
Suppose a mempool of state $\langle{}st_{i},dc_{i}\rangle{}$ receives an arriving transaction $tx_{i}$ and transitions to state $\langle{}st_{i+1},dc_{i+1}\rangle{}$.
An arriving transaction can be admitted with eviction, admitted without eviction, or declined by a mempool. The three operations are defined as follows. 1) An arriving transaction $tx_{i}$ is admitted with evicting a transaction $tx_{i}\textquoteright$ ’ from the mempool, if $st_{i+1}=st_{i}\setminus{}tx_{i}^{\prime}\cup{}tx_{i},dc_{i+1}=dc_{i}$. 2) $tx_{i}$ is admitted without evicting any transaction in the mempool, if $st_{i+1}=st_{i}\cup{}tx_{i},dc_{i+1}=dc_{i}$. 3) $tx_{i}$ is declined and does not enter the mempool, if $st_{i+1}=st_{i},dc_{i+1}=dc_{i}\cup{}tx_{i}$.
3.2 Bug Oracles
Suppose a mempool of initial state $\langle{}st_{0},dc_{0}\rangle{}$ receives a sequence of transactions, denoted by $ops$. Here, we simplify the problem by not considering arriving blocks during fuzzing. We define two bug oracles: eviction-based attacks and declining-based locking attacks.
Definition 3.1 (Eviction bug oracle)
A mempool execution from initial state $st_{0}$ to end state $st_{n}$ under transaction sequence $ops$ is a successful ADAMS eviction attack, if.f. 1) the attack damage is full, that is, end state $st_{n}$ contains only adversarial transactions and initial state $st_{0}$ contains only benign transactions, and 2) the attack cost is low, that is, the total adversarial transaction fees in the end states are smaller, by a multiplicative factor $\epsilon$, than the attack damage measured by the victim transaction fees in the initial state. Formally,
$st_{0}\cap{}(st_{n})$ $=$ $\varnothing$
$\sum_{tx\in{}st_{n}}tx.fee$ $<$ $\epsilon\sum_{tx\in{}st_{0}}tx.fee$ (1)
Definition 3.2 (Locking bug oracle)
A mempool execution from initial state $\langle{}st_{0},dc_{0}\rangle{}$ to end state $\langle{}st_{n},dc_{n}\rangle{}$ under transaction sequence $ops$ is a successful ADAMS locking attack if.f. 1) the attack damage is full, that is, the victim normal transactions in $ops$ are completely declined by the mempool storing adversarial transactions, where normal transactions and adversarial transactions are sent by different accounts, and 2) the attack cost is low, that is, the total adversarial transaction fees in the mempool are smaller, by a multiplicative factor $\lambda$, than the attack damage measured by the fees of victim transactions declined. Formally,
$\cup_{txd\in{}dc_{n}}{txd.sender}$$\cap$ $\cup_{tx\in{}s_{n}\cup{}ops\setminus{}dc_{n}}{tx.sender}=\varnothing$
$\sum_{tx\in{}st_{n}}tx.fee$ $<$ $\lambda\sum_{tx\in{}dc_{n}}tx.fee$ (2)
4 Stateful Mempool Fuzzing by mpfuzz
4.1 Motivating Observations
(In)feasibility of code-coverage-only stateless fuzzing: To fuzz a mempool, a baseline design is to use a code-coverage-based fuzzer. In each iteration, the fuzzer generates a bit string, parses it into a sequence of transactions and sends them to the tested mempool for execution. Upon each end state, it checks whether the test oracle defined above is satisfied. This design takes code coverage as the only feedback and is stateless because the feedback does not consider the end state.
The stateless design would be ineffective in fuzzing mempool or finding ADAMS exploits. For example, consider fuzzing a three-slot mempool to find a DETER-X exploit consisting of three future transactions [22]. Suppose the current mempool input is a sequence of one future transaction. After mutation, it may try an input of two future transactions, which, however, does not increase code coverage (the second future transaction is executed by the same code path in a mempool as the first future transaction). Thus, the fuzzer discards the two-future-transaction input and misses finding the three-future-transaction exploit. We validate the ineffectiveness of stateless fuzzer by experiments in § 6.
A more promising design is stateful mempool fuzzing, in which the mempool state is included in the feedback to guide next-iteration fuzzing. In the previous example of finding a three-future-transaction exploit (DETER-X), sending two future transactions leads to a different mempool state than sending one. A simple state-coverage-based fuzzer would view this as positive feedback and further explore this direction toward finding the three-future-transaction exploit.
(In)feasibility of existing consensus fuzzers: Existing consensus fuzzers [19, 23, 30] cannot detect mempool DoS or ADAMS bugs. Specifically, Loki’s test oracle is detecting system crashes, while a successful ADAMS attack does not necessarily trigger system crashes. Using Loki to detect ADAMS would cause many false negatives.
Fluffy detects the unsynchronizable difference of post-consensus states as vulnerability. However, the mempools in two clients, say Geth and Besu under benign transactions, would be different. In other words, post-consensus state differences across clients could indicate insecurity, but the difference in pre-consensus states across clients, like mempool, does not mean insecurity. Using Fluffly to detect ADAMS could cause false positives.
Tyr’s test oracle detects property violations on post-consensus states. Particularly, while Tyr models state liveness, their definition “valid transactions should be executed eventually” is, unfortunately, over-simplifying and fails to model the legitimate pre-consensus cases in real Ethereum clients; for instance, valid transactions of low Gas price can be dropped by Ethereum mempools and will not be eventually executed. Besides, they model double-spending transactions as the only type of invalid transaction. In Ethereum, double-spending transactions are of the same nonce. And there are many more sophisticated invalid transactions in Ethereum that Tyr does not model, including future transactions, latent overdrafts, etc.
We analyze the case of existing consensus fuzzers in detecting known DETER attacks [22]. Loki cannot detect DETER because DETER does not cause system crashes. Fluffy does not model future transactions and can not find the DETER attacks that rely on future transactions. Similarly, Tyr that does not model sophisticated invalid transactions would miss detecting DETER attacks.
Challenges: In stateful mempool fuzzing, the essential design problem is how to generate the next input transaction given a current “state”. State $st$ is the collection of transactions in the mempool and those included in the blockchain.
The key challenge is the large input space of transactions. Recall that an Ethereum transaction consists of multiple attributes (sender, nonce, Gas price, value, etc.), each defined in a large domain (e.g., 64-bit string). Exploring the raw transaction space is hard; the comprehensive search is inefficient, and randomly trying transactions as done in the state-of-the-art blockchain fuzzers (in § 2) is ineffective.
4.2 Proposed Fuzzer Design
4.2.1 Transaction Symbolization
Intuition: We propose symbolizing transactions for efficient and effective mempool stateful fuzzing. Our key idea is to map each group of concrete transactions triggering an equivalent mempool behavior into a distinct symbol so that searching one transaction is sufficient to cover all other transactions under the same symbol. Transaction symbolization is expected to reduce the search space from possible concrete transactions to the symbol space.
Table 2: Symbols, transactions and associated costs. Tx refers to the instantiated transaction under a given symbol.
Symbol Description
Tx
$cost()$ $opcost()$
$\mathcal{N}$ Benign
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B*}&{1}\\ {*}&{3}\end{bmatrix}$}}$
$3$ $3$
$\mathcal{F}$ Future
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+1}\end{bmatrix}$}}$
00 00
$\mathcal{P}$ Parent
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A>r}&{1}\\ {1}&{[4,m]}\end{bmatrix}$}}$
$[4,m]$ $[4,m]$
$\mathcal{C}$ Child
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$
$m+1$ $1$
$\mathcal{O}$ Overdraft
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m+1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$
00 00
$\mathcal{L}$ Latent overdraft
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$
00 00
$\mathcal{R}$ Replacement
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {1}&{m+1}\end{bmatrix}$}}$
00 00
Specification: Suppose the current state contains adversarial transactions sent from $r$ accounts $A_{1},\dots,A_{r}$. These accounts have an initial balance of $m$ Ether, where $m$ equals the mempool capacity (say $m=1000$).
Symbol $\mathcal{N}$ defines a transaction subspace covering any normal transactions sent from any benign account and of any nonce, namely $tx\begin{bmatrix}{B*}&{*}\\ {*}&{*}\end{bmatrix}$ In mpfuzz, Symbol $\mathcal{N}$ is instantiated to concrete transactions of fixed Gas price $3$ wei and of value $1$ Ether. That is, symbol $\mathcal{N}$ is instantiated to transaction $tx\begin{bmatrix}{B*}&{1}\\ {*}&{3}\end{bmatrix}$ , as shown in Table 2. Among all symbols, $\mathcal{N}$ is the only symbol associated with benign sender accounts.
Symbol $\mathcal{F}$ defines the transaction subspace of any future transaction sent from an adversarial account; the future transaction is defined w.r.t. the current state. mpfuzz instantiates symbol $\mathcal{F}$ to concrete transactions of nonce $m+1$, value $1$ Ether, Gas price $m+1$ wei, and any adversarial account $*A*$. The instantiating pattern for symbol $\mathcal{F}$ is $tx\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+1}\end{bmatrix}$.
Symbol $\mathcal{P}$ defines the transaction subspace of any parent transaction from an adversarial account w.r.t. the current state. mpfuzz instantiates $\mathcal{P}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A\geq{r}}&{1}\\ {1}&{[4,m]}\end{bmatrix}$}}$. Here, the transaction is fixed with a Gas price in the range $[4,m]$ wei, value at $1$ Ether, and nonce at $1$.
Symbol $\mathcal{C}$ defines the transaction subspace of any adversarial child transaction w.r.t. the state. mpfuzz instantiates $\mathcal{C}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$. The instantiated child transaction is fixed at Gas price of $m+1$
Symbol $\mathcal{O}$ defines the transaction subspace of any adversarial overdraft transaction w.r.t. state $st$. mpfuzz instantiates $\mathcal{O}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m+1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$. Recall that all adversarial accounts have an initial balance of $m$ Ether.
Symbol $\mathcal{L}$ defines the transaction subspace of any adversarial latent-overdraft transaction w.r.t. state $st$. mpfuzz instantiates $\mathcal{L}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$. All adversarial accounts have an initial balance of $m$ Ether.
Symbol $\mathcal{R}$ defines the transaction subspace of any adversarial “replacement” transaction w.r.t., the state. Unlike the symbols above, the transactions under Symbol $\mathcal{R}$ must share the same sender and nonce with an existing transaction in the current state $st$. mpfuzz instantiates $\mathcal{R}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {1}&{m+1}\end{bmatrix}$}}$. Here, we simplify the problem and only consider replacing the transaction with nonce $1$.
Figure 1: Symbols and transaction space reduction.
Given that a concrete transaction in this work consists of four attributes, transactions are defined in a four-dimensional space. We visualize transaction symbols (with an incomplete view) in two two-dimensional spaces in Figure 1, that is, one of transaction sender and nonce and the other of sender and value. For each symbol, the figures depict the transaction subspace covered by the symbol (in white shapes of black lines) and the transaction pattern instantiated by mpfuzz under the symbol (in red lines or shapes). In our design, the instantiated transactions are a much smaller subset of the defining transaction space. For instance, given state $st$ of transactions sent from account $A_{1},\dots,A_{r}$, transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ is within the defining space of symbol $\mathcal{F}$, but mpfuzz does not instantiate $\mathcal{F}$ by transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ (but instead to transactions $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+1}\end{bmatrix}$}}$).
4.2.2 State Search Algorithm
Algorithm 1 mpfuzz(SeedCorpus $sdb$, Mempool $mp$)
1:$sdb$.init($mp$);
2:while !$sdb$.is_empty() or timeout do
3:    $st,ops$ = $sdb$.next(); $\triangleright$ Selection by energy
4:    for all $ops^{\prime},st^{\prime}$ = mutateExec($ops,st$do
5:         if is_ADAMS($st^{\prime}$then $\triangleright$ Test oracle
6:             emit($ops^{\prime}$, $st^{\prime}$, "Found an exploit");
7:         else
8:             if !$sdb$.feedback($st^{\prime}$, $sdb$then
9:                 $sdb$.add($st^{\prime}$, $ops^{\prime}$);
10:             end if
11:         end if
12:    end for
13:end while
Algorithm overview: We propose a stateful fuzzing algorithm, listed in Algorithm 1, to selectively explore the input space in a way that prioritizes new and promising mempool states towards triggering ADAMS oracle.
The core data structure is a seed corpus or $sdb$ which stores a list of input-state pairs or seeds. Upon running the algorithm, the corpus maintains the mempool states covered so far by the algorithm execution. The algorithm runs an outer loop that continues until the corpus is empty, or only the states of zero energy are left, or timeout. In each iteration, the algorithm retrieves the next seed from the corpus based on how promising the seed is in reaching a state triggering the oracle (a.k.a., the energy [16]). A seed consists of a symbolic input $ops$, which consists of transaction symbols, and a symbolic state $st$ reached by running a transaction sequence instantiated from the input $ops$ (Line 3). The algorithm runs an inner loop, in each iteration of which it mutates the input in the current seed ($\langle{}st,in\rangle{}$) and executes the mutated input against a reinitialized empty mempool, producing end state $st^{\prime}$ (Line 4). If the end state satisfies ADAMS conditions (i.e., the test oracle), it emits the mutated input-state pair $ops^{\prime},st^{\prime}$ as a newly found exploit. The algorithm further checks the feedback: The feedback is positive if the mutated input $ops^{\prime}$ brings the mempool state $st^{\prime}$ closer to triggering the test oracle than $st$. This entails mutated state $st^{\prime}$ to be different from state $st$ (i.e., increased state coverage) and state $st^{\prime}$ to achieve larger damage or lower attack cost. In case of positive feedback, the algorithm would add mutated input-state pair $ops^{\prime},st^{\prime}$ to the corpus.
This algorithm assumes the mempool is deterministic. That is, given a seed $\langle{}ops,st\rangle{}$, running the same input $ops$ against an empty mempool multiple times always results in the same end state $st$. In practice, we generate inputs to avoid the non-deterministic behavior of real mempool implementations.
The algorithm can be configured with initial seeds and input-mutation strategies. By default, we use one initial seed whose input fills up the mempool with normal transactions. The default input-mutation strategy is to append the current input with a newly generated transaction.
Algorithm 2 mutateExec(SymbolInput $ops$, SymbolState $st$)
1:$ops^{\prime}$=mutateSymbol($ops$);
2:$st_{c}$=executeUnappended($ops,ops^{\prime}$);
3:$st_{c}^{\prime}$=executeAppended($ops^{\prime},st_{c},st$);
4:return $st_{c}^{\prime}$
Input mutation: Given a symbolized input $ops$ and symbolized state $st$, the mutation algorithm is to explore each “slightly” different input $ops^{\prime}$ and its associated state $st^{\prime}$, such that the next stage can find the inputs producing positive feedback.
Internally, the algorithm proceeds at two levels, that is, symbols and concrete value. Specifically, as shown in Algorithm 2, 1) it appends the symbolized input $ops$ with a previously untried symbol, generating the “mutated” symbolized input $ops^{\prime}$ (Line 1 in Algorithm 2). In this step, the algorithm tries different symbols in the following order: $\mathcal{P},\mathcal{L},\mathcal{C}$. 2) It instantiates symbolized $ops^{\prime}$ to a transaction sequence and executes the sequence in the tested mempool to obtain the concrete end state (Line 2 in Algorithm 2). This step is necessary for instantiating and executing the mutation in the next step. Specifically, in this step, the algorithm instantiates symbolized input $ops$ to a concrete input $in_{c}$, which is a sequence of transactions. It then drives the transactions to a reinitialized mempool for execution. It returns the resulting concrete state $st_{c}$. 3) The algorithm instantiates the appended symbols under the context of the previous state $st$ and $st_{c}$. At last, it executes the appended transaction on mempool $st_{c}$ to obtain the concrete end state $st^{\prime}_{c}$ under the mutated input (Line 3 in Algorithm 2).
Mutation feedback: We describe how mempool states are symbolized before presenting state-based feedback. In mpfuzz, a symbolized state $st$ is a list of transaction symbols, ordered first partially as follows: $\mathcal{N}\preceq{}\mathcal{E}\preceq{}\mathcal{F}\preceq{}\{\mathcal{P},\mathcal{L},\mathcal{C}\}$. $\mathcal{E}$ refers to an empty slot. State slots of the same symbols are independently instantiated into transactions, except for $\mathcal{P},\mathcal{C},\mathcal{L}$. Child transactions (i.e., $\mathcal{C},\mathcal{L}$) are appended to their parent transaction of the same sender (i.e., $\mathcal{P}$). Across different senders, symbols are ordered by the parent’s GasPrice. For instance, suppose a mempool stores four concrete transactions, $tx\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$, $tx\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$, $tx\begin{bmatrix}{A_{2}}&{*}\\ {1}&{5}\end{bmatrix}$, $tx\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10001}\end{bmatrix}$ , they are mapped to four symbols, $\mathcal{P},\mathcal{F},\mathcal{P},\mathcal{C}$, and are further ordered in a symbolic state by $\mathcal{F}\mathcal{P}\mathcal{C}\mathcal{P}$.
In mpfuzz, the feedback of an input $ops^{\prime}$ is based on the symbolic end state $st^{\prime}$. Positive feedback on state $st^{\prime}$ is determined conjunctively by two metrics: 1) State coverage that indicates state $st^{\prime}$ is not covered in the corpus, and 2) state promising-ness that indicates how promising the current state is to reach a state satisfying bug oracles.
$feedback(st^{\prime},st,sdb)=1$ , if $\text{if }st\_coverage(st^{\prime},sdb)=1\land{}$ (3)
$st\_promising(st^{\prime},st)=1$
Specifically, state coverage is determined by straightforwardly comparing symbolic state $st^{\prime}$, as an ordered list of symbols, with all the symbolic states in the corpus. For instance, state $\mathcal{F}\mathcal{P}\mathcal{C}\mathcal{P}$ is different from $\mathcal{F}\mathcal{P}\mathcal{P}\mathcal{C}$. This also implies concrete transaction sequence $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{2}}&{*}\\ {1}&{5}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10001}\end{bmatrix}$}}$ is the same with $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{2}}&{*}\\ {1}&{6}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10}\end{bmatrix}$}}$, as they are both mapped to the same symbolic state.
How promising a state $st^{\prime}$ is (i.e., $st\_promising(st^{\prime})$) is determined as follows: A mutated state $st^{\prime}$ is more promising than an unmutated state $st$ if any one of the three conditions is met: 1) State $st^{\prime}$ stores fewer normal transactions (under symbol $\mathcal{N}$) than state $st$, implying more transactions evicted and higher damage (i.e., $evict\_normal(st^{\prime},st)=1$). 2) State $st^{\prime}$ declines (speculatively) more incoming normal transactions than state $st$, also implying higher damage (i.e., $decline\_normal(st^{\prime},st)=1$). 3) The total fees of adversarial transactions in state $st^{\prime}$ are lower than those in state $st$, implying lower attack costs. We consider two forms of costs: concrete cost and symbolic cost. The former, denoted by $cost(st^{\prime})$, is simply the total fees of adversarial transactions instantiated from a symbolic state $st^{\prime}$. The latter, denoted by $opcost(st^{\prime})$, optimistically estimates the cost of transactions in the current state $st^{\prime}$ that contributes to a future state triggering test oracle. We will describe the symbolic cost next. Formally, state promising-ness is calculated by Equation 4.
$st\_promising(st^{\prime},st)=$ $1,$ if $\text{if }evict\_normal(st^{\prime},st)=1\lor{}$ (4)
$decline\_normal(st^{\prime},st)=1\lor{}$
$cost(st^{\prime})<cost(st)\lor{}$
$opcost(st^{\prime})<opcost(st)$
Seed energy: Recall that in mpfuzz, the next seed is selected from the corpus based on energy. By intuition, the energy of a state is determined based on how promising the state is in triggering the test oracle. In addition to the state promising-ness used in deciding feedback, state energy incorporates fuzz runtime information, such as how many times the state has been selected.
Specifically, the energy of a seed $\langle{}ops,st\rangle{}$ is determined by Equation 5.
$Energy(in,st)$ $=$ $b/opcost(st)$ (5)
First, each seed in the corpus records how many times it has been mutated. The more symbols it has mutated in the past, the less energy the seed currently has and the lower priority it will be selected next. $b$ can be configured differently to traverse the state tree differently. In particular, breadth-first search (BFS) is by the following configuration: $b=1$ if at least one symbol has not been tried (for mutation) in the current seed. Otherwise, $b=0$.
Second, we use the symbolic cost to estimate how promising a state is and use it in seed energy.
Estimate state cost: When mpfuzz determines how promising a state is, it needs to look beyond the current state and into all possible descendant states. We propose a heuristic that optimistically estimates the descendant-state costs given a current state. The key intuition is the following: In Ethereum, the validity of a child transaction (under symbol $\mathcal{C}$) depends on its parent/ancestor transaction. Thus, even though a transaction in the current state is valid, the transaction can be “turned” into an invalid one in subsequent states. We thus attribute, optimistically, the cost of transaction $\mathcal{C}$ to value $1$, so that it is preferable to a parent transaction $\mathcal{P}$ in input mutation and seed selection, and turning $\mathcal{C}$ into an invalid one like $\mathcal{L}$ or $\mathcal{F}$ is also encouraged. The cost profiles used in mpfuzz are summarized in Table 2.
5 Found Exploits
Running mpfuzz on six leading Ethereum execution-layer clients, namely, Geth, Besu, Nethermind, Erigon, Reth, and OpenEthereum, leads to the discovery of $22$ bugs on all six Ethereum clients, including $7$ conforming to the known DETER attacks on clients of historical versions and $15$ new bugs on the clients of the latest versions. We describe two selected patterns of these bugs as follows.
5.1 Found Eviction Attacks (Selected)
Exploit $XT_{4}$: Indirect eviction by valid-turned-overdraft transactions: This eviction attack works in two steps: 1) The attacker first sends valid transactions at high fees to evict normal transactions initially stored in the mempool. These transactions are sent from $k$ accounts, each of which sends $l$ transactions. 2) She then sends $k$ transactions; each of the transactions is of nonce $1$, at a high fee, of high value $v_{1}$, and from the same $k$ accounts in Step 1). The transactions would replace the transaction of the same sender and nonce sent in Step 1). Once the replacement is finished, they turn their child transactions into latent overdraft. Specifically, if a sender’s balance is $bal$ and the Ether value of transaction of nonce 2 is $v_{2}$, then $v_{1}$ is carefully crafted to enable turned latent overdraft, that is, $v_{1}<bal$ and $v_{1}+v_{2}>bal$.
In practice, many clients are found vulnerable under this exploit, including Geth $<v1.11.4$, Erigon, Besu, Nethermind, and OpenEthereum.
Exploit $XT_{6}$: Compositional indirect eviction (by multi-round valid-turned-future transactions): The latest Geth $\geq{}v1.11.4$ is patched against direct eviction attacks. To evade the protection, this eviction attack uses indirect eviction.
The attack works in two steps: 1) It first control-evicts the mempool using valid-turned-future transactions to avoid the triggering condition of its limit of transactions per sender. 2) It then conducts another run of valid-turned-future transactions, this time with transactions sent from one account, namely $1\times{}l$. It can succeed since the limit is not triggered. The overall attack cost is the fee of one transaction; compared to the single run of valid-turned-future attack, $XT_{6}$’s cost is much lower.
5.2 Found Locking Attacks (Selected)
[…]
While a FIFQ mempool can not be vulnerable to eviction attacks, it can be easily locked by Exploit $XT_{8}$ as follows: Whenever the attacker observes an empty slot present in the mempool, she sends a pending transaction of a minimally necessary fee to occupy the slot. Any transactions that arrive after will encounter a full mempool and are declined.
6 Performance Evaluation of mpfuzz
Preliminary of GoFuzz: Upon each fuzzing iteration, GoFuzz [10] generates a bit-string of variable length, feeds it to our fuzzer code, and receives from our fuzzer a binary value of feedback, be either positive or negative. The GoFuzz further checks if this run increases the code coverage in the Geth client (as will be described, our fuzzer invokes functions in Geth’s txpool). If the feedback is positive and code coverage increases, GoFuzz adds the current bit-string to the seeds before running into the next iteration.
Baseline B1: Stateless fuzzer. We implement a stateless fuzzer in the GoFuzz framework described above. 1) Given a bit string generated by GoFuzz, our fuzzer code parses it into a sequence of transactions; the length of the sequence depends on the length of the bit string. It skips a bitstring that is too short. 2) The fuzzer then sets up and initializes a Geth txpool instance with $m$ normal transactions, each sent from a distinct account, of nonce $1$, of value $1$ Ether (lower than the account balance), and of Gas price $3$ wei. 3) The fuzzer sends the transaction sequence to the initialized mempool for execution. 4) It checks the mempool state after execution: If the eviction-based test oracle (Equation 1) is met, it emits the current transaction sequence as an exploit.
Baseline B2: Concrete-state-coverage fuzzer. We also build the second baseline, a mempool fuzzer that takes concrete state coverage as feedback. Specifically, in each iteration, the fuzzer appends to the current transaction sequence a new transaction. Given the $m$-slot mempool, the fuzzer tries $m$ values for senders, $m$ values for nonces, $m$ values for Gas price, and $m$ values for Ether amount. After sending to the mempool the current transaction sequence, it sorts the transactions in the mempool by senders and nonces. It then hashes the sorted transactions in the mempool and checks the presence of the hash digest in the seeds. If no hash exists, the current transaction sequence would increase the concrete-state coverage and is inserted to the corpus. The corpus is a FIFO queue where the seed inserted earlier has high priority to be de-queued.
Baseline B3: Invalid-tx energy fuzzer. We build the last baseline that is similar to the concrete-state-based fuzzer except for the following: It uses as the energy the number of invalid transactions on a mempool state. The seed with more invalid transactions in the mempool has higher energy or priority to be selected for the next round of fuzzing.
We implement Baseline B1 in Go/GoFuzz and implement B2 and B3 in Python. We use a test oracle that can detect DETER attacks, that is, the number of invalid transactions in a mempool state equal $m-1$.
We run the three baselines and mpfuzz against a Geth client under two settings: A small setting where the mempool is resized to $6$ slots $m=6$ and all fuzzers run for two hours (if test oracle is not triggered), and a medium setting where $m=16$ and fuzzers run for $16$ hours.
Table 3: Fuzzing Geth $v1.11.3$’s mempool by different approaches to detect ADAMS exploit.
Settings
B1 B2 B3 mpfuzz
6-slot-2h
timeout $54$ (min) $8$ (min)
$0.11$ (min)
16-slot-16h
timeout timeout $447$ (min)
$0.84$ (min)
Table 3 reports the results under the small and medium settings. For the small setting with $6$ slots, the stateless fuzzer (baseline B1) cannot find any DETER attack in two hours, concrete-state-coverage fuzzer (baseline B2) can find a DETER exploit in 54 minutes, and the stateful fuzzer based on invalid-transaction counter finds a DETER exploit in $8$ minutes. By contrast, mpfuzz finds a DETER exploit in $0.11$ minutes.
For the medium setting, with $6$ slots, both baseline B1 and B2 cannot find any DETER exploit in $16$ hours, and baseline B3 can find a DETER exploit in $7.4$ hours. By contrast, mpfuzz finds a DETER exploit in under a minute.
We further run our mpfuzz on different clients, including Geth $v1.10.11$, Geth $v1.11.4$, Erigon $v2.42.0$, and Nethermind $v1.18.0$. The tested mempool is configured at $16$ slots. We report the number of exploits found by mpfuzz in a $16$-hour period. Figure 3 presents the result that mpfuzz finds $4096$ exploits in the first $8$ hours on Geth-$v1.10.11$. On Geth-$v1.11.4$, mpfuzz quickly finds one exploit (i.e., exploit $XT_{6}$ as described in § 5.1) within $2$ minutes and does not find anymore in the next $16$ hours. On Nethermind, it finds one exploit (exploit $XT_{1}$) within two minutes and finds the next one (exploit $XT_{4}$) near the end of $16$-th hour. On Erigon, it finds one exploit ($XT_{4}$) in the $16$-th hour.
Figure 2: # exploits found
Figure 2: # exploits found
Figure 3: # states covered
Figure 3 shows the number of states that are explored by mpfuzz on a $16$-slot mempool in 16 hours. mpfuzz explores $749$ and $91428$ states on Geth $v1.10.11$ in $2$ minutes and $16$ hours respectively. On Geth $v1.11.4$, mpfuzz explores $372$ and $23286$ states in $2$ minutes and $16$ hours respectively. However, the performance of Erigon and Nethermind is much lower than that of Geth. mpfuzz explores $12$ and $1239$ states on Erigon in $2$ minutes and $16$ hours respectively. On Nethermind, mpfuzz explores $15$ and $1433$ states in $2$ minutes and $16$ hours, respectively.
The reason mpfuzz is more performant on Geth is that we implemented an external API on Geth to initialize the mempool; in each iteration of fuzzing on Geth, mpfuzz calls the API to initialize the mempool. In contrast, on Nethermind and Erigon, mpfuzz restarts the client in each iteration, which is consuming.
7 Evaluation of Discovered Attacks
This section shows the evaluation of turning attacks and locking attacks.
7.1 Evaluation of Turning Attacks
We design experiments to measure the success of ADAMS attacks in real Ethereum networks.
7.1.1 Evaluation in Testnet
Setup: We set up our experiment platform by connecting the attacker node to the Goerli testnet. The attacker node runs an instrumented Geth client at the execution layer which sends crafted transactions to the testnet.
To monitor transaction propagation, we launch an independent supernode (called measurement node) in the same testnet. The measurement node runs a reconfigured Geth client where the limit of peers/neighbors is removed and the measurement node can connect to as many neighbors as possible, hence a super node. The measurement node is not (directly) connected to the attacker node. When setting up our experiments, we ran the measurement node in Goerli for seven days and found the node is stabilized at $290$ neighbors.
(a) Transaction propagation in six messages
(c) Etherscan screenshot of the blocks generated during the attack on Goerli (txs in six messages)
Figure 4: Mounting $XT_{4}$ attacks on Goerli: Emptying blocks and propagating malicious transactions to the entire network.
Experiment results: The instrumented measurement node is able to log the received messages from different neighbors; these messages include those of transactions and of transaction hashes. The measurement node could receive the same transaction from different neighbors, and the log stores the transaction-neighbor pairs. We started logging the received messages one month before the experiment.
To do an experiment, we make the attacker node send $XT_{4}$ transactions using $384$ accounts. Each account sends $16$ valid pending transactions, followed by a replacement transaction. In total, the attacker node sends $6144$ valid transactions and $384$ replacement transactions. The Gas prices of the valid and replacement transactions are set to be $8$ Gwei and $130$ Gwei, respectively. In one setting, we split the $6528$ transactions into six messages (i.e., the first five being $1229$ transactions each and the last one with $384$ transactions). We send these messages in order and wait for $1.5$ seconds between two consecutive messages. For comparison, in another setting, we send the $6144$ valid transactions in one message and the $384$ replacement transactions in another message. Finally, we wrap up the experiment by waiting after the attacker node sends all messages and all replacement transactions are included in the blockchain. We then turn off the measurement node’s logging and search for the presence of all $6528$ transactions across neighbor nodes in the log. Given a transaction nonce, we report the percentage of neighbor nodes in the testnet that have propagated it to the measurement node.
Table 4: Types of transactions included in the four blocks during the attack in Figure 4c. ‘# private tx’ means the number of private transactions.
# block
# private txs
# public txs
# attack txs
$8040010$
$5$
00
$384$
$8040009$
$78$
$4$
$384$
$8040008$
$1$
00
$384$
$8040007$
$5$
00
$384$
We run experiments of six-message and two-message propagation, repeat each experiment by $3$ times, and report the average result of node percentage in Figure 4a and Figure 4b. Figure 4a shows the transaction propagation with six messages. More than 90% of the testnet nodes have received all of the $15$ valid transactions in a sequence and the replacement transactions. The only exception is the relatively low percentage for the $16$-th transaction of the maximal nonce. By comparison, when sending the transactions in two messages, the propagation fails as shown in Figure 4b where nodes did not propagate any transaction of nonce larger than $5$.
Additionally, Figure 4c shows the generated blocks in the experiment propagating six messages. We took the screenshot from etherscan.io and labeled it in red with information regarding the attack. Before the attack begins, the Goerli testnet normally utilizes $24-47\%$ of the Gas in a block. For ethical consideration, we limit our attack to be short, specifically, four blocks long.
Right after the launch of the attack, in block $8040007$, the Gas used by normal transactions drops to $3.19\%$ and the Gas used by adversarial transactions (denoted by red bars in Figure 4c) is $26.88\%$. The included adversarial transactions are $384$ replacement transactions sent in the second round of $XT_{4}$. The $6144-384=5760$ child transactions sent in the first round are not included in the block. Other blocks during the attack, namely $8040008$ and $8040010$, are similar. In the third block $8040009$, $73.1\%$ Gas is spent on including $82$ normal transactions as shown in Table 4 not sent by our attacker node. To explain it, we inspect the raw history of transactions arriving at and logged by our measurement supernode and found that out of the $82$ included normal transactions, only $4$ are present in the log, implying the other $78$ normal transactions are private ones that were not broadcast to the measurement node. Furthermore, the private mempool in today’s Ethereum clients [14] has a capability of no limit, which, while being feasible under today’s low private-transaction workloads, can cause resource-exhaustion vulnerability. Notice that we did not discover specific critical nodes to attack as needed by DETER attacks [22].
8 Discussions
[…]
Defenses: Rule-based mitigation can fix some of our discovered bugs: $XT_{4}$ is fixed by detecting latent overdraft transactions upon replacement. Locking-based $XT_{8}$ is fixed by re-enabling eviction and replacement; to avoid reintroducing eviction vulnerabilities, the fix needs to comply with the security rules against eviction attacks.
Ethical concerns: When evaluating attacks, we only mounted attacks on the testnet and did so with minimal impacts on the tested network. For instance, our attack lasts a short period of time, say no more than $4$ blocks produced. We did not test our attack on the Ethereum mainnet. When reporting bugs, we disclose to the developers the mitigation design tradeoff and the risk of fixing one attacks by enabling other attacks (described above). To prevent introducing new bugs, we did not suggest fixes against locking ($XT_{8}$).
Private and public mempools: In Ethereum, transactions can be sent privately or publicly. A private transaction received on a block builder is directly packed to a block without broadcast to other builders. By contrast, a public transaction received on a node is broadcast to other nodes before it is included in the next block.
On the one hand, whether a mempool stores private or public transactions is orthogonal to the admission/eviction policy the mempool runs. Empirically, we found three open-source Ethereum block builders, named EigenPhi builder, Flashbot builder $v1.11.5-0.2.1$ and bloXroute builder. Specifically, Flashbot and bloXrout builder run the same admission policy as Geth $v1.11.4$, while EigenPhi builder runs the same admission policy as Geth $<v1.11.4$.Therefore, these private mempools are subject to the same vulnerabilities as the public mempools.
On the other hand, a private mempool receives transactions from a portion of the web3 users, while a public mempool receives transactions from all web3 users. Thus, a disabled private mempool creates limited damage by victimizing the portion of users connected to the pool. The damage of a disable public mempool can be amplified by broadcasting the payload transactions across the network and reaching the public mempools of all nodes.
9 Conclusion
This paper presents mpfuzz, the first mempool fuzzer to find asymmetric DoS bugs by symbolically exploring mempool states and optimistically estimating the promisingness an intermediate state is in reaching bug oracles. Running mpfuzz on six major Ethereum clients leads to the discovery of new mempool vulnerabilities, which exhibit a wide variety of sophisticated patterns including stealthy mempool eviction and mempool locking.

### v3

@@
-Understanding Ethereum Mempool Security under Asymmetric DoS by Symbolic Fuzzing
+Understanding Ethereum Mempool Security under Asymmetric DoS by Symbolized Stateful Fuzzing
 Abstract
 In blockchains, mempool controls transaction flow before consensus, denial of whose service hurts the health and security of blockchain networks.
-This paper presents mpfuzz, the first mempool fuzzer to find asymmetric DoS bugs by symbolically exploring mempool state space and optimistically estimating the promisingness an intermediate state is in reaching bug oracles.
+This paper presents mpfuzz, the first mempool fuzzer to find asymmetric DoS bugs by exploring the space of symbolized mempool states and optimistically estimating the promisingness of an intermediate state in reaching bug oracles.
 Compared to the baseline blockchain fuzzers, mpfuzz achieves a >100×>100$\times$> 100 × speedup in finding known DETER exploits.
-Running mpfuzz on six major Ethereum clients leads to the discovering of new mempool vulnerabilities, which exhibit a wide variety of sophisticated patterns including stealthy mempool eviction and mempool locking.
-Rule-based mitigation schemes are proposed against newly discovered vulnerabilities.
+Running mpfuzz on major Ethereum clients leads to discovering new mempool vulnerabilities, which exhibit a wide variety of sophisticated patterns, including stealthy mempool eviction and mempool locking.
+Rule-based mitigation schemes are proposed against all newly discovered vulnerabilities.
 1 Introduction
 In Ethereum, a mempool buffers unconfirmed transactions from web3 users before they are included in the next blocks.
-Mempool provides the essential functionality to bridge the gap between varying rates of submitted transactions and rates of produced blocks, despite what types of transactions it serves, whether public or private transactions.
+Mempool provides the essential functionality to bridge the gap between varying rates of submitted transactions and rates of produced blocks, regardless of public or private transactions it serves.
 As shown in recent studies [28], denying a mempool service can force the blockchain to produce blocks of low or even zero (Gas) utilization, undermining validators’ incentives and shrinking the blockchain networks in the long run, re-introducing the $51\%$ attacks.
-Besides, a denied mempool service can drop normal transactions, cutting millions of web3 users off the blockchain and failing the decentralized applications relying on real-time blockchain access.
+Besides, a denied mempool service can prevent normal transactions from block inclusion, cutting millions of web3 users off the blockchain and failing the DApps relying on real-time blockchain access.
 Problem: Spamming the mempool to deny its service has been studied for long [23, 15, 33, 18].
-Early designs by sending spam transactions at high fees burden attackers with high costs and are of limited practicality.
+Early designs by sending spam transactions at high prices burden attackers with high costs and are of limited practicality.
 What poses a real threat is Asymmetric DeniAl of Mempool Service, coined by ADAMS, in which the mempool service is denied at an asymmetrically low cost.
-That is, the attack costs in terms of the fees of adversarial transactions are significantly lower than those of normal transactions victimized by the denied mempool.
-In the existing literature, DETER [22] is the first ADAMS attack which works by sending invalid transactions to directly evict normal transactions in the mempool.
-MemPurge [29] is a similar mempool attack that finds a way to send overdraft transactions into Geth’s pending-transaction pool and cause eviction there.
-These known attacks are easy to detect (i.e., following the same direct-eviction pattern), and in fact, the DETER bugs reported in 2021 have been successfully fixed in all major Ethereum clients as of Fall 2023, including Geth, Nethermind, Erigon, and Besu.
+That is, the attack costs, in terms of the fees of adversarial transactions, are significantly lower than those of normal transactions victimized by the denied mempool.
+In the existing literature, DETER [28] is the first ADAMS attack, and it works by sending invalid transactions to directly evict normal transactions in the mempool.
+MemPurge [35] is a similar mempool attack that finds a way to send overdraft transactions into Geth’s pending transaction pool and causes eviction there.
+These known attacks are easy to detect (i.e., following the same direct-eviction pattern).
+In fact, the DETER bugs reported in 2021 have been successfully fixed in all major Ethereum clients as of Fall 2023, including Geth, Nethermind, Erigon, and Besu.
 Given this state of affairs, we pose the following research question: Are there new ADAMS vulnerabilities in the latest Ethereum clients already patched against direct-eviction based attacks?
-This work takes an automatic and systematic approach to discovering ADAMS vulnerabilities, unlike the existing DETER bugs that are manually found.
-Fuzzing mempool implementations is promising but also poses unique challenges: Unlike the consensus implementation that reads only valid confirmed transactions, the mempool residing pre-consensus needs to handle various unconfirmed transactions, imposing a much larger input space for the fuzzer.
+This work takes a systematic and semi-automated approach to discovering ADAMS vulnerabilities, unlike the existing DETER bugs that are manually found.
+Fuzzing mempool implementations is a promising approach but also poses unique challenges: Unlike the consensus implementation that reads only valid confirmed transactions, the mempool, which resides in the pre-consensus phase, needs to handle various unconfirmed transactions, imposing a much larger input space for the fuzzer.
 For instance, a mempool can receive invalid transactions under legitimate causes, [^1] For instance, future transactions can be caused by out-of-order information propagation in Ethereum.
-and factors such as fees are key in determining transaction admission outcomes.
+and factors such as fees or prices are key in determining transaction admission outcomes.
 Existing blockchain fuzzers including Fluffy [36], Loki [29] and Tyr [25] all focus on fuzzing consensus implementation and don’t explore the extra transaction space required by mempool fuzzing.
-As a result, the naive application of a consensus fuzzer to fuzz mempool misses detecting the DETER bugs as evaluated in § 4.1, let alone discover more sophisticated new ADAMS bugs.
-Proposed methods: To enable efficient mempool fuzzing, our key observation is that real-world mempool implementations admit transactions based on abstract “symbols”, such as pending and future transactions, instead of concrete value.
-Sending multiple concrete transactions generated under the same symbol triggers the same mempool behavior and should be avoided.
-Mutating one transaction under one symbol during fuzzing suffices to cover mempool behavior.
-We propose symbolic stateful mempool fuzzing or mpfuzz.
-We first extract seven symbols from Ethereum semantics to cover future, parent, child, overdraft, latent overdraft, and replacement transactions.
-Then, mpfuzz generates a new transaction by symbol, sends the input to execute on a target mempool implementation, observes the end state of the mempool, and uses the symbolic state coverage as feedback to guide the next round of mutation.
-By this means, mpfuzz searches the symbolic mempool state space for the one matching the bug oracles induced from ADAMS.
-To further speed up the search in a limited time, we prioritize “promising” states by optimistically estimating the costs of unconfirmed transactions whose validity is subject to change in the future.
-Found attacks: We run mpfuzz on six Ethereum execution-layer clients including Geth [2], Nethermind [5], Erigon [1], Besu [3], Reth [7], and OpenEthereum [6].
+As a result, directly re-purposing a consensus fuzzer to fuzz mempool would be unable to detect the DETER bugs as evaluated in Appendix § A, let alone discover more sophisticated new ADAMS bugs.
+Proposed methods: To efficiently fuzz mempools, our key observation is that real-world mempool implementation admits transactions based on abstract “symbols”, such as pending and future transactions, instead of concrete value.
+Thus, sending multiple transactions under one symbol would trigger the same mempool behavior repeatedly, and it suffices to explore just one transaction per symbol during fuzzing without losing the diversity of mempool behavior.
+We propose symbolized-stateful mempool fuzzing or mpfuzz.
+To begin with, mpfuzz is set up and run in a three-step workflow: It is first manually set up against a size-reduced mempool under test, which induces a much smaller search space for fuzzing.
+Second, mpfuzz is iteratively run against the reduced mempool to discover short exploits.
+Third, the short exploits are manually extended to actual ones that are functional on real mempools and Ethereum clients.
+Internally, the design of mpfuzz is based on seven transaction symbols we design out of Ethereum semantics: future, parent, overdraft, latent overdraft, and replacement transactions.
+Under these symbols, in each iteration of fuzzing, mpfuzz explores one concrete transaction per symbol, sends the generated transaction sequence to the target mempool, observes the mempool end state, and extracts the feedback of symbolized state coverage and state promisingness in reaching bug oracles to guide the next round of fuzzing.
+In particular, mpfuzz employs a novel technique to evaluate state promisingness, that is, by optimistically estimating the costs of unconfirmed transactions whose validity is subject to change in the future.
+Found attacks: We run mpfuzz on six leading execution-layer Ethereum clients deployed on the mainnet’s public-transaction path (Geth [7], Nethermind [11], Erigon [5], Besu [10], Reth [13], and OpenEthereum [12]), three PBS builders (proposer-builder separation) on the mainnet’s private-transaction and bundle path (Flashbot $v1.11.5$ [6], EigenPhi [4] and bloXroute [1]), and three Ethereum-like clients (BSC $v1.3.8$ [2] deployed on Binance Smart Chain, go-opera $v1.1.3$ [9] on Fantom, and core-geth $v1.12.19$ [3] on Ethereum Classic).
 Compared to the baseline fuzzers, mpfuzz can find exploits faster by more than two orders of magnitude.
@@
 Example strategies include composing transactions of multiple patterns in one attack, re-sending evicted transactions to evict “reversible” mempools, locking mempool patched against turning, etc.
-We propose rule-based mitigation schemes against some of our discovered bugs and identify design challenges in mitigating others, which is left for future works.
-All newly found ADAMS bugs are reported to the Ethereum Foundation with $7$ bugs confirmed and $3$ fixed in recent client releases [4].
-Bug reporting is documented on the webpage [13] which also includes demos of mpfuzz on six Ethereum clients.
+We propose rule-based mitigation schemes against all newly discovered vulnerabilities.
+All newly found ADAMS bugs are reported to the developer communities, including Ethereum Foundation, BSC, Fantom, PBS builders, etc.
+with $15$ bugs confirmed and $4$ fixed in recent client releases [8].
+Bug reporting is documented on the webpage [20], which also includes demos of mpfuzz on the tested Ethereum clients.
 mpfuzz will be open-sourced to facilitate vulnerability discovery in more and future clients.
@@
 $\bullet$ New fuzzing problem: This paper is the first to formulate the mempool-fuzzing problem to automatically find asymmetric DoS vulnerabilities as bug oracles.
-Fuzzing mempools poses new unique challenges that existing blockchain fuzzers don’t address and entails a larger search space including invalid transactions and varying fees.
-$\bullet$ Symbolic fuzzing tool: The paper presents the design and implementation of mpfuzz, a symbolic stateful mempool fuzzer.
-Given a mempool implementation, mpfuzz searches the mempool state space using the feedback of symbolic state coverage and promising-ness to reach bug oracles.
+Fuzzing mempools poses new unique challenges that existing blockchain fuzzers don’t address and entails a larger search space including invalid transactions and varying prices.
+$\bullet$ New fuzzing method: The paper presents the design and implementation of mpfuzz, a symbolized-stateful mempool fuzzer.
+Given a mempool implementation, mpfuzz defines the search space by the mempool states covered under symbolized transaction sequences and efficiently searches this space using the feedback of symbolized state coverage and the promising-ness of an intermediate state in triggering bug oracles.
 With a Python prototype and runs on real Ethereum clients, mpfuzz achieves a >100×>100$\times$> 100 × speedup in finding known DETER exploits compared to baselines.
-$\bullet$ New mempool vulnerabilities: mpfuzz discovers new asymmetric-DoS vulnerabilities in six major Ethereum clients in the mainnet.
+$\bullet$ New discovery of mempool vulnerabilities: mpfuzz discovers new asymmetric-DoS vulnerabilities in six major Ethereum clients in the mainnet.
 By evaluation under real transaction workloads, all found attacks achieve $84.6-99.6\%$ success rates and low costs such as adversarial transaction fees 100×100$\times$100 × lower than the victim transaction fees.
@@
 2) The fuzzer node uses a learned model of blockchain network protocol to generate the next message of complying format, in which the content is mutated by randomly choosing integers and bit-flipping strings.
-In other words, the mutation in Loki is unaware of application level semantics.
+That is, the mutation in Loki is unaware of application level semantics.
 For instance, if the message is about propagating an Ethereum transaction, the transaction’s nonce, Ether amount, and GasPrice are all randomly chosen; the sender, receiver, and data are bit flipped.
@@
 Concurrent to this work is MemPurge [35] posted online in June 2023; in it, the attacker reconnects her adversarial future transactions and makes them latent-overdraft transactions.
-As we tested on the latest Ethereum clients in July 2023, such as Geth $v1.11.4$, the DETER bugs have been fixed.
+The DETER bugs have been fixed on the latest Ethereum clients such as Geth $v1.11.4$, as we tested them in July 2023.
 Both DETER and MemPurge are manually discovered.
-3 Threat Model
-In the threat model, an attacker aims at disabling the mempools on one or multiple nodes and causing the disruption of the transaction propagation service.
-This can lead to damage, such as delayed or dropped transaction inclusion and low-utilization blocks (even empty ones).
-The damage can benefit certain attackers in practice, such as competing RPC services, MEV bots, or block validators.
-Besides, the attacker aims at causing the above damage at asymmetrically low costs, that is, the fees of adversarial transactions included in the blockchain are lower than the fees of victim transactions outside the mempool, as will be presented.
-The attacker has the same capability as in the DETER work [22], where the attacker joins an Ethereum network, discovers critical nodes (e.g., backend nodes of a RPC service, a node hosting MEV bot, or target block validators), and sends adversarial transactions.
+3 Background
+Ethereum mempools: In Ethereum, users send their transactions to the blockchain network, which are propagated to reach validator nodes.
+On every blockchain node, unconfirmed transactions are buffered in the mempool before they are included in the next block or evicted by another transaction.
+In Ethereum 2.0, transactions are propagated in two fashions: public transactions are broadcast among all nodes, and private transactions are forwarded to selected nodes, more specifically, selected builders and proposers (as in PBS or proposer-builder separation).
+The mempool is present on both paths, and it is inside Ethereum clients such as Geth [7] and Nethermind [11] (handling public transactions), and Flashbot [6] (private transactions).
+On these clients, a mempool also serves many other downstream modules, including MEV searchers or bots (in PBS), Gas stations, RPC queries, etc.
+Mempools on different nodes are operated independently and don’t need to synchronize as the consensus layer does.
+For instance, future transactions are not propagated, and the set of future transactions in the mempool on one node is different from another node.
+The core mempool design is transaction admission: Given an initial state, whether and how the mempool admits an arriving transaction.
+In practice, example policies include those favoring admitting transactions of higher prices (and thus evicting or declining the ones of lower prices), transactions arriving earlier, certain transaction types (e.g., parent over child transactions), and valid transactions (over future transactions).
+See different types of Ethereum transactions in § 4.1.
+Transactions and fees: In Ethereum, Gas measures the amount of computations caused by smart contract execution.
+When preparing for transaction $tx$, the sender needs to specify Gas and GasPrice; the former is the maximal amount of computations allowed for running smart contracts under $tx$, and the latter is the amount of Ether per Gas the sender is willing to pay.
+After $tx$ is included in a block, the actual amount of computations consumed by contract execution is denoted by GasUsed.
+The fees of transaction $tx$ are the product of GasUsed and GasPrice, that is, $tx.fee=\textit{GasUsed}*\textit{GasPrice}$.
+After EIP-1559 [19], GasPrice is divided into two components: $BasePrice$ and $PriorityPrice$, that is, $\textit{GasPrice}=BasePrice+PriorityPrice$.
+Blockchain nodes follow the Ethereum protocol to derive $BasePrice$ from Gas utilization in recent blocks.
+The fees associated with the $BasePrice$ are burnt upon transaction inclusion.
+$PriorityPrice$ is set by the transaction sender.
+4 Threat Model and Bug Oracle
+In the threat model, an attacker controls one or few nodes to join an Ethereum network, discovers and neighbors critical nodes (e.g., top validators, backends of an RPC Service, MEV searcher) if necessary, and sends crafted adversarial transactions to their neighbors.
+The attacker in this work has the same networking capacity as in DETER [28].
+The attacker’s goal is to disable the mempools on the critical or all nodes in an Etheurem network and cause damage such as dropped transactions in block inclusion, produced blocks of low or zero (Gas) utilization, and disabled other downstream services.
+In practice, this damage is beneficial to competing block validators, MEV searchers, and RPC services.
+Besides, the attacker aims to cause this damage at asymmetrically low costs, as will be described.
 4.1 Notations
@@
 Table 1: Notations
-Annotation
-Meaning
-Symbol
-Meaning
-$m$
-Mempool length
-$ops$
-Tx sequence
-$st$
-State of residential txs in mempool
-$dc$
-History of declined txs from mempool
-$A$
-Adversarial tx sender
-$B$
-Benign tx sender
-$tx\begin{bmatrix}{s}&{v}\\ {n}&{f}\end{bmatrix}$
-A transaction from sender $s$, of nonce $n$, transferring Ether value $v$, and with Gas Price $f$
+Symbol Meaning Symbol Meaning
+$m$ Mempool length $ops$ Tx sequence
+$st$ State of residential txs in mempool $dc$ History of declined txs from mempool
+$A$ Adversarial tx sender $B$ Benign tx sender
+$tx\begin{bmatrix}{s}&{v}\\ {n}&{f}\end{bmatrix}$ A transaction from sender $s$, of nonce $n$, transferring Ether value $v$, and with GasPrice $f$
 We use each of the following symbols to represent a disjoint group of transactions.
@@
 $tx_{2}$ is a latent overdraft.
-States: We recognize two mempool-related states in Ethereum, the collection of transactions stored in mempool $st$, and the collection of declined transactions $dc$.
+States and transitions: We recognize two mempool-related states in Ethereum, the collection of transactions stored in mempool $st$, and the collection of declined transactions $dc$.
 Suppose a mempool of state $\langle{}st_{i},dc_{i}\rangle{}$ receives an arriving transaction $tx_{i}$ and transitions to state $\langle{}st_{i+1},dc_{i+1}\rangle{}$.
@@
 The three operations are defined as follows.
-1) An arriving transaction $tx_{i}$ is admitted with evicting a transaction $tx_{i}\textquoteright$ ’ from the mempool, if $st_{i+1}=st_{i}\setminus{}tx_{i}^{\prime}\cup{}tx_{i},dc_{i+1}=dc_{i}$.
+1) An arriving transaction $tx_{i}$ is admitted with evicting a transaction $tx_{i}^{\prime}$ from the mempool, if $st_{i+1}=st_{i}\setminus{}tx_{i}^{\prime}\cup{}tx_{i},dc_{i+1}=dc_{i}$.
 2) $tx_{i}$ is admitted without evicting any transaction in the mempool, if $st_{i+1}=st_{i}\cup{}tx_{i},dc_{i+1}=dc_{i}$.
 3) $tx_{i}$ is declined and does not enter the mempool, if $st_{i+1}=st_{i},dc_{i+1}=dc_{i}\cup{}tx_{i}$.
-3.2 Bug Oracles
-Suppose a mempool of initial state $\langle{}st_{0},dc_{0}\rangle{}$ receives a sequence of transactions, denoted by $ops$.
-Here, we simplify the problem by not considering arriving blocks during fuzzing.
-We define two bug oracles: eviction-based attacks and declining-based locking attacks.
+Figure 1: Overview of exploit discovery and evaluation workflow: 0) mpfuzz setup, 1) mempool reduction, 2) fuzzing on reduced mempool under test (MUT) to discover short exploits, and 3) exploit extension.
+The extended exploits are 4) evaluated on the actual mempool of the original size.
+Green means automated tasks, and gray requires manual effort.
+4.2 Exploit Discovery Workflow
+To set up the stage, we present the workflow overview in this work.
+The workflow is depicted in Figure 1.
+To discover vulnerabilities and exploits in an actual mempool, one has to 0) set up mpfuzz by various parameters (as will be introduced), 1) reduce the mempool from its original configuration to a much smaller version, named mempool under test (MUT).
+2) mpfuzz is run on the reduced MUT deployed locally (with settings described in § 7.1).
+Fuzzing leads to the discovery of short exploits.
+3) Short exploits are extended to the longer ones, that is, actual exploits.
+4) The success of the actual exploit is evaluated on the actual mempools deployed in close-to-operational settings, such as using popular Ethereum testnets like Goerli (see § 6.3.1) or a network we set up locally running unmodified Ethereum clients (see § 4 and § D).
+Fuzzing a reduced mempool (MUT) instead of an original one is necessary to ensure the efficient execution of the mpfuzz, which entails many iterations of re-initiating and populating the mempool.
+We discuss and evaluate the possible false positives that could be introduced by fuzzing smaller MUT in § 7.2.
+In this workflow, fully automated is Step 2), that is, the discovery of short exploits on a given MUT (colored green in Figure 1).
+Other steps colored in gray require manual efforts and are described below: In Step 0), mpfuzz setup entails setting parameters in bug oracles (e.g., $\epsilon$ and $\lambda$) and fixed symbols in guiding fuzzing.
+The setup is usually one-time.
+In Step 1), mempool reduction entails reconfiguring the mempool to a much smaller capacity (e.g., MUT length $m=6$ or $m=16$, compared to the original size like $m^{\prime}=6144$ as in Geth).
+In addition, policy-specific parameters in the actual mempool are tuned down proportionally.
+For instance, the Geth mempool buffers up to $py_{1}^{\prime}=1024$ future transactions, and when buffering more than $py^{\prime}_{3}=5120$ pending transactions, the mempool triggers a protective policy to limit the number of transactions sent from the same sender by $py_{2}^{\prime}=16$ In a reduced MUT of $m=6$, these parameters are reduced while retaining the same proportion, such as $py_{1}=1,py_{3}=5,py_{2}=2$.
+The mempool reduction requires expert understanding and manual efforts, and it is by nature the best effort.
+In Step 3), it first de-duplicates the raw short exploits found by mpfuzz.
+Then, for each distinct short exploit, it extends it by repeating the transaction-admission events to fit into the actual mempool, during which one may practice strategies like increasing nonce, switching sender accounts, and measuring normal transactions in the actual mempool.
+In exploit extension, one could also consider the factors not modeled in MUT or mpfuzz such as $BasePrice$, as seen in the case in § 6.4.
+4.3 Bug Oracle: Definitions & Rationale
+Definition 4.1 (Tx admission timeline)
+In a transaction admission timeline, a mempool under test is initialized at state $\langle{}st_{0},dc_{0}=\varnothing\rangle{}$, receives a sequence of arriving transactions $ops$, and ends up with state $\langle{}st_{n},dc_{n}\rangle{}$.
+Then, a validator continually builds blocks by selecting and clearing transactions in the mempool until the mempool is empty.
+This transaction timeline is denoted by $\langle{}st_{0},dc_{0}=\varnothing\rangle{},ops\Rightarrow{}\langle{}st_{n},dc_{n}\rangle{}$.
+The transaction admission timeline is simplified from the timeline that an operational mempool experiences where transaction arrival and block building can be interleaved.
+The simplification is intended to trade off the accuracy of fuzzing results for fuzzing efficiency as discussed in § 7.2.
+We use two mempool-attack templates to define two bug oracles: an eviction-based DoS where adversarial transactions evict existing victim transactions in the mempool, and a locking-based DoS where existing adversarial transactions in the mempool decline arriving victim transactions.
 Definition 4.2 (Eviction bug oracle)
-A mempool execution from initial state $st_{0}$ to end state $st_{n}$ under transaction sequence $ops$ is a successful ADAMS eviction attack, if.f.
-1) the attack damage is full, that is, end state $st_{n}$ contains only adversarial transactions and initial state $st_{0}$ contains only benign transactions, and 2) the attack cost is low, that is, the total adversarial transaction fees in the end states are smaller, by a multiplicative factor $\epsilon$, than the attack damage measured by the victim transaction fees in the initial state.
+A transaction admission timeline, $\langle{}st_{0},dc_{0}=\varnothing\rangle{},ops\Rightarrow{}\langle{}st_{n},dc_{n}\rangle{}$, is a successful ADAMS eviction attack, if.f.
+1) the admission timeline causes full damage, that is, initial state $st_{0}$ contains only benign transactions, $ops$ are adversarial transactions, and the result end state $st_{n}$ contains only adversarial transactions (in Equation 1), and 2) the attack cost is asymmetrically low, that is, the total adversarial transaction fees in the end state $st_{n}$ to be charged are smaller, by a multiplicative factor $\epsilon$, than the attack damage measured by the fees of evicted transactions in the initial state $st_{0}$ (in Equations 2 and 3).
 Formally,
-$st_{0}\cap{}(st_{n})$ $=$ $\varnothing$
-$\sum_{tx\in{}st_{n}}tx.fee$ $<$ $\epsilon\sum_{tx\in{}st_{0}}tx.fee$ (1)
+$st_{0}\cap{}st_{n}$ $=$ $\varnothing$ (1)
+$asym_{E}(st_{0},ops)$ ]$\stackrel{{\scriptstyle[}}{{]}}{def}{=}$ $\frac{\sum_{tx\in{}st_{n}}tx.fee}{\sum_{tx\in{}st_{0}}tx.fee}$ (2)
+$asym_{E}(st_{0},ops)$ $<$ $\epsilon$ (3)
 Definition 4.3 (Locking bug oracle)
-A mempool execution from initial state $\langle{}st_{0},dc_{0}\rangle{}$ to end state $\langle{}st_{n},dc_{n}\rangle{}$ under transaction sequence $ops$ is a successful ADAMS locking attack if.f.
-1) the attack damage is full, that is, the victim normal transactions in $ops$ are completely declined by the mempool storing adversarial transactions, where normal transactions and adversarial transactions are sent by different accounts, and 2) the attack cost is low, that is, the total adversarial transaction fees in the mempool are smaller, by a multiplicative factor $\lambda$, than the attack damage measured by the fees of victim transactions declined.
+A transaction admission timeline, $\langle{}st_{0}=\varnothing,dc_{0}=\varnothing\rangle{},ops\Rightarrow{}\langle{}st_{n},dc_{n}\rangle{}$, is a successful ADAMS locking attack, if.f.
+1) the admission timeline causes full damage, that is, the mempool is first occupied by adversarial transactions (i.e., $st_{n}$ contains only adversarial transactions) and then declines the arriving normal transactions (i.e., $dc_{n}$ contains only normal transactions); the normal transactions and adversarial transactions are sent by different accounts (i.e., Equation 4), and 2) the attack cost is asymmetrically low, that is, the average adversarial transaction fees in the mempool are smaller, by a multiplicative factor $\lambda$, than the attack damage measured by the average fees of victim transactions declined (i.e., Equations 5 and 6).
 Formally,
-$\cup_{txd\in{}dc_{n}}{txd.sender}$$\cap$ $\cup_{tx\in{}s_{n}\cup{}ops\setminus{}dc_{n}}{tx.sender}=\varnothing$
-$\sum_{tx\in{}st_{n}}tx.fee$ $<$ $\lambda\sum_{tx\in{}dc_{n}}tx.fee$ (2)
+$\cup_{tx\in{}dc_{n}}{tx.sender}$$\cap$ $\cup_{tx\in{}st_{n}}{tx.sender}=\varnothing$ (4)
+$asym_{D}(ops)$ ]$\stackrel{{\scriptstyle[}}{{]}}{def}{=}$ $\frac{\sum_{tx\in{}st_{n}}tx.fee/\|st_{n}\|}{\sum_{tx\in{}dc_{n}}tx.fee/\|dc_{n}\|}$ (5)
+$asym_{D}(ops)$ $<$ $\lambda$ (6)
+Design rationale: Both definitions use the normal transaction fees to measure the attack damage because these fees are not collected by the validator in the timeline under attack and are collectible without attack.
+Specifically, the damage in Definition 4.2 is measured by the fees of normal transactions in $st_{0}$ which are evicted in end state $st_{n}$ and which would not have been evicted had there been no adversarial transactions in $ops$.
+The damage in Definition 4.3 is measured by the fees of normal transactions in $ops$ which are declined from end state $st_{n}$ and which would not have been declined had there been no adversarial transactions in $ops$.
+Both definitions are strict and require causing full damage, such as evicting or declining all normal transactions.
+Strict definitions are necessary to ensure that what is found as exploits on the small MUT is true positive and works on the actual mempool.
+The targeted attacks by this work are denial of mempool service in feeding downstream validators with valid transactions sent from benign users.
+They are different from other forms of DoS in blockchains, including resource exhaustion via executing malicious smart contracts [35, 27].
+Particularly, in the threat model of mempool DoS, we assume the validators are benign and functional in building blocks.
+Because in all Ethereum clients we know, transactions are admitted into mempool without execution (i.e., executing the smart contracts they invoke), GasUsed is not a factor in transaction admission.
+[^2] The further design rationale for admitting transactions without execution is that an Ethereum transaction $tx$’s GasUsed is non-deterministic unless the ordering of $tx$ w.r.t.
+other transactions in the block is fixed and $tx$ is executed.
+Thus, our bug oracle does not vary GasUsed.
+Instead, all the transactions are fixed at $21000$ Gas (i.e., as if they don’t invoke smart contracts).
+[^3] We may use price and fees interchangeably.
+This design significantly reduces the search space without loss in covering different mempool behaviors.
+Besides, our bug oracle does not model the effect of EIP1559 or $BasePrice$.
+Without modeling $BasePrice$, our bug oracle can still capture the alternative transaction fee (the product of GasPrice and $21000$ Gas), which is still proportional to actual $PriorityPrice$ or damage on validator revenue.
 5 Stateful Mempool Fuzzing by mpfuzz
-4.1 Motivating Observations
-(In)feasibility of code-coverage-only stateless fuzzing: To fuzz a mempool, a baseline design is to use a code-coverage-based fuzzer.
-In each iteration, the fuzzer generates a bit string, parses it into a sequence of transactions and sends them to the tested mempool for execution.
-Upon each end state, it checks whether the test oracle defined above is satisfied.
-This design takes code coverage as the only feedback and is stateless because the feedback does not consider the end state.
-The stateless design would be ineffective in fuzzing mempool or finding ADAMS exploits.
-For example, consider fuzzing a three-slot mempool to find a DETER-X exploit consisting of three future transactions [22].
-Suppose the current mempool input is a sequence of one future transaction.
-After mutation, it may try an input of two future transactions, which, however, does not increase code coverage (the second future transaction is executed by the same code path in a mempool as the first future transaction).
-Thus, the fuzzer discards the two-future-transaction input and misses finding the three-future-transaction exploit.
-We validate the ineffectiveness of stateless fuzzer by experiments in § 6.
-A more promising design is stateful mempool fuzzing, in which the mempool state is included in the feedback to guide next-iteration fuzzing.
-In the previous example of finding a three-future-transaction exploit (DETER-X), sending two future transactions leads to a different mempool state than sending one.
-A simple state-coverage-based fuzzer would view this as positive feedback and further explore this direction toward finding the three-future-transaction exploit.
-(In)feasibility of existing consensus fuzzers: Existing consensus fuzzers [19, 23, 30] cannot detect mempool DoS or ADAMS bugs.
-Specifically, Loki’s test oracle is detecting system crashes, while a successful ADAMS attack does not necessarily trigger system crashes.
-Using Loki to detect ADAMS would cause many false negatives.
-Fluffy detects the unsynchronizable difference of post-consensus states as vulnerability.
-However, the mempools in two clients, say Geth and Besu under benign transactions, would be different.
-In other words, post-consensus state differences across clients could indicate insecurity, but the difference in pre-consensus states across clients, like mempool, does not mean insecurity.
-Using Fluffly to detect ADAMS could cause false positives.
-Tyr’s test oracle detects property violations on post-consensus states.
-Particularly, while Tyr models state liveness, their definition “valid transactions should be executed eventually” is, unfortunately, over-simplifying and fails to model the legitimate pre-consensus cases in real Ethereum clients; for instance, valid transactions of low Gas price can be dropped by Ethereum mempools and will not be eventually executed.
-Besides, they model double-spending transactions as the only type of invalid transaction.
-In Ethereum, double-spending transactions are of the same nonce.
-And there are many more sophisticated invalid transactions in Ethereum that Tyr does not model, including future transactions, latent overdrafts, etc.
-We analyze the case of existing consensus fuzzers in detecting known DETER attacks [22].
-Loki cannot detect DETER because DETER does not cause system crashes.
-Fluffy does not model future transactions and can not find the DETER attacks that rely on future transactions.
-Similarly, Tyr that does not model sophisticated invalid transactions would miss detecting DETER attacks.
-Challenges: In stateful mempool fuzzing, the essential design problem is how to generate the next input transaction given a current “state”.
-State $st$ is the collection of transactions in the mempool and those included in the blockchain.
-The key challenge is the large input space of transactions.
+5.1 Transaction Symbolization
+The key challenge in designing mpfuzz is the large input space of transaction sequences.
 Recall that an Ethereum transaction consists of multiple attributes (sender, nonce, GasPrice, value, etc.), each defined in a large domain (e.g., 64-bit string).
 Exploring the raw transaction space is hard; the comprehensive search is inefficient, and randomly trying transactions as done in the state-of-the-art blockchain fuzzers (in § 2) is ineffective.
-4.2 Proposed Fuzzer Design
-4.2.1 Transaction Symbolization
 Intuition: We propose symbolizing transactions for efficient and effective mempool stateful fuzzing.
-Our key idea is to map each group of concrete transactions triggering an equivalent mempool behavior into a distinct symbol so that searching one transaction is sufficient to cover all other transactions under the same symbol.
+Our key idea is to map each group of concrete transactions, triggering an equivalent mempool behavior into a distinct symbol so that searching for one transaction is sufficient to cover all other transactions under the same symbol.
 Transaction symbolization is expected to reduce the search space from possible concrete transactions to the symbol space.
-Table 2: Symbols, transactions and associated costs. Tx refers to the instantiated transaction under a given symbol.
-Symbol Description
-Tx
-$cost()$ $opcost()$
-$\mathcal{N}$ Benign
-$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B*}&{1}\\ {*}&{3}\end{bmatrix}$}}$
-$3$ $3$
-$\mathcal{F}$ Future
-$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+1}\end{bmatrix}$}}$
-00 00
-$\mathcal{P}$ Parent
-$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A>r}&{1}\\ {1}&{[4,m]}\end{bmatrix}$}}$
-$[4,m]$ $[4,m]$
-$\mathcal{C}$ Child
-$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$
-$m+1$ $1$
-$\mathcal{O}$ Overdraft
-$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m+1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$
-00 00
-$\mathcal{L}$ Latent overdraft
-$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$
-00 00
-$\mathcal{R}$ Replacement
-$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {1}&{m+1}\end{bmatrix}$}}$
-00 00
+In this work, we manually design seven symbols to represent transactions based on the Ethereum “semantics”, that is, how different transactions are admitted by the mempool.
+Table 2: Symbols, transactions, and associated costs. Tx refers to the instantiated transaction under a given symbol.
+Symbol Description Tx $cost()$ $opcost()$
+$\mathcal{N}$ Benign $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B*}&{1}\\ {*}&{3}\end{bmatrix}$}}$ $3$ $3$
+$\mathcal{F}$ Future $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+4}\end{bmatrix}$}}$ 00 00
+$\mathcal{P}$ Parent $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A>r}&{1}\\ {1}&{[4,m+3]}\end{bmatrix}$}}$ $[4,m+3]$ $[4,m+3]$
+$\mathcal{C}$ Child $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{1}\\ {\geq{}2}&{m+4}\end{bmatrix}$}}$ $m+4$ $1$
+$\mathcal{O}$ Overdraft $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m+1}\\ {\geq{}2}&{m+4}\end{bmatrix}$}}$ 00 00
+$\mathcal{L}$ Latent overdraft $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {\geq{}2}&{m+4}\end{bmatrix}$}}$ 00 00
+$\mathcal{R}$ Replacement $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {1}&{m+4}\end{bmatrix}$}}$ $m+4$ 00
 Specification: Suppose the current state contains adversarial transactions sent from $r$ accounts $A_{1},\dots,A_{r}$.
 These accounts have an initial balance of $m$ Ether, where $m$ equals the mempool capacity (say $m=1000$).
+The rationale is that these accounts can send at most $m$ adversarial transactions, each minimally spending one Ether, to just occupy a mempool of $m$ slots.
+A larger value of attacker balance is possible but increases the search space.
 Symbol $\mathcal{N}$ defines a transaction subspace covering any normal transactions sent from any benign account and of any nonce, namely $tx\begin{bmatrix}{B*}&{*}\\ {*}&{*}\end{bmatrix}$ In mpfuzz, Symbol $\mathcal{N}$ is instantiated to concrete transactions of fixed GasPrice $3$ wei and of value $1$ Ether.
@@
 the current state.
-mpfuzz instantiates symbol $\mathcal{F}$ to concrete transactions of nonce $m+1$, value $1$ Ether, Gas price $m+1$ wei, and any adversarial account $*A*$.
-The instantiating pattern for symbol $\mathcal{F}$ is $tx\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+1}\end{bmatrix}$.
+mpfuzz instantiates symbol $\mathcal{F}$ to concrete transactions of nonce $m+1$, value $1$ Ether, GasPrice $m+4$ wei, and any adversarial account $A*$.
+The instantiating pattern for symbol $\mathcal{F}$ is $tx\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+4}\end{bmatrix}$.
 Symbol $\mathcal{P}$ defines the transaction subspace of any parent transaction from an adversarial account w.r.t.
 the current state.
-mpfuzz instantiates $\mathcal{P}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A\geq{r}}&{1}\\ {1}&{[4,m]}\end{bmatrix}$}}$.
-Here, the transaction is fixed with a Gas price in the range $[4,m]$ wei, value at $1$ Ether, and nonce at $1$.
+mpfuzz instantiates $\mathcal{P}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A\geq{r}}&{1}\\ {1}&{[4,m+3]}\end{bmatrix}$}}$.
+Here, the transaction is fixed with a GasPrice in the range $[4,m+3]$ wei, value at $1$ Ether, and nonce at $1$.
+The GasPrice of a future transaction ($\mathcal{F}$) is $m+4$ wei, which is higher than the price of a parent transaction ($\mathcal{P}$) in $[4,m+3]$ wei.
+The purpose is to ensure that a future transaction can evict any parent transaction during fuzzing.
 Symbol $\mathcal{C}$ defines the transaction subspace of any adversarial child transaction w.r.t.
 the state.
-mpfuzz instantiates $\mathcal{C}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$.
-The instantiated child transaction is fixed at Gas price of $m+1$
+mpfuzz instantiates $\mathcal{C}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{1}\\ {\geq{}2}&{m+4}\end{bmatrix}$}}$.
+The instantiated child transaction is fixed at GasPrice of $m+4$
 Symbol $\mathcal{O}$ defines the transaction subspace of any adversarial overdraft transaction w.r.t.
 state $st$.
-mpfuzz instantiates $\mathcal{O}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m+1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$.
+mpfuzz instantiates $\mathcal{O}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m+1}\\ {\geq{}2}&{m+4}\end{bmatrix}$}}$.
 Recall that all adversarial accounts have an initial balance of $m$ Ether.
@@
 state $st$.
-mpfuzz instantiates $\mathcal{L}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$.
+mpfuzz instantiates $\mathcal{L}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {\geq{}2}&{m+4}\end{bmatrix}$}}$.
 All adversarial accounts have an initial balance of $m$ Ether.
@@
 Unlike the symbols above, the transactions under Symbol $\mathcal{R}$ must share the same sender and nonce with an existing transaction in the current state $st$.
-mpfuzz instantiates $\mathcal{R}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {1}&{m+1}\end{bmatrix}$}}$.
+mpfuzz instantiates $\mathcal{R}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {1}&{m+4}\end{bmatrix}$}}$.
 Here, we simplify the problem and only consider replacing the transaction with nonce $1$.
@@
 In our design, the instantiated transactions are a much smaller subset of the defining transaction space.
-For instance, given state $st$ of transactions sent from account $A_{1},\dots,A_{r}$, transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ is within the defining space of symbol $\mathcal{F}$, but mpfuzz does not instantiate $\mathcal{F}$ by transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ (but instead to transactions $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+1}\end{bmatrix}$}}$).
+For instance, given state $st$ of transactions sent from account $A_{1},\dots,A_{r}$, transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ is within the defining space of symbol $\mathcal{F}$, but mpfuzz does not instantiate $\mathcal{F}$ by transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ (but instead to transactions $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+4}\end{bmatrix}$}}$).
+The design rationale of transaction symbolization is deferred to Appendix A.1.
 5.1.1 State Search Algorithm
@@
 In each iteration, the algorithm retrieves the next seed from the corpus based on how promising the seed is in reaching a state triggering the oracle (a.k.a., the energy [22]).
-A seed consists of a symbolic input $ops$, which consists of transaction symbols, and a symbolic state $st$ reached by running a transaction sequence instantiated from the input $ops$ (Line 3).
+A seed consists of a symbolized input $ops$, which consists of transaction symbols, and a symbolized state $st$ reached by running a transaction sequence instantiated from the input $ops$ (Line 3).
 The algorithm runs an inner loop, in each iteration of which it mutates the input in the current seed ($\langle{}st,in\rangle{}$) and executes the mutated input against a reinitialized empty mempool, producing end state $st^{\prime}$ (Line 4).
@@
 Across different senders, symbols are ordered by the parent’s GasPrice.
-For instance, suppose a mempool stores four concrete transactions, $tx\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$, $tx\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$, $tx\begin{bmatrix}{A_{2}}&{*}\\ {1}&{5}\end{bmatrix}$, $tx\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10001}\end{bmatrix}$ , they are mapped to four symbols, $\mathcal{P},\mathcal{F},\mathcal{P},\mathcal{C}$, and are further ordered in a symbolic state by $\mathcal{F}\mathcal{P}\mathcal{C}\mathcal{P}$.
-In mpfuzz, the feedback of an input $ops^{\prime}$ is based on the symbolic end state $st^{\prime}$.
+For instance, suppose a mempool stores four concrete transactions, $tx\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$, $tx\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$, $tx\begin{bmatrix}{A_{2}}&{*}\\ {1}&{5}\end{bmatrix}$, $tx\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10001}\end{bmatrix}$ , they are mapped to four symbols, $\mathcal{P},\mathcal{F},\mathcal{P},\mathcal{C}$, and are further ordered in a symbolized state by $\mathcal{F}\mathcal{P}\mathcal{C}\mathcal{P}$.
+In mpfuzz, the feedback of an input $ops^{\prime}$ is based on the symbolized end state $st^{\prime}$.
 Positive feedback on state $st^{\prime}$ is determined conjunctively by two metrics: 1) State coverage that indicates state $st^{\prime}$ is not covered in the corpus, and 2) state promising-ness that indicates how promising the current state is to reach a state satisfying bug oracles.
-$feedback(st^{\prime},st,sdb)=1$ , if $\text{if }st\_coverage(st^{\prime},sdb)=1\land{}$ (3)
-$st\_promising(st^{\prime},st)=1$
-Specifically, state coverage is determined by straightforwardly comparing symbolic state $st^{\prime}$, as an ordered list of symbols, with all the symbolic states in the corpus.
+$feedback(st^{\prime},st,sdb)$ $=$ $st\_coverage(st^{\prime},sdb)==1\land{}$ (7)
+$st\_promising(st^{\prime},st)==1$
+Specifically, state coverage is determined by straightforwardly comparing symbolized state $st^{\prime}$, as an ordered list of symbols, with all the symbolized states in the corpus.
 For instance, state $\mathcal{F}\mathcal{P}\mathcal{C}\mathcal{P}$ is different from $\mathcal{F}\mathcal{P}\mathcal{P}\mathcal{C}$.
-This also implies concrete transaction sequence $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{2}}&{*}\\ {1}&{5}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10001}\end{bmatrix}$}}$ is the same with $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{2}}&{*}\\ {1}&{6}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10}\end{bmatrix}$}}$, as they are both mapped to the same symbolic state.
+This also implies concrete transaction sequence $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{2}}&{*}\\ {1}&{5}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10001}\end{bmatrix}$}}$ is the same with $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{2}}&{*}\\ {1}&{6}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10}\end{bmatrix}$}}$, as they are both mapped to the same symbolized state.
 How promising a state $st^{\prime}$ is (i.e., $st\_promising(st^{\prime})$) is determined as follows: A mutated state $st^{\prime}$ is more promising than an unmutated state $st$ if any one of the three conditions is met: 1) State $st^{\prime}$ stores fewer normal transactions (under symbol $\mathcal{N}$) than state $st$, implying more transactions evicted and higher damage (i.e., $evict\_normal(st^{\prime},st)=1$).
@@
 3) The total fees of adversarial transactions in state $st^{\prime}$ are lower than those in state $st$, implying lower attack costs.
-We consider two forms of costs: concrete cost and symbolic cost.
-The former, denoted by $cost(st^{\prime})$, is simply the total fees of adversarial transactions instantiated from a symbolic state $st^{\prime}$.
+We consider two forms of costs: concrete cost and symbolized cost.
+The former, denoted by $cost(st^{\prime})$, is simply the total fees of adversarial transactions instantiated from a symbolized state $st^{\prime}$.
 The latter, denoted by $opcost(st^{\prime})$, optimistically estimates the cost of transactions in the current state $st^{\prime}$ that contributes to a future state triggering test oracle.
-We will describe the symbolic cost next.
+We will describe the symbolized cost next.
 Formally, state promising-ness is calculated by Equation 8.
-$st\_promising(st^{\prime},st)=$ $1,$ if $\text{if }evict\_normal(st^{\prime},st)=1\lor{}$ (4)
-$decline\_normal(st^{\prime},st)=1\lor{}$
+$st\_promising(st^{\prime},st)$ $=$ $evict\_normal(st^{\prime},st)==1\lor{}$ (8)
+$decline\_normal(st^{\prime},st)==1\lor{}$
 $cost(st^{\prime})<cost(st)\lor{}$
@@
 Otherwise, $b=0$.
-Second, we use the symbolic cost to estimate how promising a state is and use it in seed energy.
+Second, we use the symbolized cost to estimate how promising a state is and use it in seed energy.
 Estimate state cost: When mpfuzz determines how promising a state is, it needs to look beyond the current state and into all possible descendant states.
@@
 6 Found Exploits
-Running mpfuzz on six leading Ethereum execution-layer clients, namely, Geth, Besu, Nethermind, Erigon, Reth, and OpenEthereum, leads to the discovery of $22$ bugs on all six Ethereum clients, including $7$ conforming to the known DETER attacks on clients of historical versions and $15$ new bugs on the clients of the latest versions.
-We describe two selected patterns of these bugs as follows.
-5.1 Found Eviction Attacks (Selected)
-Exploit $XT_{4}$: Indirect eviction by valid-turned-overdraft transactions: This eviction attack works in two steps: 1) The attacker first sends valid transactions at high fees to evict normal transactions initially stored in the mempool.
+Table 3: ADAMS exploit patterns found by mpfuzz across Ethereum clients; $XT_{1-7}$ are eviction based, and $XT_{8-9}$ are locking based.
+$XT_{1-3}$ are known patterns in DETER [28], while others are new.
+✗ indicates the presence of a bug, indicates the fixing of a bug after our reporting, and ✓ indicates the fixing of a bug by previous works.
+$XT_{1}$ $XT_{2}$ $XT_{3}$ $XT_{4}$ $XT_{5}$ $XT_{6}$ $XT_{7}$ $XT_{8}$ $XT_{9}$
+[…]
+Geth $<v1.11.4$, EigenPhi
+go-opera $\leq{}v1.1.3$
+Erigon $\leq{}v2.42.0$
+Besu $\geq{}v22.7.4$
+Besu $<v22.7.4$
+Nethermind $\geq{}v1.18.0$
+Nethermind $<v1.18.0$
+[…]
+[…]
+OpenEthereum $\leq{}v3.3.5$
+We have run mpfuzz across a variety of Ethereum clients, including six leading execution-layer clients on the public-transaction path of Ethereum mainnet (Geth, Besu, Nethermind, Erigon, Reth, and OpenEthereum), three PBS clients (proposer-builder separation) on the mainnet’s private-transaction and bundle path (Flashbot builder $v1.11.5$ [6], EigenPhi builder [4] and bloXroute builder-ws [1]), and the clients deployed on three operational Ethereum-like networks (BSC $v1.3.8$ [2] deployed on Binance Smart Chain, go-opera $v1.1.3$ [9] on Fantom, and core-geth $v1.12.19$ [3] on Ethereum Classic).
+On the six public-transaction clients, mpfuzz leads to the discovery of $22$ bugs, as listed in Table 3, including $7$ conforming to the known DETER attacks on clients of historical versions and $15$ new bugs on the clients of the latest versions.
+Other clients, including the three PBS clients and three clients on Ethereum-like networks, are mostly forks of the Geth clients except for minor code changes [^4] For instance, go-opera adopts its own fix against $XT_{1}$ based on the fork of Geth $<v1.11.4$., and on them, mpfuzz discovered $13$ bugs of a similar nature to those found on Geth (of the historical and latest versions), as seen in Table 3.
+We describe the patterns of these bugs as follows.
+6.1 Found Eviction Attacks
+Exploit $XT_{1}$: Direct eviction by future transactions: In this attack, given a mempool’s initial state storing normal transactions, the attacker sends future transactions at high GasPrice to evict the normal transactions.
+This exploit is essentially the DETER-X attack [28].
+In practice, Geth ($\leq{}v1.11.4$), Nethermind, Besu, and EigenPhi are vulnerable under this exploit, as in Table 3.
+Exploit $XT_{2}$: Direct eviction by latent overdraft transactions: In this eviction attack, the attacker sends latent-overdraft transactions at high GasPrice to evict the normal transactions initially stored in the target mempool.
+These transactions are sent from $k$ accounts, each of which sends $l$ transactions, denoted by $k\times{}l$.
+The intention is to evade the limit on the number of transactions per sender.
+The evasion increases the attacker cost from one pending transaction to multiple.
+This exploit is essentially the DETER-Z attack [28].
+In practice, Geth ($\leq{}v1.11.4$), Besu, EigenPhi, and go-opera are found vulnerable under this exploit.
+Exploit $XT_{3}$: Compositional direct eviction (by combining $XT_{1}$ and $XT_{2}$): In some Ethereum clients, notably Geth, the limit on the number of transactions per sender is triggered under the condition that the mempool stores enough pending transactions (e.g., more than 5120 transactions in Geth).
+This eviction attack combines $XT_{1}$ and $XT_{2}$ to avoid triggering the condition and evade the protection.
+It works by maximizing the eviction of mempool under $XT_{1}$ until it is about to trigger the condition.
+It then conducts $XT_{2}$ by sending latent overdraft transactions under one sender, that is, $1\times{}l$.
+Compared to $XT_{2}$, the combined exploit $XT_{3}$ achieves lower costs.
+In practice, Geth ($\leq{}v1.11.4$) is vulnerable under this exploit where its mempool of capacity of 6144 slots triggers the limit of 16 transactions per sender when there are more than 5120 pending transactions.
+That is, $XT_{3}$ is configured with $l=5120$.
+Other clients including EigenPhi and go-opera are also vulnerable.
+Exploit $XT_{4}$: Indirect eviction by valid-turned-overdraft transactions: This eviction attack works in two steps: 1) The attacker first sends valid transactions at high GasPrice to evict normal transactions initially stored in the mempool.
 These transactions are sent from $k$ accounts, each of which sends $l$ transactions.
-2) She then sends $k$ transactions; each of the transactions is of nonce $1$, at a high fee, of high value $v_{1}$, and from the same $k$ accounts in Step 1).
+2) She then sends $k$ transactions; each of the transactions is of nonce $1$, at a high GasPrice, of high value $v_{1}$, and from the same $k$ accounts in Step 1).
 The transactions would replace the transaction of the same sender and nonce sent in Step 1).
@@
 Specifically, if a sender’s balance is $bal$ and the Ether value of transaction of nonce 2 is $v_{2}$, then $v_{1}$ is carefully crafted to enable turned latent overdraft, that is, $v_{1}<bal$ and $v_{1}+v_{2}>bal$.
-In practice, many clients are found vulnerable under this exploit, including Geth $<v1.11.4$, Erigon, Besu, Nethermind, and OpenEthereum.
-Exploit $XT_{6}$: Compositional indirect eviction (by multi-round valid-turned-future transactions): The latest Geth $\geq{}v1.11.4$ is patched against direct eviction attacks.
-To evade the protection, this eviction attack uses indirect eviction.
-The attack works in two steps: 1) It first control-evicts the mempool using valid-turned-future transactions to avoid the triggering condition of its limit of transactions per sender.
-2) It then conducts another run of valid-turned-future transactions, this time with transactions sent from one account, namely $1\times{}l$.
-It can succeed since the limit is not triggered.
-The overall attack cost is the fee of one transaction; compared to the single run of valid-turned-future attack, $XT_{6}$’s cost is much lower.
-5.2 Found Locking Attacks (Selected)
+In practice, many clients are found vulnerable under this exploit, including Geth $<v1.11.4$, Erigon, Besu, Nethermind, OpenEthereum, EigenPhi, and go-opera.
+Exploit $XT_{5}$: Indirect eviction by valid-turned-future transactions: This eviction attack works in two steps: 1) The attacker first sends valid transactions at high fees to evict normal transactions initially stored in the mempool.
+These transactions are sent by $k\times{}l$, that is, from $k$ accounts, each with $l$ transactions.
+For each sender, the transaction of nonce 1 has a fee, say $f_{1}$, slightly lower than the transactions of other nonces, that is, $f_{1}<f_{2}$.
+2) She then sends $k$ transactions; each of the transactions is from a distinct sender from those used in Step 1) and of fee $f^{\prime}$ that $f_{1}<f^{\prime}<f_{2}$.
+The intent is that the $k$ transactions in Step 2) evict the transactions of nonce $1$ sent in Step 1), turning other child transactions sent in Step 1) into future transactions.
+Exploit $XT_{6}$: Compositional indirect eviction (by multi-round valid-turned-future transactions): This exploit is an adaptive attack to the Geth $\geq{}v1.11.4$.
+Geth $\geq{}v1.11.4$ is patched against $XT_{1-2}$ and adopts the following admission policies: The mempool of $m^{\prime}=6144$ slots admits up to $py_{1}^{\prime}=1024$ future transactions.
+When containing more than $py_{3}^{\prime}=5120$ pending transactions, the mempool starts to limit that no more than $py_{2}^{\prime}=16$ pending transactions from the same sender can be admitted.
+Exploit $XT_{6}$ works in three steps: 1) it first evicts the initial mempool of normal transactions with $m^{\prime}/py_{2}^{\prime}=384$ sequences, each of $16$ transactions sent from one unique sender, 2) it then sends $py_{1}^{\prime}/(py_{2}^{\prime}-1)=65$ transactions to evict the parent transactions and creates $py_{1}^{\prime}=1024$ future transactions in the mempool, 3) it sends one sequence of $5120$ transactions from one sender to evict the normal transactions sent in Step 1), and 4) it sends one transaction to evict the parent transaction in Step 3, leaving the mempool of just one valid transaction.
+Step 3) can succeed because the precondition (w.r.t.
+$py_{3}^{\prime}=5120$) of limiting transactions of the same sender does not hold.
+In practice, Geth of all versions and the Geth forks, including Flashbot, bloXroute, BSC, and Ethereum Classic are found vulnerable under $XT_{6}$.
+Exploit $XT_{7}$: Reversible evictions: Suppose a mempool of state $st$ admits an arriving transaction $tx$ by evicting an existing transaction $tx^{\prime}$, transitioning its state to $st^{\prime}$.
+A mempool is reversible if one sends $tx^{\prime}$ to state $st^{\prime}$, and the mempool admits $tx^{\prime}$ by evicting $tx$, transitioning its state back to $st$.
+The mempool on Nethermind $<v1.18.0$ can be reversible: Minimally, suppose a two-slot mempool stores $tx_{1}$ of low GasPrice from one sender and $tx_{2}$ of medium GasPrice from another sender $A2$ and receives $tx_{3}$ as a child of $tx_{1}$ and of high GasPrice.
+For instance, $tx_{1}\begin{bmatrix}{A}&{1}\\ {1}&{1}\end{bmatrix}$, $tx_{2}\begin{bmatrix}{B}&{1}\\ {1}&{3}\end{bmatrix}$, $tx_{3}\begin{bmatrix}{A}&{1}\\ {2}&{5}\end{bmatrix}$.
+The Nethermind $<v1.18.0$ mempool would admit $tx_{3}$ and evict $tx_{2}$.
+After that, if one resends the evicted $tx_{2}$ back to the mempool storing $tx_{1}$ and $tx_{3}$, the Nethermind $<v1.18.0$ mempool would admit $tx_{2}$ and evict $tx_{3}$, looping back to its initial state.
+An attacker observing reversible mempool mounts Exploit $XT_{7}$ to evict the mempool while bringing down attack costs.
+On Nethermind $<v1.18.0$, the attack works in two steps: 1) She first sends $k$ transaction sequences from $k$ senders, [^5] Nethermind does not limit transactions per sender, thus $k=1$ in actual attacks.
+each of $l$ transactions (i.e., $k\times{}l$).
+In each $l$-transaction sequence, the nonce-$1$ transaction has a low GasPrice, say $f_{1}$, while all its child transactions are of high GasPrice, evicting normal transactions initially stored in the mempool.
+2) The attacker sends $k\times(l-1)$ transactions from $k\times(l-1)$ new senders, all with GasPrice slightly higher than $f_{1}$.
+The mempool would admit these transactions to the mempool by evicting all child transactions sent in Step 1), bringing down the attack costs.
+6.2 Found Locking Attacks
 […]
@@
 Any transactions that arrive after will encounter a full mempool and are declined.
-6 Performance Evaluation of mpfuzz
+Exploit $XT_{9}$: Locking mempool of no turning: Because of the risk of evicting a parent transaction that could lead to turning, certain mempool is designed to restrict eviction victims to child transactions; that is, only the transaction of maximal nonce under a given sender can be evicted.
+For instance, given an arriving transaction $tx$, the OpenEthereum mempool finds an existing child transaction $tx^{\prime}$ with a lower GasPrice than $tx$ and evicts it to admit $tx$.
+In Exploit $XT_{9}$, the attacker observes empty slots in the mempool (e.g., created by arriving blocks) and sends the same number of transactions to occupy them.
+The transactions are sent from one account where the child transaction has a higher GasPrice $f_{1}$ than normal transactions, while all its parent transactions have minimal fees.
+A normal transaction that arrives subsequently is declined because $f_{1}$ is higher than a normal transaction.
+Because each of its parent transactions has a low GasPrice, the mempool is locked at a low cost.
+6.3 Evaluation of Found Turning Exploits
+This and next subsections show the evaluation of turning attacks and locking attacks, and due to the space limit, the evaluation of other ADAMS attacks is deferred to Appendix D.
+We design experiments to measure the success of ADAMS attacks in real Ethereum networks.
+6.3.1 Evaluation in Testnet
+Experimental settings: We set up our experiment platform by connecting the attacker node to the Goerli testnet.
+The attacker node runs an instrumented Geth client at the execution layer which sends crafted transactions to the testnet.
+To monitor transaction propagation, we launch an independent supernode (called measurement node) in the same testnet.
+The measurement node runs a reconfigured Geth client where the limit of peers/neighbors is removed and can connect to as many neighbors as possible.
+The measurement node is not (directly) connected to the attacker node.
+When setting up our experiments, we ran the measurement node in Goerli for seven days and found the node is stabilized at $290$ neighbors.
+Figure 3: Mounting $XT_{4}$ attacks on Goerli: Etherscan screenshot of the blocks generated during the attack
+Experiment results: The instrumented measurement node is able to log the received messages from different neighbors; these messages include those of transactions and of transaction hashes.
+The measurement node could receive the same transaction from different neighbors, and the log stores the transaction-neighbor pairs.
+We started logging the received messages one month before the experiment.
+To do an experiment, we make the attacker node send $XT_{4}$ transactions using $384$ accounts.
+Each account sends $16$ valid pending transactions, followed by a replacement transaction.
+In total, the attacker node sends $6144$ valid transactions and $384$ replacement transactions.
+The GasPrice of the valid and replacement transactions are set to be $8$ Gwei and $130$ Gwei, respectively.
+Finally, we wrap up the experiment by waiting after the attacker node sends all messages and all replacement transactions are included in the blockchain.
+Figure 3 shows the generated blocks in the experiment.
+We took the screenshot from etherscan.io and labeled it in red with information regarding the attack.
+Before the attack begins, the testnet normally utilizes $24-47\%$ of the Gas in a block.
+For ethics, our attacks are short (lasting four blocks).
+Right after the launch of the attack, in block $xx07$, the Gas used by normal transactions drops to $3.19\%$ and the Gas used by adversarial transactions (denoted by red bars in Figure 3) is $26.88\%$.
+The included adversarial transactions are $384$ replacement transactions sent in the second round of $XT_{4}$.
+The $6144-384=5760$ child transactions sent in the first round are not included in the block.
+Other blocks during the attack, namely $xx08$ and $xx10$, are similar.
+In the third block $xx09$, $73.1\%$ Gas is spent on including $82$ normal transactions.To explain it, we inspect the raw history of transactions arriving at and logged by our measurement supernode and found that out of the $82$ included normal transactions, only $4$ are present in the log, implying the other $78$ normal transactions are private ones that were not broadcast to the measurement node.
+Notice that our attacks require no discovery of critical nodes as needed in [28].
+6.4 Evaluation of Found Locking Exploits
+6.4.1 Exploit Extension
+[…]
+Actual exploit $XT_{8a}$: In $XT_{8}$, the attack cost increases with the block $BasePrice$.
+We manually propose a method to decrease the block $BasePrice$ in Ethereum by mounting an eviction attack on the previous block.
+Specifically, in Ethereum (after EIP1559), given a recently produced block $i$, the block $BasePrice$ of block $i+1$ is calculated dynamically as follows:
+$BasePrice(i+1)$ $=$ $BasePrice(i)*[\frac{7}{8}+\frac{1}{4}*\frac{\textit{GasUsed}(i)}{BlockLimit}]$ (10)
+Therefore, if an eviction attack can persistently lower the Gas utilization of recent blocks, the current $BasePrice$ can be reduced, which decreases the costs of locking attacks on the current block.
+In Exploit $XT_{8a}$, we mount a series of eviction attacks first and then a locking attack against the Reth node.
+Specifically, suppose in a network of a Reth node and a block validator node, the $XT_{8a}$ attacker first keeps sending eviction attacks directly to the validator node for several consecutive blocks.
+When observing the block $BasePrice$ drop sufficiently, the attacker mounts the regular locking attack $XT_{8}$ against the Reth node.
+6.4.2 Attack Evaluation on Reth
+Experimental settings: We set up an experiment platform for evaluating locking attacks, on which an attacker node is connected to a victim non-validator node, which is also connected to a workload generator node and a victim validator node.
+The attacker node also maintains a direct connection to the victim validator.
+The network topology is depicted in Figure 4.
+The non-validator node runs the Geth $v1.11.4$ client.
+Figure 4: Experimental setup for locking attacks on Reth
+Figure 5: Evaluation of locking attack $XT_{8}$ on Reth
+Evaluation of $XT_{8}$: To evaluate $XT_{8}$, we set up the workload generator that sends normal transactions to the validator Geth node.
+The non-validator Reth node has not yet joined the network.
+We let the attacker node first send eviction attacks ($XT_{6}$) directly to the victim validator for $35$ consecutive blocks until it observes the block $BasePrice$ drops sufficiently to $1$ Gwei, which occurs at the $45$-th in our experiment.
+The Reth node then joins the network with an empty mempool.
+The attacker node mounts a locking attack $XT_{8}$ with transactions of price $5$ Gwei to occupy the Reth node’s mempool.
+We report the total fees of normal transactions included in the blocks – the lower the fees are, the more successful the locking attack is.
+We also report the adversarial transaction fees as the attack cost shown in Figure 5.
+When the locking attack begins at the $45$-th block, the normal transaction fees increase because of the empty mempool on the Reth node which accepts some normal transactions in addition to the adversarial locking transactions.
+Then, after three blocks, the normal transaction fees quickly drop to near-zero Ether, showing the success of locking attacks.
+Interestingly, unlike the eviction attacks that cause low Gas utilization per block, the locking attacks just reduce the total Ether fees without reducing Gas utilization (i.e., the blocks produced under locking attacks from height $45$ to $60$ use the Gas of almost $100\%$ block limit).
+Due to the high Gas utilization, block $BasePrice$ keeps increasing during the locking attack.
+When reaching the $60$-th block, the $BasePrice$ is higher than the maximal transaction price tolerable by the attacker.
+In our experiment, the locking attacker stops at $60$-th block, and the normal transaction fees immediately recover to the “normal” level.
+In practice, the attacker can repeat sending eviction attacks to reduce the $BasePrice$ before mounting the locking attack again.
+7 Evaluation of mpfuzz
+7.1 Performance
 Preliminary of GoFuzz: Upon each fuzzing iteration, GoFuzz [16] generates a bit-string of variable length, feeds it to our fuzzer code, and receives from our fuzzer a binary value of feedback, be either positive or negative.
-The GoFuzz further checks if this run increases the code coverage in the Geth client (as will be described, our fuzzer invokes functions in Geth’s txpool).
+The GoFuzz further checks if this run increases the code coverage.
 If the feedback is positive and code coverage increases, GoFuzz adds the current bit-string to the seeds before running into the next iteration.
@@
 It skips a bitstring that is too short.
-2) The fuzzer then sets up and initializes a Geth txpool instance with $m$ normal transactions, each sent from a distinct account, of nonce $1$, of value $1$ Ether (lower than the account balance), and of Gas price $3$ wei.
+2) The fuzzer sets up and initializes a Geth txpool instance with $m$ normal transactions, each sent from a distinct account, of nonce $1$, of value $1$ Ether (lower than the account balance), and of GasPrice $3$ wei.
 3) The fuzzer sends the transaction sequence to the initialized mempool for execution.
-4) It checks the mempool state after execution: If the eviction-based test oracle (Equation 1) is met, it emits the current transaction sequence as an exploit.
+4) It checks the mempool state after execution: If the test oracle is met, it emits the current transaction sequence as an exploit.
 Baseline B2: Concrete-state-coverage fuzzer.
-We also build the second baseline, a mempool fuzzer that takes concrete state coverage as feedback.
-Specifically, in each iteration, the fuzzer appends to the current transaction sequence a new transaction.
-Given the $m$-slot mempool, the fuzzer tries $m$ values for senders, $m$ values for nonces, $m$ values for Gas price, and $m$ values for Ether amount.
-After sending to the mempool the current transaction sequence, it sorts the transactions in the mempool by senders and nonces.
+We build the second baseline, a mempool fuzzer that takes concrete state coverage as feedback.
+Specifically, in each iteration, the fuzzer appends a new transaction to the current transaction sequence.
+Given the $m$-slot mempool, the fuzzer tries $m$ values for senders, nonces, GasPrice, and Ether amount.
+After sending the current transaction sequence to the mempool, it sorts the transactions in the mempool by senders and nonces.
 It then hashes the sorted transactions in the mempool and checks the presence of the hash digest in the seeds.
@@
 Baseline B3: Invalid-tx energy fuzzer.
-We build the last baseline that is similar to the concrete-state-based fuzzer except for the following: It uses as the energy the number of invalid transactions on a mempool state.
+Baseline B3 is similar to B2 except that it uses as the energy the number of invalid transactions on a mempool state.
 The seed with more invalid transactions in the mempool has higher energy or priority to be selected for the next round of fuzzing.
-We implement Baseline B1 in Go/GoFuzz and implement B2 and B3 in Python. We use a test oracle that can detect DETER attacks, that is, the number of invalid transactions in a mempool state equal $m-1$.
-We run the three baselines and mpfuzz against a Geth client under two settings: A small setting where the mempool is resized to $6$ slots $m=6$ and all fuzzers run for two hours (if test oracle is not triggered), and a medium setting where $m=16$ and fuzzers run for $16$ hours.
-Table 3: Fuzzing Geth $v1.11.3$’s mempool by different approaches to detect ADAMS exploit.
-Settings
-B1 B2 B3 mpfuzz
-6-slot-2h
-timeout $54$ (min) $8$ (min)
-$0.11$ (min)
-16-slot-16h
-timeout timeout $447$ (min)
-$0.84$ (min)
-Table 3 reports the results under the small and medium settings.
-For the small setting with $6$ slots, the stateless fuzzer (baseline B1) cannot find any DETER attack in two hours, concrete-state-coverage fuzzer (baseline B2) can find a DETER exploit in 54 minutes, and the stateful fuzzer based on invalid-transaction counter finds a DETER exploit in $8$ minutes.
-By contrast, mpfuzz finds a DETER exploit in $0.11$ minutes.
-For the medium setting, with $6$ slots, both baseline B1 and B2 cannot find any DETER exploit in $16$ hours, and baseline B3 can find a DETER exploit in $7.4$ hours.
-By contrast, mpfuzz finds a DETER exploit in under a minute.
+Baseline B4: feedback of no promising states: We design Baseline B4 for an ablation study on the effectiveness of using promising states as feedback in improving fuzzing performance.
+B4 is the same with mpfuzz except that it removes the state promising-ness from its feedback.
+Specifically, the feedback formula in B4 includes only state coverage ($st\_coverage$) and excludes state promisingness ($st\_promising$ as in Equation 7).
+We implement Baseline B1 in Go/GoFuzz and implement B2, B3 and B4 in Python.
+We use a test oracle that can detect DETER attacks, that is, the number of invalid transactions in a mempool state equal $m-1$.
+Experimental settings of fuzzing: We run the four baselines and mpfuzz against a MUT in a Geth client reconfigured under two settings: A small setting where the mempool is resized to $6$ slots $m=6$ and all fuzzers run for two hours (if test oracle is not triggered), and a medium setting where $m=16$ and fuzzers run for $16$ hours.
+The fuzzing experiments are run on a local machine with an Intel i7-7700k CPU of $4$ cores and $64$ GB RAM.
+Table 4: Fuzzing Geth $v1.11.3$’s mempool (in minutes) by different approaches to detect Exploit $XT_{3}$.
+Settings B1 B2 B3 B4 mpfuzz
+$6$slot-$2$h Timeout $54$ $8$ $1.22$ $0.03$
+$16$slot-$16$h Timeout Timeout $447$ Timeout $0.06$
+Results: Table 4 reports the results under the small and medium MUT settings.
+For the small MUT, B1 cannot find any DETER attack in two hours.
+B2, B3 and B4 can find $XT_{3}$ in 54 minutes, $8$ minutes and $1.22$ minutes respectively.
+By contrast, mpfuzz finds the same exploit in $0.03$ minutes.
+For the medium setting, with $16$ slots, baselines B1, B2, and B4 cannot find any DETER exploit in $16$ hours, and B3 can find Exploit $XT_{3}$ in $7.4$ hours.
+By contrast, mpfuzz finds the same exploit in under a minute.
+The evaluation on different clients is in Appendix C.
+7.2 True/False Positive Rates
+This section evaluates the true/false positives of the short exploits found by mpfuzz.
+Recall that mpfuzz runs against a small mempool (MUT) and may introduce false positives, that is, the exploits that satisfy bug oracles in MUT but do not in the actual larger mempool.
+We formally define true and false positives as follows:
+Definition 7.1 (TP/FP exploits)
+Consider a short exploit found by mpfuzz on a MUT, $\langle{}st_{0},dc_{0}\rangle{},ops\Rightarrow{}\langle{}st_{n},dc_{n}\rangle{}$, and a (manually) extended exploit on an actual mempool, $\langle{}st_{0}^{\prime},dc_{0}^{\prime},ops^{\prime}\Rightarrow{}\langle{}st_{n}^{\prime},dc_{n}^{\prime}\rangle{}$.
+The pair of exploits constitute a true positive eviction attack w.r.t.
+$\epsilon$ ($\lambda$), if.f.
+$st_{0}\cap{}st_{n}=\varnothing\land{}asym_{E}(st_{0},ops)<\epsilon\land{}\ st_{0}^{\prime}\cap{}st_{n}^{\prime}=\varnothing\land{}asym_{E}(st_{0}^{\prime},ops^{\prime})<\epsilon$
+The pair of exploits constitute a false positive eviction attack w.r.t.
+$\epsilon$ ($\lambda$), if.f.
+$st_{0}\cap{}st_{n}=\varnothing\land{}asym_{E}(st_{0},ops)<\epsilon\land{}\ (st_{0}^{\prime}\cap{}st_{n}^{\prime}\neq{}\varnothing\lor{}asym_{E}(st_{0}^{\prime},ops^{\prime})>\epsilon)$.
+True/false positives on locking attacks can be similarly defined from bug-oracle definitions 4.2 and 4.3.
+We evaluate the false positives of the exploits found by mpfuzz under varying $\epsilon$ and $\lambda$.
+Specifically, we design a two-step experiment: First, given a MUT, we run mpfuzz with varying $\epsilon$ (and $\lambda$).
+Increasing $\epsilon$ allows mpfuzz to find more short exploits; for each distinct short exploit, we manually extend it to an actual exploit.
+Second, we evaluate the extended exploits on an actual mempool (in the same local experimental setting as described in § 9).
+Table 5: True/false positives (TP/FP) of eviction and locking attacks discovered with varying $\epsilon$ and $\lambda$. ✓means satisfied.
+Exploit $\epsilon$ $m=16$ (MUT) Default $m$ TP?
+Eq. 1 $asym_{E}$ Eq. 1 $asym_{E}^{\prime}$
+Geth-$XT_{1}$ $10^{-4}$ 00 00
+Geth-$XT_{3}$09.09.09.09 $0.083$ $0.0002$
+Geth-$XT_{6}$125.125.125.125 $0.125$ $0.0003$
+Geth-$XT_{2}$ $0.2$ $0.167$ $0.0698$
+Geth-$XT_{4}$ $0.23$ $0.208$ $0.0768$
+Geth-$XT_{5}$ $0.23$ $0.208$ $0.0768$
+Nethermind-$XT_{1}$ $10^{-4}$ 00 00
+Nethermind-$XT_{4}$ $0.11$ $0.104$ $0.0008$
+Nethermind-$XT_{7}$ $0.36$ $0.355$ $0.0012$
+Besu-$XT_{2}$ $0.2$167.167.167.167 $0.0754$
+Besu-$XT_{4}$ $0.23$208.208.208.208 $0.0083$
+Erigon-$XT_{4}$ $0.23$ $0.208$ $0.0894$
+Exploit $\lambda$ $m=16$ (MUT) Default $m$ TP?
+Eq. 4 $asym_{D}$ Eq. 4 $asym_{D}^{\prime}$
+Reth-$XT_{8}$ $0.34$ $0.34$ $0.015$
+OpenEthereum-$XT_{9}$ $0.46$ $0.46$ $0.0439$
+Table 5 presents the details of exploits discovered by mpfuzz.
+For instance, when setting $epsilon$ at $0.0001$ to find eviction attacks on Geth, mpfuzz discovered one short exploit $XT_{1}$ with $asym_{E}=0$ on MUT, which can be extended to an actual exploit on Geth mempool (of $m=6144$) with $asym_{E}^{\prime}=0<0.0001=\epsilon$.
+This makes it a true positive (marked by ✓in the table).
+Increasing $\epsilon$ to $0.23$ leads to discovering all six exploits $XT_{1-6}$ on Geth.
+All these exploits are true positives.
+The table also includes the results of eviction exploits found on other clients and locking exploits.
+Overall, the true-positive rate of mpfuzz remains at $100\%$ for $\epsilon\leq{}0.36$ and for $\lambda\leq{}0.46$ across Ethereum clients.
+8 Discussions
+Root causes of the exploits: We attribute all found exploits $XT_{1-9}$ to four distinct causes: 1) the presence of certain admission patterns (including $XT_{1}/XT_{2}/XT_{3}/XT_{4}/XT_{6}/XT_{7}$).
+For instance, $XT_{4}$ allows transaction replacement to cause latent overdrafts in place of valid transactions.
+2) the absence of certain admission patterns to cause mempool locking (including $XT_{9}$ and $XT_{8}$).
+For instance, the cause of $XT_{8}$ on Reth is that Reth mempool is a FIFO queue, and it disallows transaction eviction of any kind, which is risky.
+3) Inconsistency across multiple admission patterns, notably $XT_{7}$.
+In $XT_{7}$, the cause is that a mempool allows an eviction and its reversed eviction at the same time.
+4) Evadable conditions to context-sensitive admission patterns, including $XT_{3}$ and $XT_{6}$.
+Both exploits work on Geth, where the mempool triggers the limit of transactions per send under an evadable condition (when the mempool stores more than $py_{3}^{\prime}=5120$ pending transactions).
+Mitigation: We propose schemes to mitigate all newly discovered exploits comprehensively.
+In principle, given an exploit under a known cause, our mitigation design is to negate the cause.
+For instance, to mitigate $XT_{4}$, we propose declining transaction replacements that cause latent overdrafts.
+We have implemented this mitigation strategy on Geth, which was merged into the release of Geth $v1.11.4$ [8] .
+Similarly, to mitigate $XT_{7}$, we propose ensuring the consistency between eviction events and reversed events (e.g., if admitting $tx_{1}$ by evicting $tx_{2}$ is allowed, then admitting $tx_{2}$ by evicting $tx_{1}$ is not allowed).
+$XT_{1}/XT_{2}$ are mitigated by disallowing eviction of valid transactions by future/latent-overdraft transactions.
+$XT_{3}/XT_{6}$ are mitigated by removing the triggering condition of transaction sender limit (i.e., making $py_{3}^{\prime}=0$).
+$XT_{8}$ can be mitigated by re-enabling transaction evictions in certain cases, such as price-based transaction eviction (i.e., a pending transaction arriving later and of higher price can evict another pending transaction of lower price and admitted earlier).
+If an exploit relies on multiple causes (e.g., $XT_{3}$ on Causes 1 and 4), negating one cause suffices to mitigate the exploit.
+There are more sophisticated cases where one exploit relies on the presence of an admission pattern, and another exploit relies on the absence of the pattern.
+For instance, $XT_{5}$ relies on the presence of an admission policy named $SAP_{5}$ that turns child transactions into future ones (as in Geth), and the success of $XT_{9}$ relies on the absence of the same policy, that is, $\neg{}SAP_{5}$ (as in OpenEthereum).
+To mitigate both $XT_{5}$ and $XT_{9}$, we propose a defense strategy in which the admission decision regarding $SAP_{5}$ is non-deterministic or randomized so that $SAP$ is not always on or off, making it hard for either $XT_{5}$ or $XT_{9}$ to always succeed.
+[…]
+Ethical concerns: When evaluating attacks, we only mounted attacks on the testnet and did so with minimal impacts on the tested network.
+For instance, our attack lasts a short period of time, say no more than $4$ blocks produced.
+We did not test our attack on the Ethereum mainnet.
+We also mask part of the block numbers in the screenshots (e.g., Figure 3) to prevent detailed attack inspection and reproduction.
+When reporting bugs, we disclose to the developers the mitigation design tradeoff and the risk of fixing one attack by enabling another attack (described above).
+To prevent introducing new bugs, we did not suggest fixes against turning-based locking ($XT_{5}$, $XT_{6}$) and locking ($XT_{8}$).
+9 Conclusion and Future Works
+Conclusion: This paper presents mpfuzz, the first mempool fuzzer to find asymmetric DoS bugs by exploring symbolized mempool states and optimistically estimating the promisingness of an intermediate state in reaching bug oracles.
+Running mpfuzz on popular Ethereum clients discovers new mempool-DoS vulnerabilities, which exhibit various sophisticated patterns, including stealthy mempool eviction and mempool locking.
+Limitations and future works: The exploit generation in this work is not fully automated.
+Manual tasks include mempool reduction, mpfuzz setup (configuring $\epsilon$ and symbols), exploit extension, etc.
+Automating these tasks is the future work.
+mpfuzz’s bug oracles neither captures complete dependencies among concrete transactions nor guarantees completeness in finding vulnerabilities.
+Certified mempool security with completeness is also an open problem.
+mpfuzz targets a victim mempool of limited size: the vulnerabilities found in this work are related to transaction eviction, which does not occur in a mempool of infinite capacity.
+Thus, the mpfuzz workflow cannot find exploits in the mempool of infinite capacity.
+It is an open problem whether a mempool of large or infinite capacity has DoS vulnerabilities.
+10 Acknowledgments
+The authors thank USENIX Security reviewers and Kangjie Lu for the discussion in the early stage of this work.
+Yibo Wang, Wanning Ding, Zhihua Yang, and Yuzhe Tang were partially supported by two Ethereum Foundation academic grants and NSF grants CNS-2139801, CNS-1815814, DGE-2104532.
+Kai Li is partially supported by NSF grant CNS-2347486 and one Ethereum academic grant.
+Appendix A Observations Motivating mpfuzz Design
+(In)feasibility of code-coverage-only stateless fuzzing: To fuzz a mempool, a baseline design is to use a code-coverage-based fuzzer.
+In each iteration, the fuzzer generates a bit string, parses it into a sequence of transactions and sends them to the tested mempool for execution.
+Upon each end state, it checks whether the test oracle defined above is satisfied.
+This design takes code coverage as the only feedback and is stateless because the feedback does not consider the end state.
+The stateless design would be ineffective in fuzzing mempool or finding ADAMS exploits.
+For example, consider fuzzing a three-slot mempool to find a DETER-X exploit consisting of three future transactions [28].
+Suppose the current mempool input is a sequence of one future transaction.
+After mutation, it may try an input of two future transactions, which, however, does not increase code coverage (the second future transaction is executed by the same code path in a mempool as the first future transaction).
+Thus, the fuzzer discards the two-future-transaction input and misses finding the three-future-transaction exploit.
+We validate the ineffectiveness of stateless fuzzer by experiments in § 7.
+A more promising design is stateful mempool fuzzing, in which the mempool state is included in the feedback to guide next-iteration fuzzing.
+In the previous example of finding a three-future-transaction exploit (DETER-X), sending two future transactions leads to a different mempool state than sending one.
+A simple state-coverage-based fuzzer would view this as positive feedback and further explore this direction toward finding the three-future-transaction exploit.
+(In)feasibility of existing consensus fuzzers: Existing consensus fuzzers [25, 29, 36] cannot detect mempool DoS or ADAMS bugs.
+Specifically, Loki’s test oracle is detecting system crashes, while a successful ADAMS attack does not necessarily trigger system crashes.
+Using Loki to detect ADAMS would cause many false negatives.
+Fluffy detects the unsynchronizable difference of post-consensus states as vulnerability.
+However, the mempools in two clients, say Geth and Besu under benign transactions, would be different.
+In other words, post-consensus state differences across clients could indicate insecurity, but the difference in pre-consensus states across clients, like mempool, does not mean insecurity.
+Using Fluffly to detect ADAMS could cause false positives.
+Tyr’s test oracle detects property violations on post-consensus states.
+Particularly, while Tyr models state liveness, their definition “valid transactions should be executed eventually” is, unfortunately, over-simplifying and fails to model the legitimate pre-consensus cases in real Ethereum clients; for instance, valid transactions of low price can be dropped by Ethereum mempools and will not be eventually executed.
+Besides, they model double-spending transactions as the only type of invalid transaction.
+In Ethereum, double-spending transactions are of the same nonce.
+And there are many more sophisticated invalid transactions in Ethereum that Tyr does not model, including future transactions, latent overdrafts, etc.
+We analyze the case of existing consensus fuzzers in detecting known DETER attacks [28].
+Loki cannot detect DETER because DETER does not cause system crashes.
+Fluffy does not model future transactions and can not find the DETER attacks that rely on future transactions.
+Similarly, Tyr that does not model sophisticated invalid transactions would miss detecting DETER attacks.
+A.1 Further Rationale of Symbolization
+We now describe the design rationale of our symbolization technique.
+  • The first observation is that real-world mempools admit transactions based on the symbolic value, not concrete value, of nonce and Ether amount.
+Our design reflects this intuition.
+For instance, symbols of valid transactions, e.g., $\mathcal{P},\mathcal{C},\mathcal{N}$, are instantiated by mpfuzz to a minimal amount $1$ Ether.
+As long as the transferred value does not exceed the sender balance, transactions remain valid (i.e., no overdraft).
+Admission decisions remain the same for two valid transactions despite their difference in nonces or values.
+    Likewise, Symbol $\mathcal{F}$ is instantiated by mpfuzz to fixed nonce $n+1$: As long as the nonce is non-consecutive, no matter what specific value it takes (e.g., $3$ or $m+1$), the mempool would deem it as a future transaction and makes the same admission decision.
+    As a side note, the nonce of Symbol $\mathcal{F}$ is fixed at a large nonce $n+1$ (we denote it by far future transactions) instead of a small nonce, say $5$ (we denote it by near future transactions), because near future transactions’ nonce could be reconnected as consecutive and lead to duplicated states.
+  • The second observation is that restricting eviction/replacement victims to only parent transactions can facilitate finding exploits quickly.
+Specifically, while any transactions can be evicted/replaced, mpfuzz evicts/replaces only parent transactions; such a strategy can cause maximal damage (i.e., the most child transactions turned) while saving the search space.
+Our symbol design reflects this idea.
+For instance, Symbol $\mathcal{R}$ covers any transaction of the same sender and nonce as a transaction in state $st$, but it is instantiated by mpfuzz to only the transaction of nonce $1$ (i.e., replacing a parent transaction).
+Appendix B Case Study: How mpfuzz Finds Exploits
+We describe how mpfuzz finds an exploit by presenting a case study on finding $XT_{6}$ in the latest Geth $\geq{}v1.11.4$.
+Mempool reduction: Recall that a Geth mempool has a capacity of $m^{\prime}=6144$ slots, and its transaction-admission policies are characterized by three essential parameters: admitting up to $py_{1}^{\prime}=1024$ future transactions and limiting up to $py_{2}^{\prime}=16$ pending transactions from any senders when more than $py_{3}^{\prime}=5120$ pending transactions are residing in the mempool.
+We set up the MUT to run the same codebase or the same admission policy but with smaller parameters: $m=3,py_{1}=1,py_{2}=2,py_{3}=2$ What follows is a description of how mpfuzz find a short exploit on this MUT.
+Fuzzing: How mpfuzz automatically finds exploits: Initially, the mempool is filled with $m=3$ normal transactions.
+That is, the initial symbolized state is $\mathcal{NNN}$.
+The seed corpus $sdb$ initially contains an empty string.
+mpfuzz retrieves the empty string and appends to it with different symbols $\mathcal{O}_{0}$, $\mathcal{C}_{0}$, $\mathcal{P}_{0}$.
+Because of the initial state $\mathcal{NNN}$, only Symbol $\mathcal{P}_{0}$ is feasible.
+It generates the mutated input $\mathcal{P}$ and instantiates it to a parent transaction of a higher GasPrice than normal transactions (as described in § 5).
+Sending the input to the mempool gets the transaction admitted, leading to transitioned state $st_{1}=\mathcal{NNP}$.
+This is a new state that is not in the corpus, and it evicts more normal transactions $\mathcal{N}$ than the previous state $st_{0}$; the input produces positive feedback, and the associated input-state pair $\langle{}P,st_{1}=\mathcal{NNP}\rangle{}$ is added to the corpus.
+Figure 6: Snapshot of the mpfuzz state-search tree when finding Exploit $XT_{6}$ on Geth $v1.11.4$.
+Next, mpfuzz retrieves from $sdb$ a seed by high energy.
+According to Table 2, the energy of state $\mathcal{NNN}$ is $\frac{1}{3*3}*0$, and the energy of state $\mathcal{NNP}$ is $\frac{1}{3*2+4}*1=1/10>0$.
+Hence, seed $\mathcal{NNP}$ is selected.
+mpfuzz tries input mutation and appends to the selected input $\mathcal{P}$ one of four new symbols, that is, $\mathcal{O}_{1}$, $\mathcal{C}_{1}$, $\mathcal{P}_{0}$ or $\mathcal{P}_{1}$.
+On state $\mathcal{NNP}$, 1) mutation transaction instantiated from symbol $\mathcal{O}_{1}$ is declined admission.
+2) Mutation transaction from $\mathcal{C}_{1}$ is admitted, transitioning the state to $\mathcal{NPC}$, which produces positive feedback and is added to $sdb$.
+3) Likewise, transaction $\mathcal{P}_{0}$ is admitted and produces state $\mathcal{NPP}$ of positive feedback; the mutated input is also added to $sdb$.
+4) Mutation $\mathcal{P}_{1}$ is admitted but produces an identical state with mutation $\mathcal{P}_{0}$; thus, the mutated input is not added to $sdb$.
+Now, there are four seeds in $sdb$: $\mathcal{NNN}(0)$, $\mathcal{NNP}(0)$, $\mathcal{NPC}(1/8)$, and $\mathcal{NPP}(1/12)$.
+In parentheses are their energy numbers.
+The seed of the highest energy $\mathcal{NPC}$ is selected ( 1 in Figure 6).
+mpfuzz then mutates $\mathcal{NPC}$ with $4$ possibilities, that is, $\mathcal{O}_{1}$, $\mathcal{C}_{1}$, $\mathcal{P}_{0}$, and $\mathcal{P}_{1}$, which produce two end states with positive-feedback, that is, $\mathcal{PPC}$ and $\mathcal{PCP}$.
+They are added to the $sdb$ with energy $\mathcal{PPC}(1/10)$ and $\mathcal{PCP}(1/10)$.
+Let’s say $\mathcal{PPC}(1/10)$ is selected ( 2).
+mpfuzz mutates input $\mathcal{P}\mathcal{C}_{1}\mathcal{P}_{0}$ with $7$ mutation transactions, that is, $\mathcal{O}_{1}$, $\mathcal{O}_{2}$, $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, $\mathcal{P}_{0}$, $\mathcal{P}_{1}$, $\mathcal{P}_{2}$, which produces one end state with positive feedback, that is, mutation $\mathcal{C}_{1}$ is admitted and transitions state to $\mathcal{FPC}$.
+After that, $\mathcal{FPC}$ is the one with the highest energy and is chosen for the next-round fuzzing ( 4).
+mpfuzz mutates input $\mathcal{P}\mathcal{C}_{1}\mathcal{P}_{0}\mathcal{C}_{1}$ with $4$ mutations and produces $2$ state transitions with positive feedback.
+Upon state $\mathcal{FPC}$, mutation $\mathcal{C}_{1}$ is admitted and leads to state $\mathcal{FEE}$, which satisfies the bug oracle of eviction attacks under $\epsilon=0$.
+The algorithm then emits the found short exploit: $\langle{}st_{0}=\mathcal{NNN},dc_{0}=\varnothing\rangle{},ops=\mathcal{P}\mathcal{C}_{1}\mathcal{P}_{0}\mathcal{C}_{1}\mathcal{C}_{1}$ (recall Definition 4.1).
+The snapshot of the state-search tree is depicted in Figure 6.
+Exploit extension: Given the short exploit automatically found on MUT (with $m=3,py_{1}=1,py_{2}=2,py_{3}=2$), the next step is to extend it to a longer exploit functional on the actual Geth mempool (with $m^{\prime}=6144,py_{1}^{\prime}=1024,py_{2}^{\prime}=16,py_{3}^{\prime}=5120$).
+Exploit extension requires manual efforts: After identifying exploit $\langle{}st_{0}=\mathcal{NNN},dc_{0}=\varnothing\rangle{},ops=\mathcal{P}\mathcal{C}_{1}\mathcal{P}_{0}\mathcal{C}_{1}\mathcal{C}_{1}$ is unique, we extend it to the longer exploit by ensuring the same admission event occurs on the actual mempool as on the smaller MUT.
+In this process, it tries the next transaction of the same sender with the previous one but with an incremented nonce.
+If it fails, it switches to the next sender.
+It also tries transaction fees/prices based on measuring the fees of actual normal transactions, which may not be fixed as $3$ in the mpfuzz setting.
+Appendix C Additional Eval. of mpfuzz Performance
 We further run our mpfuzz on different clients, including Geth $v1.10.11$, Geth $v1.11.4$, Erigon $v2.42.0$, and Nethermind $v1.18.0$.
@@
 We report the number of exploits found by mpfuzz in a $16$-hour period.
-Figure 3 presents the result that mpfuzz finds $4096$ exploits in the first $8$ hours on Geth-$v1.10.11$.
+Figure 8 presents the result that mpfuzz finds $4096$ short exploits in the first $8$ hours on Geth-$v1.10.11$.
 On Geth-$v1.11.4$, mpfuzz quickly finds one exploit (i.e., exploit $XT_{6}$ as described in § 6.1) within $2$ minutes and does not find anymore in the next $16$ hours.
@@
 On Erigon, it finds one exploit ($XT_{4}$) in the $16$-th hour.
-Figure 2: # exploits found
 Figure 7: # exploits found
@@
 In contrast, on Nethermind and Erigon, mpfuzz restarts the client in each iteration, which is consuming.
-7 Evaluation of Discovered Attacks
-This section shows the evaluation of turning attacks and locking attacks.
-7.1 Evaluation of Turning Attacks
-We design experiments to measure the success of ADAMS attacks in real Ethereum networks.
-7.1.1 Evaluation in Testnet
-Setup: We set up our experiment platform by connecting the attacker node to the Goerli testnet.
-The attacker node runs an instrumented Geth client at the execution layer which sends crafted transactions to the testnet.
-To monitor transaction propagation, we launch an independent supernode (called measurement node) in the same testnet.
-The measurement node runs a reconfigured Geth client where the limit of peers/neighbors is removed and the measurement node can connect to as many neighbors as possible, hence a super node.
-The measurement node is not (directly) connected to the attacker node.
-When setting up our experiments, we ran the measurement node in Goerli for seven days and found the node is stabilized at $290$ neighbors.
-(a) Transaction propagation in six messages
-(c) Etherscan screenshot of the blocks generated during the attack on Goerli (txs in six messages)
-Figure 4: Mounting $XT_{4}$ attacks on Goerli: Emptying blocks and propagating malicious transactions to the entire network.
-Experiment results: The instrumented measurement node is able to log the received messages from different neighbors; these messages include those of transactions and of transaction hashes.
-The measurement node could receive the same transaction from different neighbors, and the log stores the transaction-neighbor pairs.
-We started logging the received messages one month before the experiment.
-To do an experiment, we make the attacker node send $XT_{4}$ transactions using $384$ accounts.
-Each account sends $16$ valid pending transactions, followed by a replacement transaction.
-In total, the attacker node sends $6144$ valid transactions and $384$ replacement transactions.
-The Gas prices of the valid and replacement transactions are set to be $8$ Gwei and $130$ Gwei, respectively.
-In one setting, we split the $6528$ transactions into six messages (i.e., the first five being $1229$ transactions each and the last one with $384$ transactions).
-We send these messages in order and wait for $1.5$ seconds between two consecutive messages.
-For comparison, in another setting, we send the $6144$ valid transactions in one message and the $384$ replacement transactions in another message.
-Finally, we wrap up the experiment by waiting after the attacker node sends all messages and all replacement transactions are included in the blockchain.
-We then turn off the measurement node’s logging and search for the presence of all $6528$ transactions across neighbor nodes in the log.
-Given a transaction nonce, we report the percentage of neighbor nodes in the testnet that have propagated it to the measurement node.
-Table 4: Types of transactions included in the four blocks during the attack in Figure 4c. ‘# private tx’ means the number of private transactions.
-# block
-# private txs
-# public txs
-# attack txs
-$8040010$
-$5$
-00
-$384$
-$8040009$
-$78$
-$4$
-$384$
-$8040008$
-$1$
-00
-$384$
-$8040007$
-$5$
-00
-$384$
-We run experiments of six-message and two-message propagation, repeat each experiment by $3$ times, and report the average result of node percentage in Figure 4a and Figure 4b.
-Figure 4a shows the transaction propagation with six messages.
-More than 90% of the testnet nodes have received all of the $15$ valid transactions in a sequence and the replacement transactions.
-The only exception is the relatively low percentage for the $16$-th transaction of the maximal nonce.
-By comparison, when sending the transactions in two messages, the propagation fails as shown in Figure 4b where nodes did not propagate any transaction of nonce larger than $5$.
-Additionally, Figure 4c shows the generated blocks in the experiment propagating six messages.
-We took the screenshot from etherscan.io and labeled it in red with information regarding the attack.
-Before the attack begins, the Goerli testnet normally utilizes $24-47\%$ of the Gas in a block.
-For ethical consideration, we limit our attack to be short, specifically, four blocks long.
-Right after the launch of the attack, in block $8040007$, the Gas used by normal transactions drops to $3.19\%$ and the Gas used by adversarial transactions (denoted by red bars in Figure 4c) is $26.88\%$.
-The included adversarial transactions are $384$ replacement transactions sent in the second round of $XT_{4}$.
-The $6144-384=5760$ child transactions sent in the first round are not included in the block.
-Other blocks during the attack, namely $8040008$ and $8040010$, are similar.
-In the third block $8040009$, $73.1\%$ Gas is spent on including $82$ normal transactions as shown in Table 4 not sent by our attacker node.
-To explain it, we inspect the raw history of transactions arriving at and logged by our measurement supernode and found that out of the $82$ included normal transactions, only $4$ are present in the log, implying the other $78$ normal transactions are private ones that were not broadcast to the measurement node.
-Furthermore, the private mempool in today’s Ethereum clients [14] has a capability of no limit, which, while being feasible under today’s low private-transaction workloads, can cause resource-exhaustion vulnerability.
-Notice that we did not discover specific critical nodes to attack as needed by DETER attacks [22].
-8 Discussions
+Appendix D Additional Attack Evaluation
+D.1 Experiment on a Single Victim Node
+Figure 9: Experimental setup
+Evaluation settings: Our goal is to evaluate the success rate and cost of different ADAMS attacks on a single victim node.
+Because some ADAMS attacks are sensitive to the workload of normal transactions, we first collect transaction workloads from the mainnet.
+Specifically, we instrumented a Geth client (denoted by Geth-m) to log every message it receives from every neighbor.
+The logged messages contain transactions, transaction hashes (announcements), and blocks.
+When the client receives the same message from multiple neighbors, it logs it as multiple message-neighbor pairs.
+We also log the arrival time of a transaction or a block.
+Workload collection: We launched a Geth-m node in the mainnet on May 17, 2023, turned on logging for $5$ hours, and collected the transactions propagated to it.
+We make the collected transactions replayable as follows: We use the account balances and nonces on the mainnet to set up the initial state locally.
+We then replace the original sender in the collected transactions with the public keys that we generated.
+By this means, we know the secret keys of transaction senders and are able to send the otherwise same transactions for experiments.
+For experiments, we set up three nodes, an attack node sending crafted transactions, a workload node sending normal transactions collected, and a victim node receiving transactions from the other two nodes.
+The victim node is connected to both the attack and workload nodes.
+There is no direct connection between the attack node and the workload node.
+The attacker node runs an instrumented Geth $v1.11.4$ client that can propagate invalid transactions.
+The victim node runs the tested client.
+The workload node runs a vanilla Geth $v1.11.4$ client.
+On each node, we also run a Prysm $v3.3.0$ client at the consensus layer.
+The experiment platform is denoted in Figure 9.
+Among the three nodes, only the victim node is staked as a validator and would propose or produce blocks.
+In each experiment, we first run the above “attacked” setup.
+We then run a “regular” setup that excludes the attack node.
+Under the regular setup, the workload node sends the normal transactions and blocks to the victim node.
+We compare the experiment results under the attacked setup and regular setup to show the success of the attack.
+In the “attacked” setup, we collect the blocks produced and, given a block, we report two metrics: 1) total fees of benign transactions included, 2) total fees of attack transactions included.
+We also re-run the workload and victim nodes with the “regular” setup.
+We collect the blocks produced, and, given a block, we report two metrics: 3) total fees of transactions included, and 4) the cost of a baseline spamming attack with 100% success rate.
+For the latter, given a block, we select the transaction of the highest GasPrice and report the GasPrice multiplied by the Ethereum block Gas limit.
+Table 6: Attack success rate and cost (Ether/block)
+Clients Exploit Success rate Cost Baseline
+Geth $v1.10.25$ $XT_{1}$ $99.80\%$ 00 $11.39$
+$XT_{2}$ $99.42\%$ $0.725$ $11.39$
+$XT_{3}$ $93.02\%$ $0.0021$ $11.39$
+Geth $v1.11.4$ $XT_{4}$ $99.42\%$ $0.806$ $11.39$
+$XT_{5}$ $92.65\%$ $0.806$ $11.39$
+$XT_{6}$ $94.74\%$ $0.0022$ $11.39$
+Erigon $v2.42.0$ $XT_{4}$ $92.53\%$ $1.172$ $17.7$
+Nethermind $v1.18.0$ $XT_{4}$ $99.60\%$ $0.0021$ $10.75$
+$XT_{7}$ $84.63\%$ $0.20$ $10.75$
+Besu $v22.7.4$ $XT_{2}$ $99.63\%$ $1.04$ $17.7$
+$XT_{4}$ $99.60\%$ $1.06$ $17.7$
 […]
-Defenses: Rule-based mitigation can fix some of our discovered bugs: $XT_{4}$ is fixed by detecting latent overdraft transactions upon replacement.
-Locking-based $XT_{8}$ is fixed by re-enabling eviction and replacement; to avoid reintroducing eviction vulnerabilities, the fix needs to comply with the security rules against eviction attacks.
-Ethical concerns: When evaluating attacks, we only mounted attacks on the testnet and did so with minimal impacts on the tested network.
-For instance, our attack lasts a short period of time, say no more than $4$ blocks produced.
-We did not test our attack on the Ethereum mainnet.
-When reporting bugs, we disclose to the developers the mitigation design tradeoff and the risk of fixing one attacks by enabling other attacks (described above).
-To prevent introducing new bugs, we did not suggest fixes against locking ($XT_{8}$).
-Private and public mempools: In Ethereum, transactions can be sent privately or publicly.
-A private transaction received on a block builder is directly packed to a block without broadcast to other builders.
-By contrast, a public transaction received on a node is broadcast to other nodes before it is included in the next block.
-On the one hand, whether a mempool stores private or public transactions is orthogonal to the admission/eviction policy the mempool runs.
-Empirically, we found three open-source Ethereum block builders, named EigenPhi builder, Flashbot builder $v1.11.5-0.2.1$ and bloXroute builder.
-Specifically, Flashbot and bloXrout builder run the same admission policy as Geth $v1.11.4$, while EigenPhi builder runs the same admission policy as Geth $<v1.11.4$.Therefore, these private mempools are subject to the same vulnerabilities as the public mempools.
-On the other hand, a private mempool receives transactions from a portion of the web3 users, while a public mempool receives transactions from all web3 users.
-Thus, a disabled private mempool creates limited damage by victimizing the portion of users connected to the pool.
-The damage of a disable public mempool can be amplified by broadcasting the payload transactions across the network and reaching the public mempools of all nodes.
-9 Conclusion
-This paper presents mpfuzz, the first mempool fuzzer to find asymmetric DoS bugs by symbolically exploring mempool states and optimistically estimating the promisingness an intermediate state is in reaching bug oracles.
-Running mpfuzz on six major Ethereum clients leads to the discovery of new mempool vulnerabilities, which exhibit a wide variety of sophisticated patterns including stealthy mempool eviction and mempool locking.
+Flashbot builder $v1.11.5$ $XT_{6}$ $94.74\%$ $0.0022$ $11.39$
+EigenPhi $XT_{1}$ $99.80\%$ 00 $11.39$
+builder $XT_{2}$ $99.42\%$ $0.725$ $11.39$
+$XT_{3}$ $93.02\%$ $0.0021$ $11.39$
+$XT_{4}$ $99.42\%$ $0.806$ $11.39$
+$XT_{6}$ $94.74\%$ $0.0022$ $11.39$
+bloXroute builder-ws $XT_{6}$ $94.74\%$ $0.0022$ $11.39$
+go-opera $v1.1.3$ $XT_{2}$ $99.12\%$ $0.201$ $11.39$
+$XT_{3}$ $92.60\%$ $0.0021$ $11.39$
+$XT_{4}$ $99.13\%$ $0.221$ $11.39$
+$XT_{6}$ $93.76\%$ $0.0022$ $11.39$
+BSC $v1.3.8$ $XT_{6}$ $94.74\%$ $0.0022$ $11.39$
+core-geth $v1.12.18$ $XT_{6}$ $94.74\%$ $0.0022$ $11.39$
+[…]
+OpenEthereum $v3.3.5$ $XT_{4}$ $99.56\%$ $0.233$ $11.39$
+$XT_{9}$ $98.36\%$ $0.177$ $18.35$
+The evaluation results of all attacks on all clients are in Table 6.
+On the six Ethereum clients in the public transaction path, the attack success rates are all higher than $84.63\%$, and the attack costs are all lower than $1.172$ Ether per block, which is significantly lower than the baseline.
+We also evaluate the attacks found on PBS and Ethereum-like clients.
+Under the same experimental settings as § D.1, the results show similar success rates and costs with the attacks on Geth (recall these clients are Geth forks).
+More specifically, Table 6 shows that the success rates are higher than $92.60\%$, and attack costs are lower than $0.806$ Ether per block.

### v4

@@
-Understanding Ethereum Mempool Security under Asymmetric DoS by Symbolized Stateful Fuzzing
+Towards Automated Discovery of Asymmetric Mempool DoS in Blockchains
 Abstract
@@
 As shown in recent studies [1], denying a mempool service can force the blockchain to produce blocks of low or even zero (Gas) utilization, undermining validators’ incentives and shrinking the blockchain networks in the long run, re-introducing the $51\%$ attacks.
-Besides, a denied mempool service can prevent normal transactions from block inclusion, cutting millions of web3 users off the blockchain and failing the DApps relying on real-time blockchain access.
+Besides, a denied mempool service can prevent normal transactions from block inclusion, cutting millions of web3 users off the blockchain and failing the decentralized applications relying on real-time blockchain access.
 Problem: Spamming the mempool to deny its service has been studied for long [2, 3, 4, 5].
@@
 Existing blockchain fuzzers including Fluffy [7], Loki [8] and Tyr [9] all focus on fuzzing consensus implementation and don’t explore the extra transaction space required by mempool fuzzing.
-As a result, directly re-purposing a consensus fuzzer to fuzz mempool would be unable to detect the DETER bugs as evaluated in Appendix § A, let alone discover more sophisticated new ADAMS bugs.
+As a result, directly re-purposing a consensus fuzzer to fuzz mempool would be unable to detect the DETER bugs, let alone discover more sophisticated new ADAMS bugs.
 Proposed methods: To efficiently fuzz mempools, our key observation is that real-world mempool implementation admits transactions based on abstract “symbols”, such as pending and future transactions, instead of concrete value.
@@
 Second, mpfuzz is iteratively run against the reduced mempool to discover short exploits.
-Third, the short exploits are manually extended to actual ones that are functional on real mempools and Ethereum clients.
+Third, mpfuzz automatically extends the short exploits to actual ones that are functional on real mempools and Ethereum clients.
 Internally, the design of mpfuzz is based on seven transaction symbols we design out of Ethereum semantics: future, parent, overdraft, latent overdraft, and replacement transactions.
@@
 In particular, mpfuzz employs a novel technique to evaluate state promisingness, that is, by optimistically estimating the costs of unconfirmed transactions whose validity is subject to change in the future.
+After finding a short exploit against the reduced mempool, mpfuzz automatically extends the small exploit to a more extensive exploit that is effective within a larger mempool.
+Specifically, mpfuzz extracts all the unique admission event patterns in the small exploit and aligns these patterns to construct the attack transaction sequence on the larger-size mempool.
 Found attacks: We run mpfuzz on six leading execution-layer Ethereum clients deployed on the mainnet’s public-transaction path (Geth [10], Nethermind [11], Erigon [12], Besu [13], Reth [14], and OpenEthereum [15]), three PBS builders (proposer-builder separation) on the mainnet’s private-transaction and bundle path (Flashbot $v1.11.5$ [16], EigenPhi [17] and bloXroute [18]), and three Ethereum-like clients (BSC $v1.3.8$ [19] deployed on Binance Smart Chain, go-opera $v1.1.3$ [20] on Fantom, and core-geth $v1.12.19$ [21] on Ethereum Classic).
-Compared to the baseline fuzzers, mpfuzz can find exploits faster by more than two orders of magnitude.
 On Ethereum clients of historical versions, mpfuzz can rediscover known DETER bugs.
@@
 Example strategies include composing transactions of multiple patterns in one attack, re-sending evicted transactions to evict “reversible” mempools, locking mempool patched against turning, etc.
-We propose rule-based mitigation schemes against all newly discovered vulnerabilities.
 All newly found ADAMS bugs are reported to the developer communities, including Ethereum Foundation, BSC, Fantom, PBS builders, etc.
@@
 mpfuzz will be open-sourced to facilitate vulnerability discovery in more and future clients.
+Systematic evaluation: We conducted an ablation study to evaluate the performance of mpfuzz, which incorporates three distinct techniques.
+We aim to identify the contribution and significance of each component to the overall mpfuzz.
+We implement four baseline fuzzers to compare against mpfuzz.
+The first baseline is a stateless fuzzer built within the GoFuzz framework [24].
+For the remaining three baselines, each one selectively disables a single technique from mpfuzz: symbolized state, symbolized input, and promising feedback and energy.
+The experimental result shows that the stateless fuzzer is the least effective in discovering ADAMS exploits.
+Among the three distinct techniques, the symbolized input technique improves efficiency more compared to that of the promising feedback and symbolized state coverage, and the state promisingness in feedback is more beneficial for performance than symbolized state coverage.
+We systematically evaluate the attack success rate and cost of all the exploits found by mpfuzz on the Ethereum testnet and a local network.
+Specifically, we evaluate a newly discovered eviction attack on Goerli testnet.
+The result of the eviction attack shows the Gas used by normal transactions significantly drops to $3.19\%$ after the launch of the attack, which indicates that only a few normal transactions are included in the blocks following the attack.
+The locking attack evaluated on Goerli testnet shows the normal transactions received from the victim node drop essentially to zero, which indicates the locking attack is highly successful.
+Additionally, we evaluate other exploits on a local network.
+The evaluation results of all attacks on the Ethereum clients, PBS and Ethereum-like clients show that the attack success rates are all higher than $84.63\%$, and the attack costs are all lower than $1.172$ Ether per block.
 Contributions: The paper makes contributions as follows:
@@
 $\bullet$ New discovery of mempool vulnerabilities: mpfuzz discovers new asymmetric-DoS vulnerabilities in six major Ethereum clients in the mainnet.
+By evaluation under real transaction workloads, all found attacks achieve $84.6-99.6\%$ success rates and low costs such as adversarial transaction fees 100×100$\times$100 × lower than the victim transaction fees.
+$\bullet$ New discovery of mempool vulnerabilities: mpfuzz discovers new asymmetric-DoS vulnerabilities in six major Ethereum clients, PBS and Ethereum-like clients.
+We found and reported $24$ newly discovered ADAMS bugs.
+After our reporting, $15$ bugs are confirmed and $3$ bugs are fixed.
+$\bullet$ New found attack evaluation: This paper presents the evaluation results of the attack success rate and cost of all the exploits found by mpfuzz on the Ethereum testnet and a local network.
 By evaluation under real transaction workloads, all found attacks achieve $84.6-99.6\%$ success rates and low costs such as adversarial transaction fees 100×100$\times$100 × lower than the victim transaction fees.
@@
 2) The fuzzer node uses a learned model of blockchain network protocol to generate the next message of complying format, in which the content is mutated by randomly choosing integers and bit-flipping strings.
-That is, the mutation in Loki is unaware of application level semantics.
+In other words, the mutation in Loki is unaware of application level semantics.
 For instance, if the message is about propagating an Ethereum transaction, the transaction’s nonce, Ether amount, and Gas price are all randomly chosen; the sender, receiver, and data are bit flipped.
@@
 In practice, example policies include those favoring admitting transactions of higher prices (and thus evicting or declining the ones of lower prices), transactions arriving earlier, certain transaction types (e.g., parent over child transactions), and valid transactions (over future transactions).
-See different types of Ethereum transactions in § 4.1.
+See different types of Ethereum transactions in § 4-A.
 Transactions and fees: In Ethereum, Gas measures the amount of computations caused by smart contract execution.
@@
 The attacker in this work has the same networking capacity as in DETER [1].
-The attacker’s goal is to disable the mempools on the critical or all nodes in an Etheurem network and cause damage such as dropped transactions in block inclusion, produced blocks of low or zero (Gas) utilization, and disabled other downstream services.
+The attacker’s goal is to disable the mempools on the critical nodes or all nodes in an Etheurem network and cause damage such as deployed or dropped transactions in block inclusion, produced blocks of low or zero (Gas) utilization, and disabled other downstream services like MEV searchers.
 In practice, this damage is beneficial to competing block validators, MEV searchers, and RPC services.
 Besides, the attacker aims to cause this damage at asymmetrically low costs, as will be described.
-4.1 Notations
+4-A Notations
 Transactions: In Ethereum, we consider two types of accounts: a benign account of index $i$ denoted by $B_{i}$, and an adversarial account of index $i$ denoted by $A_{i}$.
@@
 This work does not consider the “data” field in Ethereum transactions or smart contracts because smart contracts do not affect the validity of the Ethereum transaction.
-The notations are summarized in Table 1.
-Table 1: Notations
+The notations are summarized in Table I.
+TABLE I: Notations
 Symbol Meaning Symbol Meaning
@@
 Green means automated tasks, and gray requires manual effort.
-4.2 Exploit Discovery Workflow
+4-B Exploit Discovery Workflow
 To set up the stage, we present the workflow overview in this work.
@@
 To discover vulnerabilities and exploits in an actual mempool, one has to 0) set up mpfuzz by various parameters (as will be introduced), 1) reduce the mempool from its original configuration to a much smaller version, named mempool under test (MUT).
-2) mpfuzz is run on the reduced MUT deployed locally (with settings described in § 7.1).
+2) mpfuzz is run on the reduced MUT deployed locally (with settings described in § 8-A).
 Fuzzing leads to the discovery of short exploits.
 3) Short exploits are extended to the longer ones, that is, actual exploits.
-4) The success of the actual exploit is evaluated on the actual mempools deployed in close-to-operational settings, such as using popular Ethereum testnets like Goerli (see § 6.3.1) or a network we set up locally running unmodified Ethereum clients (see § 4 and § D).
+4) The success of the actual exploit is evaluated on the actual mempools deployed in close-to-operational settings, such as using popular Ethereum testnets like Goerli (see § 7-A1) or a network we set up locally running unmodified Ethereum clients (see § 7-B2, § 7-C and § 7-D ).
 Fuzzing a reduced mempool (MUT) instead of an original one is necessary to ensure the efficient execution of the mpfuzz, which entails many iterations of re-initiating and populating the mempool.
-We discuss and evaluate the possible false positives that could be introduced by fuzzing smaller MUT in § 7.2.
-In this workflow, fully automated is Step 2), that is, the discovery of short exploits on a given MUT (colored green in Figure 1).
+In this workflow, fully automated is Step 2) and Step 3), that is, the discovery of short exploits on a given MUT and extension of the short exploit to an actual exploit (colored green in Figure 1).
 Other steps colored in gray require manual efforts and are described below: In Step 0), mpfuzz setup entails setting parameters in bug oracles (e.g., $\epsilon$ and $\lambda$) and fixed symbols in guiding fuzzing.
@@
 The mempool reduction requires expert understanding and manual efforts, and it is by nature the best effort.
-In Step 3), it first de-duplicates the raw short exploits found by mpfuzz.
-Then, for each distinct short exploit, it extends it by repeating the transaction-admission events to fit into the actual mempool, during which one may practice strategies like increasing nonce, switching sender accounts, and measuring normal transactions in the actual mempool.
-In exploit extension, one could also consider the factors not modeled in MUT or mpfuzz such as $BasePrice$, as seen in the case in § 6.4.
-4.3 Bug Oracle: Definitions & Rationale
+4-C Bug Oracle: Definitions & Rationale
 Definition 4.1 (Tx admission timeline)
@@
 The transaction admission timeline is simplified from the timeline that an operational mempool experiences where transaction arrival and block building can be interleaved.
-The simplification is intended to trade off the accuracy of fuzzing results for fuzzing efficiency as discussed in § 7.2.
 We use two mempool-attack templates to define two bug oracles: an eviction-based DoS where adversarial transactions evict existing victim transactions in the mempool, and a locking-based DoS where existing adversarial transactions in the mempool decline arriving victim transactions.
@@
 Besides, our bug oracle does not model the effect of EIP1559 or $BasePrice$.
-Without modeling $BasePrice$, our bug oracle can still capture the alternative transaction fee (the product of GasPrice and $21000$ Gas), which is still proportional to actual $PriorityPrice$ or damage on validator revenue.
+Without modeling $BasePrice$, our bug oracle can still capture the alternative transaction fee (the product of GasPrice and $21000$ Gas), which is still proportional to actual $tip$ or damage on validator revenue.
 5 Stateful Mempool Fuzzing by mpfuzz
-5.1 Transaction Symbolization
+5-A Transaction Symbolization
 The key challenge in designing mpfuzz is the large input space of transaction sequences.
@@
 In this work, we manually design seven symbols to represent transactions based on the Ethereum “semantics”, that is, how different transactions are admitted by the mempool.
-Table 2: Symbols, transactions, and associated costs. Tx refers to the instantiated transaction under a given symbol.
+TABLE II: Symbols, transactions, and associated costs. Tx refers to the instantiated transaction under a given symbol.
 Symbol Description Tx $cost()$ $opcost()$
@@
 Symbol $\mathcal{N}$ defines a transaction subspace covering any normal transactions sent from any benign account and of any nonce, namely $tx\begin{bmatrix}{B*}&{*}\\ {*}&{*}\end{bmatrix}$ In mpfuzz, Symbol $\mathcal{N}$ is instantiated to concrete transactions of fixed GasPrice $3$ wei and of value $1$ Ether.
-That is, symbol $\mathcal{N}$ is instantiated to transaction $tx\begin{bmatrix}{B*}&{1}\\ {*}&{3}\end{bmatrix}$ , as shown in Table 2.
+That is, symbol $\mathcal{N}$ is instantiated to transaction $tx\begin{bmatrix}{B*}&{1}\\ {*}&{3}\end{bmatrix}$ , as shown in Table II.
 Among all symbols, $\mathcal{N}$ is the only symbol associated with benign sender accounts.
@@
 mpfuzz instantiates $\mathcal{C}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{1}\\ {\geq{}2}&{m+4}\end{bmatrix}$}}$.
-The instantiated child transaction is fixed at GasPrice of $m+4$
+The instantiated child transaction is fixed at GasPrice of $m+4$.
 Symbol $\mathcal{O}$ defines the transaction subspace of any adversarial overdraft transaction w.r.t.
@@
 For instance, given state $st$ of transactions sent from account $A_{1},\dots,A_{r}$, transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ is within the defining space of symbol $\mathcal{F}$, but mpfuzz does not instantiate $\mathcal{F}$ by transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ (but instead to transactions $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+4}\end{bmatrix}$}}$).
-The design rationale of transaction symbolization is deferred to Appendix A.1.
-5.1.1 State Search Algorithm
+5-A1 State Search Algorithm
 Algorithm 1 mpfuzz(SeedCorpus $sdb$, Mempool $mp$)
@@
 5:         if is_ADAMS($st^{\prime}$then $\triangleright$ Test oracle
-6:             emit($ops^{\prime}$, $st^{\prime}$, "Found an exploit");
+6:             emit($ops^{\prime}$, $st^{\prime}$, ”Found an exploit”);
 7:         else
@@
 The algorithm runs an inner loop, in each iteration of which it mutates the input in the current seed ($\langle{}st,in\rangle{}$) and executes the mutated input against a reinitialized empty mempool, producing end state $st^{\prime}$ (Line 4).
-If the end state satisfies ADAMS conditions (i.e., the test oracle), it emits the mutated input-state pair $ops^{\prime},st^{\prime}$ as a newly found exploit.
+If the end state satisfies ADAMS conditions, it emits the mutated input-state pair $ops^{\prime},st^{\prime}$ as a newly found exploit.
 The algorithm further checks the feedback: The feedback is positive if the mutated input $ops^{\prime}$ brings the mempool state $st^{\prime}$ closer to triggering the test oracle than $st$.
@@
 We thus attribute, optimistically, the cost of transaction $\mathcal{C}$ to value $1$, so that it is preferable to a parent transaction $\mathcal{P}$ in input mutation and seed selection, and turning $\mathcal{C}$ into an invalid one like $\mathcal{L}$ or $\mathcal{F}$ is also encouraged.
-The cost profiles used in mpfuzz are summarized in Table 2.
+The cost profiles used in mpfuzz are summarized in Table II.
+5-B Exploit Extension
+The initial short exploit identified on MUT necessitates further automatic development to create a more extensive exploit that is effective within a larger mempool.
+The key idea in extending the small exploit to a more comprehensive one is to ensure the same admission event observed in the smaller MUT.
+We first define the admission event as given an arriving transaction $tx_{i}$ at a mempool of initial state $st_{0}$, the admission of $tx_{i}$ transitions the mempool into an end state $tx_{n}$.
+The admission event is denoted by $f(st_{0},tx_{i})\Rightarrow{}st_{n}$.
+We extract the admission event patterns, representing a state transition from the initial state to the end state.
+In the pattern, we only symbolize the mempool slots that are updated during the state transition.
+A pattern is denoted as $\langle{}st_{0}^{\prime},tx_{i},st_{n}^{\prime}\rangle{}$.
+For example, in the state transition from $\mathcal{NNPC}$ to $\mathcal{NPPC}$ by admitting a transaction $\mathcal{P}_{0}$, the symbolized state of a mempool slots transitions from $\mathcal{N}$ to $\mathcal{P}$.
+Thus, we extract the pattern as $\langle{}\mathcal{N},\mathcal{P},\mathcal{P}\rangle{}$.
+we extract all the unique admission event patterns in the small exploit and store them into an admission event pattern set ($aps$).
+We then construct the attack transaction sequence on the larger-size mempool by aligning the patterns identified from the small exploit with corresponding admission events.
+This entails validating that the admission event pattern of a given event $f(st_{0},tx_{i})\Rightarrow{}st_{n}$ on the larger-size mempool exists within the $aps$.
+Specifically, mpfuzz first selects $ap_{i}$ from the set $aps$ whose initial state $st_{0}$ matches the state of the large mempool.
+We then instantiate $tx_{i}$ of the $ap_{i}$ and send it to the large mempool.
+If the state transition of the large mempool $\langle{}st_{0}^{\prime},tx_{i},st_{n}^{\prime}\rangle{}$ aligns with the pattern $ap_{i}$, this positive outcome prompts mpfuzz to proceed with the extension process accordingly.
+Otherwise, mpfuzz backtracks by rolling back the mempool state to its state prior to the admission of $tx_{i}$ and tries the next pattern.
+mpfuzz tries the patterns in the decreasing order of $opcost()$, which means if multiple admission event patterns in $aps$ yield positive results, mpfuzz proceeds the extension process by applying the pattern with the lower $opcost()$.
 6 Found Exploits
-Table 3: ADAMS exploit patterns found by mpfuzz across Ethereum clients; $XT_{1-7}$ are eviction based, and $XT_{8-9}$ are locking based.
+TABLE III: ADAMS exploit patterns found by mpfuzz across Ethereum clients; $XT_{1-7}$ are eviction based, and $XT_{8-9}$ are locking based.
 $XT_{1-3}$ are known patterns in DETER [1], while others are new.
@@
 We have run mpfuzz across a variety of Ethereum clients, including six leading execution-layer clients on the public-transaction path of Ethereum mainnet (Geth, Besu, Nethermind, Erigon, Reth, and OpenEthereum), three PBS clients (proposer-builder separation) on the mainnet’s private-transaction and bundle path (Flashbot builder $v1.11.5$ [16], EigenPhi builder [17] and bloXroute builder-ws [18]), and the clients deployed on three operational Ethereum-like networks (BSC $v1.3.8$ [19] deployed on Binance Smart Chain, go-opera $v1.1.3$ [20] on Fantom, and core-geth $v1.12.19$ [21] on Ethereum Classic).
-On the six public-transaction clients, mpfuzz leads to the discovery of $22$ bugs, as listed in Table 3, including $7$ conforming to the known DETER attacks on clients of historical versions and $15$ new bugs on the clients of the latest versions.
-Other clients, including the three PBS clients and three clients on Ethereum-like networks, are mostly forks of the Geth clients except for minor code changes [^4] For instance, go-opera adopts its own fix against $XT_{1}$ based on the fork of Geth $<v1.11.4$., and on them, mpfuzz discovered $13$ bugs of a similar nature to those found on Geth (of the historical and latest versions), as seen in Table 3.
+On the six public-transaction clients, mpfuzz leads to the discovery of $22$ bugs, as listed in Table III, including $7$ conforming to the known DETER attacks on clients of historical versions and $15$ new bugs on the clients of the latest versions.
+Other clients, including the three PBS clients and three clients on Ethereum-like networks, are mostly forks of the Geth clients, and on them, mpfuzz discovered $13$ bugs of a similar nature to those found on Geth (of the historical and latest versions), as seen in Table III.
 We describe the patterns of these bugs as follows.
-6.1 Found Eviction Attacks
-Exploit $XT_{1}$: Direct eviction by future transactions: In this attack, given a mempool’s initial state storing normal transactions, the attacker sends future transactions at high GasPrice to evict the normal transactions.
+6-A Found Eviction Attacks
+Exploit $XT_{1}$: Direct eviction by future transactions: In this attack, given a mempool’s initial state storing normal transactions, the attacker sends future transactions at high fees to evict the normal transactions.
 This exploit is essentially the DETER-X attack [1].
-In practice, Geth ($\leq{}v1.11.4$), Nethermind, Besu, and EigenPhi are vulnerable under this exploit, as in Table 3.
-Exploit $XT_{2}$: Direct eviction by latent overdraft transactions: In this eviction attack, the attacker sends latent-overdraft transactions at high GasPrice to evict the normal transactions initially stored in the target mempool.
+In practice, Geth ($\leq{}v1.11.4$), Nethermind, Besu, and EigenPhi are vulnerable under this exploit, as in Table III.
+Exploit $XT_{2}$: Direct eviction by latent overdraft transactions: In this eviction attack, the attacker sends latent-overdraft transactions at high fees to evict the normal transactions initially stored in the target mempool.
 These transactions are sent from $k$ accounts, each of which sends $l$ transactions, denoted by $k\times{}l$.
@@
 Other clients including EigenPhi and go-opera are also vulnerable.
-Exploit $XT_{4}$: Indirect eviction by valid-turned-overdraft transactions: This eviction attack works in two steps: 1) The attacker first sends valid transactions at high GasPrice to evict normal transactions initially stored in the mempool.
+Exploit $XT_{4}$: Indirect eviction by valid-turned-overdraft transactions: This eviction attack works in two steps: 1) The attacker first sends valid transactions at high fees to evict normal transactions initially stored in the mempool.
 These transactions are sent from $k$ accounts, each of which sends $l$ transactions.
-2) She then sends $k$ transactions; each of the transactions is of nonce $1$, at a high GasPrice, of high value $v_{1}$, and from the same $k$ accounts in Step 1).
+2) She then sends $k$ transactions; each of the transactions is of nonce $1$, at a high fee, of high value $v_{1}$, and from the same $k$ accounts in Step 1).
 The transactions would replace the transaction of the same sender and nonce sent in Step 1).
@@
 When containing more than $py_{3}^{\prime}=5120$ pending transactions, the mempool starts to limit that no more than $py_{2}^{\prime}=16$ pending transactions from the same sender can be admitted.
-Exploit $XT_{6}$ works in three steps: 1) it first evicts the initial mempool of normal transactions with $m^{\prime}/py_{2}^{\prime}=384$ sequences, each of $16$ transactions sent from one unique sender, 2) it then sends $py_{1}^{\prime}/(py_{2}^{\prime}-1)=65$ transactions to evict the parent transactions and creates $py_{1}^{\prime}=1024$ future transactions in the mempool, 3) it sends one sequence of $5120$ transactions from one sender to evict the normal transactions sent in Step 1), and 4) it sends one transaction to evict the parent transaction in Step 3, leaving the mempool of just one valid transaction.
+Exploit $XT_{6}$ works in three steps: 1) it first evicts the initial mempool of normal transactions with $m^{\prime}/py_{2}^{\prime}=384$ sequences, each of $16$ transactions sent from one unique sender, 2) it then sends $py_{1}^{\prime}/py_{2}^{\prime}=64$ transactions to evict the parent transactions and creates $py_{1}^{\prime}=1024$ future transactions in the mempool, 3) it sends one sequence of $5120$ transactions from one sender to evict the normal transactions sent in Step 1), and 4) it sends one transaction to evict the parent transaction in Step 3, leaving the mempool of just one valid transaction.
 Step 3) can succeed because the precondition (w.r.t.
@@
 A mempool is reversible if one sends $tx^{\prime}$ to state $st^{\prime}$, and the mempool admits $tx^{\prime}$ by evicting $tx$, transitioning its state back to $st$.
-The mempool on Nethermind $<v1.18.0$ can be reversible: Minimally, suppose a two-slot mempool stores $tx_{1}$ of low GasPrice from one sender and $tx_{2}$ of medium GasPrice from another sender $A2$ and receives $tx_{3}$ as a child of $tx_{1}$ and of high GasPrice.
+The mempool on Nethermind $<v1.18.0$ can be reversible: Minimally, suppose a two-slot mempool stores $tx_{1}$ of low fee from one sender and $tx_{2}$ of medium fee from another sender $A2$ and receives $tx_{3}$ as a child of $tx_{1}$ and of high fee.
 For instance, $tx_{1}\begin{bmatrix}{A}&{1}\\ {1}&{1}\end{bmatrix}$, $tx_{2}\begin{bmatrix}{B}&{1}\\ {1}&{3}\end{bmatrix}$, $tx_{3}\begin{bmatrix}{A}&{1}\\ {2}&{5}\end{bmatrix}$.
@@
 each of $l$ transactions (i.e., $k\times{}l$).
-In each $l$-transaction sequence, the nonce-$1$ transaction has a low GasPrice, say $f_{1}$, while all its child transactions are of high GasPrice, evicting normal transactions initially stored in the mempool.
-2) The attacker sends $k\times(l-1)$ transactions from $k\times(l-1)$ new senders, all with GasPrice slightly higher than $f_{1}$.
+In each $l$-transaction sequence, the nonce-$1$ transaction has a low fee, say $f_{1}$, while all its child transactions are of high fees, evicting normal transactions initially stored in the mempool.
+2) The attacker sends $k\times(l-1)$ transactions from $k\times(l-1)$ new senders, all with fees slightly higher than $f_{1}$.
 The mempool would admit these transactions to the mempool by evicting all child transactions sent in Step 1), bringing down the attack costs.
-6.2 Found Locking Attacks
+6-B Found Locking Attacks
 […]
@@
 Exploit $XT_{9}$: Locking mempool of no turning: Because of the risk of evicting a parent transaction that could lead to turning, certain mempool is designed to restrict eviction victims to child transactions; that is, only the transaction of maximal nonce under a given sender can be evicted.
-For instance, given an arriving transaction $tx$, the OpenEthereum mempool finds an existing child transaction $tx^{\prime}$ with a lower GasPrice than $tx$ and evicts it to admit $tx$.
+For instance, given an arriving transaction $tx$, the OpenEthereum mempool finds an existing child transaction $tx^{\prime}$ with a lower fee than $tx$ and evicts it to admit $tx$.
 In Exploit $XT_{9}$, the attacker observes empty slots in the mempool (e.g., created by arriving blocks) and sends the same number of transactions to occupy them.
-The transactions are sent from one account where the child transaction has a higher GasPrice $f_{1}$ than normal transactions, while all its parent transactions have minimal fees.
-A normal transaction that arrives subsequently is declined because $f_{1}$ is higher than a normal transaction.
-Because each of its parent transactions has a low GasPrice, the mempool is locked at a low cost.
-6.3 Evaluation of Found Turning Exploits
-This and next subsections show the evaluation of turning attacks and locking attacks, and due to the space limit, the evaluation of other ADAMS attacks is deferred to Appendix D.
+The transactions are sent from one account where the child transaction has a higher fee $f_{1}$ than normal transactions, while all its parent transactions have minimal fees.
+A normal transaction that arrives subsequently is declined because $f_{1}$ is higher than the fee of a normal transaction.
+Because each of its parent transactions has a low fee, the mempool is locked at a low cost.
+7 Attack Evaluation
+7-A Evaluation of Found Eviction Exploits on Testnet
 We design experiments to measure the success of ADAMS attacks in real Ethereum networks.
-6.3.1 Evaluation in Testnet
+7-A1 Evaluation in Testnet
 Experimental settings: We set up our experiment platform by connecting the attacker node to the Goerli testnet.
@@
 To monitor transaction propagation, we launch an independent supernode (called measurement node) in the same testnet.
-The measurement node runs a reconfigured Geth client where the limit of peers/neighbors is removed and can connect to as many neighbors as possible.
+The measurement node runs a reconfigured Geth client where the limit of peers/neighbors is removed and the measurement node can connect to as many neighbors as possible, hence a super node.
 The measurement node is not (directly) connected to the attacker node.
@@
 Notice that our attacks require no discovery of critical nodes as needed in [1].
-6.4 Evaluation of Found Locking Exploits
-6.4.1 Exploit Extension
+7-B Evaluation of Found Locking Exploits on Reth
+Given Exploit $XT_{8}$ found on Reth, we manually extended it to two actual exploits on unmodified Reth’s mempool, $XT_{8a}$, and $XT_{8b}$.
+7-B1 Exploit Extension
 […]
-Actual exploit $XT_{8a}$: In $XT_{8}$, the attack cost increases with the block $BasePrice$.
-We manually propose a method to decrease the block $BasePrice$ in Ethereum by mounting an eviction attack on the previous block.
-Specifically, in Ethereum (after EIP1559), given a recently produced block $i$, the block $BasePrice$ of block $i+1$ is calculated dynamically as follows:
+Actual exploit $XT_{8a}$: In $XT_{8}$, the attack cost increases with the block base fee.
+We manually propose a method to decrease the block base fee in Ethereum by mounting an eviction attack on the previous block.
+Specifically, in Ethereum (after EIP1559), given a recently produced block $i$, the base block fee of block $i+1$ is calculated dynamically as follows:
 $BasePrice(i+1)$ $=$ $BasePrice(i)*[\frac{7}{8}+\frac{1}{4}*\frac{\textit{GasUsed}(i)}{BlockLimit}]$ (10)
-Therefore, if an eviction attack can persistently lower the Gas utilization of recent blocks, the current $BasePrice$ can be reduced, which decreases the costs of locking attacks on the current block.
+Therefore, if an eviction attack can persistently lower the Gas utilization of recent blocks, the current block fees can be reduced, which decreases the costs of locking attacks on the current block.
 In Exploit $XT_{8a}$, we mount a series of eviction attacks first and then a locking attack against the Reth node.
 Specifically, suppose in a network of a Reth node and a block validator node, the $XT_{8a}$ attacker first keeps sending eviction attacks directly to the validator node for several consecutive blocks.
-When observing the block $BasePrice$ drop sufficiently, the attacker mounts the regular locking attack $XT_{8}$ against the Reth node.
-6.4.2 Attack Evaluation on Reth
+When observing the block base fees drop sufficiently, the attacker mounts the regular locking attack $XT_{8}$ against the Reth node.
+7-B2 Attack Evaluation on Reth
 Experimental settings: We set up an experiment platform for evaluating locking attacks, on which an attacker node is connected to a victim non-validator node, which is also connected to a workload generator node and a victim validator node.
@@
 Figure 4: Experimental setup for locking attacks on Reth
-Figure 5: Evaluation of locking attack $XT_{8}$ on Reth
-Evaluation of $XT_{8}$: To evaluate $XT_{8}$, we set up the workload generator that sends normal transactions to the validator Geth node.
+Figure 5: Evaluation of locking attack $XT_{8a}$ on Reth
+Evaluation of $XT_{8a}$: To evaluate $XT_{8a}$, we set up the workload generator that sends normal transactions to the validator Geth node.
 The non-validator Reth node has not yet joined the network.
-We let the attacker node first send eviction attacks ($XT_{6}$) directly to the victim validator for $35$ consecutive blocks until it observes the block $BasePrice$ drops sufficiently to $1$ Gwei, which occurs at the $45$-th in our experiment.
+We let the attacker node first send eviction attacks ($XT_{6}$) directly to the victim validator for $35$ consecutive blocks until it observes the block base fees drop sufficiently to $1$ Gwei, which occurs at the $45$-th in our experiment.
 The Reth node then joins the network with an empty mempool.
@@
 Then, after three blocks, the normal transaction fees quickly drop to near-zero Ether, showing the success of locking attacks.
-Interestingly, unlike the eviction attacks that cause low Gas utilization per block, the locking attacks just reduce the total Ether fees without reducing Gas utilization (i.e., the blocks produced under locking attacks from height $45$ to $60$ use the Gas of almost $100\%$ block limit).
-Due to the high Gas utilization, block $BasePrice$ keeps increasing during the locking attack.
-When reaching the $60$-th block, the $BasePrice$ is higher than the maximal transaction price tolerable by the attacker.
+Interestingly, unlike the eviction attacks that cause low Gas utilization per block, the locking attacks just reduce the total Ether fees without reducing Gas utilization (i.e., the blocks produced under locking attacks from height $45$ to $60$ use the Gas of almost $100\%$ block limit as shown in Figure 6).
+Due to the high Gas utilization, block base fees keep increasing during the locking attack.
+When reaching the $60$-th block, the base fee is higher than the maximal transaction price tolerable by the attacker.
 In our experiment, the locking attacker stops at $60$-th block, and the normal transaction fees immediately recover to the “normal” level.
-In practice, the attacker can repeat sending eviction attacks to reduce the $BasePrice$ before mounting the locking attack again.
-7 Evaluation of mpfuzz
-7.1 Performance
-Preliminary of GoFuzz: Upon each fuzzing iteration, GoFuzz [16] generates a bit-string of variable length, feeds it to our fuzzer code, and receives from our fuzzer a binary value of feedback, be either positive or negative.
-The GoFuzz further checks if this run increases the code coverage.
-If the feedback is positive and code coverage increases, GoFuzz adds the current bit-string to the seeds before running into the next iteration.
-Baseline B1: Stateless fuzzer.
-We implement a stateless fuzzer in the GoFuzz framework described above.
-1) Given a bit string generated by GoFuzz, our fuzzer code parses it into a sequence of transactions; the length of the sequence depends on the length of the bit string.
-It skips a bitstring that is too short.
-2) The fuzzer sets up and initializes a Geth txpool instance with $m$ normal transactions, each sent from a distinct account, of nonce $1$, of value $1$ Ether (lower than the account balance), and of GasPrice $3$ wei.
-3) The fuzzer sends the transaction sequence to the initialized mempool for execution.
-4) It checks the mempool state after execution: If the test oracle is met, it emits the current transaction sequence as an exploit.
-Baseline B2: Concrete-state-coverage fuzzer.
-We build the second baseline, a mempool fuzzer that takes concrete state coverage as feedback.
-Specifically, in each iteration, the fuzzer appends a new transaction to the current transaction sequence.
-Given the $m$-slot mempool, the fuzzer tries $m$ values for senders, nonces, GasPrice, and Ether amount.
-After sending the current transaction sequence to the mempool, it sorts the transactions in the mempool by senders and nonces.
-It then hashes the sorted transactions in the mempool and checks the presence of the hash digest in the seeds.
-If no hash exists, the current transaction sequence would increase the concrete-state coverage and is inserted to the corpus.
-The corpus is a FIFO queue where the seed inserted earlier has high priority to be de-queued.
-Baseline B3: Invalid-tx energy fuzzer.
-Baseline B3 is similar to B2 except that it uses as the energy the number of invalid transactions on a mempool state.
-The seed with more invalid transactions in the mempool has higher energy or priority to be selected for the next round of fuzzing.
-Baseline B4: feedback of no promising states: We design Baseline B4 for an ablation study on the effectiveness of using promising states as feedback in improving fuzzing performance.
-B4 is the same with mpfuzz except that it removes the state promising-ness from its feedback.
-Specifically, the feedback formula in B4 includes only state coverage ($st\_coverage$) and excludes state promisingness ($st\_promising$ as in Equation 7).
-We implement Baseline B1 in Go/GoFuzz and implement B2, B3 and B4 in Python.
-We use a test oracle that can detect DETER attacks, that is, the number of invalid transactions in a mempool state equal $m-1$.
-Experimental settings of fuzzing: We run the four baselines and mpfuzz against a MUT in a Geth client reconfigured under two settings: A small setting where the mempool is resized to $6$ slots $m=6$ and all fuzzers run for two hours (if test oracle is not triggered), and a medium setting where $m=16$ and fuzzers run for $16$ hours.
-The fuzzing experiments are run on a local machine with an Intel i7-7700k CPU of $4$ cores and $64$ GB RAM.
-Table 4: Fuzzing Geth $v1.11.3$’s mempool (in minutes) by different approaches to detect Exploit $XT_{3}$.
-Settings B1 B2 B3 B4 mpfuzz
-$6$slot-$2$h Timeout $54$ $8$ $1.22$ $0.03$
-$16$slot-$16$h Timeout Timeout $447$ Timeout $0.06$
-Results: Table 4 reports the results under the small and medium MUT settings.
-For the small MUT, B1 cannot find any DETER attack in two hours.
-B2, B3 and B4 can find $XT_{3}$ in 54 minutes, $8$ minutes and $1.22$ minutes respectively.
-By contrast, mpfuzz finds the same exploit in $0.03$ minutes.
-For the medium setting, with $16$ slots, baselines B1, B2, and B4 cannot find any DETER exploit in $16$ hours, and B3 can find Exploit $XT_{3}$ in $7.4$ hours.
-By contrast, mpfuzz finds the same exploit in under a minute.
-The evaluation on different clients is in Appendix C.
-7.2 True/False Positive Rates
-This section evaluates the true/false positives of the short exploits found by mpfuzz.
-Recall that mpfuzz runs against a small mempool (MUT) and may introduce false positives, that is, the exploits that satisfy bug oracles in MUT but do not in the actual larger mempool.
-We formally define true and false positives as follows:
-Definition 7.1 (TP/FP exploits)
-Consider a short exploit found by mpfuzz on a MUT, $\langle{}st_{0},dc_{0}\rangle{},ops\Rightarrow{}\langle{}st_{n},dc_{n}\rangle{}$, and a (manually) extended exploit on an actual mempool, $\langle{}st_{0}^{\prime},dc_{0}^{\prime},ops^{\prime}\Rightarrow{}\langle{}st_{n}^{\prime},dc_{n}^{\prime}\rangle{}$.
-The pair of exploits constitute a true positive eviction attack w.r.t.
-$\epsilon$ ($\lambda$), if.f.
-$st_{0}\cap{}st_{n}=\varnothing\land{}asym_{E}(st_{0},ops)<\epsilon\land{}\ st_{0}^{\prime}\cap{}st_{n}^{\prime}=\varnothing\land{}asym_{E}(st_{0}^{\prime},ops^{\prime})<\epsilon$
-The pair of exploits constitute a false positive eviction attack w.r.t.
-$\epsilon$ ($\lambda$), if.f.
-$st_{0}\cap{}st_{n}=\varnothing\land{}asym_{E}(st_{0},ops)<\epsilon\land{}\ (st_{0}^{\prime}\cap{}st_{n}^{\prime}\neq{}\varnothing\lor{}asym_{E}(st_{0}^{\prime},ops^{\prime})>\epsilon)$.
-True/false positives on locking attacks can be similarly defined from bug-oracle definitions 4.2 and 4.3.
-We evaluate the false positives of the exploits found by mpfuzz under varying $\epsilon$ and $\lambda$.
-Specifically, we design a two-step experiment: First, given a MUT, we run mpfuzz with varying $\epsilon$ (and $\lambda$).
-Increasing $\epsilon$ allows mpfuzz to find more short exploits; for each distinct short exploit, we manually extend it to an actual exploit.
-Second, we evaluate the extended exploits on an actual mempool (in the same local experimental setting as described in § 9).
-Table 5: True/false positives (TP/FP) of eviction and locking attacks discovered with varying $\epsilon$ and $\lambda$. ✓means satisfied.
-Exploit $\epsilon$ $m=16$ (MUT) Default $m$ TP?
-Eq. 1 $asym_{E}$ Eq. 1 $asym_{E}^{\prime}$
-Geth-$XT_{1}$ $10^{-4}$ 00 00
-Geth-$XT_{3}$09.09.09.09 $0.083$ $0.0002$
-Geth-$XT_{6}$125.125.125.125 $0.125$ $0.0003$
-Geth-$XT_{2}$ $0.2$ $0.167$ $0.0698$
-Geth-$XT_{4}$ $0.23$ $0.208$ $0.0768$
-Geth-$XT_{5}$ $0.23$ $0.208$ $0.0768$
-Nethermind-$XT_{1}$ $10^{-4}$ 00 00
-Nethermind-$XT_{4}$ $0.11$ $0.104$ $0.0008$
-Nethermind-$XT_{7}$ $0.36$ $0.355$ $0.0012$
-Besu-$XT_{2}$ $0.2$167.167.167.167 $0.0754$
-Besu-$XT_{4}$ $0.23$208.208.208.208 $0.0083$
-Erigon-$XT_{4}$ $0.23$ $0.208$ $0.0894$
-Exploit $\lambda$ $m=16$ (MUT) Default $m$ TP?
-Eq. 4 $asym_{D}$ Eq. 4 $asym_{D}^{\prime}$
-Reth-$XT_{8}$ $0.34$ $0.34$ $0.015$
-OpenEthereum-$XT_{9}$ $0.46$ $0.46$ $0.0439$
-Table 5 presents the details of exploits discovered by mpfuzz.
-For instance, when setting $epsilon$ at $0.0001$ to find eviction attacks on Geth, mpfuzz discovered one short exploit $XT_{1}$ with $asym_{E}=0$ on MUT, which can be extended to an actual exploit on Geth mempool (of $m=6144$) with $asym_{E}^{\prime}=0<0.0001=\epsilon$.
-This makes it a true positive (marked by ✓in the table).
-Increasing $\epsilon$ to $0.23$ leads to discovering all six exploits $XT_{1-6}$ on Geth.
-All these exploits are true positives.
-The table also includes the results of eviction exploits found on other clients and locking exploits.
-Overall, the true-positive rate of mpfuzz remains at $100\%$ for $\epsilon\leq{}0.36$ and for $\lambda\leq{}0.46$ across Ethereum clients.
-8 Discussions
-Root causes of the exploits: We attribute all found exploits $XT_{1-9}$ to four distinct causes: 1) the presence of certain admission patterns (including $XT_{1}/XT_{2}/XT_{3}/XT_{4}/XT_{6}/XT_{7}$).
-For instance, $XT_{4}$ allows transaction replacement to cause latent overdrafts in place of valid transactions.
-2) the absence of certain admission patterns to cause mempool locking (including $XT_{9}$ and $XT_{8}$).
-For instance, the cause of $XT_{8}$ on Reth is that Reth mempool is a FIFO queue, and it disallows transaction eviction of any kind, which is risky.
-3) Inconsistency across multiple admission patterns, notably $XT_{7}$.
-In $XT_{7}$, the cause is that a mempool allows an eviction and its reversed eviction at the same time.
-4) Evadable conditions to context-sensitive admission patterns, including $XT_{3}$ and $XT_{6}$.
-Both exploits work on Geth, where the mempool triggers the limit of transactions per send under an evadable condition (when the mempool stores more than $py_{3}^{\prime}=5120$ pending transactions).
-Mitigation: We propose schemes to mitigate all newly discovered exploits comprehensively.
-In principle, given an exploit under a known cause, our mitigation design is to negate the cause.
-For instance, to mitigate $XT_{4}$, we propose declining transaction replacements that cause latent overdrafts.
-We have implemented this mitigation strategy on Geth, which was merged into the release of Geth $v1.11.4$ [8] .
-Similarly, to mitigate $XT_{7}$, we propose ensuring the consistency between eviction events and reversed events (e.g., if admitting $tx_{1}$ by evicting $tx_{2}$ is allowed, then admitting $tx_{2}$ by evicting $tx_{1}$ is not allowed).
-$XT_{1}/XT_{2}$ are mitigated by disallowing eviction of valid transactions by future/latent-overdraft transactions.
-$XT_{3}/XT_{6}$ are mitigated by removing the triggering condition of transaction sender limit (i.e., making $py_{3}^{\prime}=0$).
-$XT_{8}$ can be mitigated by re-enabling transaction evictions in certain cases, such as price-based transaction eviction (i.e., a pending transaction arriving later and of higher price can evict another pending transaction of lower price and admitted earlier).
-If an exploit relies on multiple causes (e.g., $XT_{3}$ on Causes 1 and 4), negating one cause suffices to mitigate the exploit.
-There are more sophisticated cases where one exploit relies on the presence of an admission pattern, and another exploit relies on the absence of the pattern.
-For instance, $XT_{5}$ relies on the presence of an admission policy named $SAP_{5}$ that turns child transactions into future ones (as in Geth), and the success of $XT_{9}$ relies on the absence of the same policy, that is, $\neg{}SAP_{5}$ (as in OpenEthereum).
-To mitigate both $XT_{5}$ and $XT_{9}$, we propose a defense strategy in which the admission decision regarding $SAP_{5}$ is non-deterministic or randomized so that $SAP$ is not always on or off, making it hard for either $XT_{5}$ or $XT_{9}$ to always succeed.
-[…]
-Ethical concerns: When evaluating attacks, we only mounted attacks on the testnet and did so with minimal impacts on the tested network.
-For instance, our attack lasts a short period of time, say no more than $4$ blocks produced.
-We did not test our attack on the Ethereum mainnet.
-We also mask part of the block numbers in the screenshots (e.g., Figure 3) to prevent detailed attack inspection and reproduction.
-When reporting bugs, we disclose to the developers the mitigation design tradeoff and the risk of fixing one attack by enabling another attack (described above).
-To prevent introducing new bugs, we did not suggest fixes against turning-based locking ($XT_{5}$, $XT_{6}$) and locking ($XT_{8}$).
-9 Conclusion and Future Works
-Conclusion: This paper presents mpfuzz, the first mempool fuzzer to find asymmetric DoS bugs by exploring symbolized mempool states and optimistically estimating the promisingness of an intermediate state in reaching bug oracles.
-Running mpfuzz on popular Ethereum clients discovers new mempool-DoS vulnerabilities, which exhibit various sophisticated patterns, including stealthy mempool eviction and mempool locking.
-Limitations and future works: The exploit generation in this work is not fully automated.
-Manual tasks include mempool reduction, mpfuzz setup (configuring $\epsilon$ and symbols), exploit extension, etc.
-Automating these tasks is the future work.
-mpfuzz’s bug oracles neither captures complete dependencies among concrete transactions nor guarantees completeness in finding vulnerabilities.
-Certified mempool security with completeness is also an open problem.
-mpfuzz targets a victim mempool of limited size: the vulnerabilities found in this work are related to transaction eviction, which does not occur in a mempool of infinite capacity.
-Thus, the mpfuzz workflow cannot find exploits in the mempool of infinite capacity.
-It is an open problem whether a mempool of large or infinite capacity has DoS vulnerabilities.
-10 Acknowledgments
-The authors thank USENIX Security reviewers and Kangjie Lu for the discussion in the early stage of this work.
-Yibo Wang, Wanning Ding, Zhihua Yang, and Yuzhe Tang were partially supported by two Ethereum Foundation academic grants and NSF grants CNS-2139801, CNS-1815814, DGE-2104532.
-Kai Li is partially supported by NSF grant CNS-2347486 and one Ethereum academic grant.
-Appendix A Observations Motivating mpfuzz Design
-(In)feasibility of code-coverage-only stateless fuzzing: To fuzz a mempool, a baseline design is to use a code-coverage-based fuzzer.
-In each iteration, the fuzzer generates a bit string, parses it into a sequence of transactions and sends them to the tested mempool for execution.
-Upon each end state, it checks whether the test oracle defined above is satisfied.
-This design takes code coverage as the only feedback and is stateless because the feedback does not consider the end state.
-The stateless design would be ineffective in fuzzing mempool or finding ADAMS exploits.
-For example, consider fuzzing a three-slot mempool to find a DETER-X exploit consisting of three future transactions [28].
-Suppose the current mempool input is a sequence of one future transaction.
-After mutation, it may try an input of two future transactions, which, however, does not increase code coverage (the second future transaction is executed by the same code path in a mempool as the first future transaction).
-Thus, the fuzzer discards the two-future-transaction input and misses finding the three-future-transaction exploit.
-We validate the ineffectiveness of stateless fuzzer by experiments in § 7.
-A more promising design is stateful mempool fuzzing, in which the mempool state is included in the feedback to guide next-iteration fuzzing.
-In the previous example of finding a three-future-transaction exploit (DETER-X), sending two future transactions leads to a different mempool state than sending one.
-A simple state-coverage-based fuzzer would view this as positive feedback and further explore this direction toward finding the three-future-transaction exploit.
-(In)feasibility of existing consensus fuzzers: Existing consensus fuzzers [25, 29, 36] cannot detect mempool DoS or ADAMS bugs.
-Specifically, Loki’s test oracle is detecting system crashes, while a successful ADAMS attack does not necessarily trigger system crashes.
-Using Loki to detect ADAMS would cause many false negatives.
-Fluffy detects the unsynchronizable difference of post-consensus states as vulnerability.
-However, the mempools in two clients, say Geth and Besu under benign transactions, would be different.
-In other words, post-consensus state differences across clients could indicate insecurity, but the difference in pre-consensus states across clients, like mempool, does not mean insecurity.
-Using Fluffly to detect ADAMS could cause false positives.
-Tyr’s test oracle detects property violations on post-consensus states.
-Particularly, while Tyr models state liveness, their definition “valid transactions should be executed eventually” is, unfortunately, over-simplifying and fails to model the legitimate pre-consensus cases in real Ethereum clients; for instance, valid transactions of low price can be dropped by Ethereum mempools and will not be eventually executed.
-Besides, they model double-spending transactions as the only type of invalid transaction.
-In Ethereum, double-spending transactions are of the same nonce.
-And there are many more sophisticated invalid transactions in Ethereum that Tyr does not model, including future transactions, latent overdrafts, etc.
-We analyze the case of existing consensus fuzzers in detecting known DETER attacks [28].
-Loki cannot detect DETER because DETER does not cause system crashes.
-Fluffy does not model future transactions and can not find the DETER attacks that rely on future transactions.
-Similarly, Tyr that does not model sophisticated invalid transactions would miss detecting DETER attacks.
-A.1 Further Rationale of Symbolization
-We now describe the design rationale of our symbolization technique.
-  • The first observation is that real-world mempools admit transactions based on the symbolic value, not concrete value, of nonce and Ether amount.
-Our design reflects this intuition.
-For instance, symbols of valid transactions, e.g., $\mathcal{P},\mathcal{C},\mathcal{N}$, are instantiated by mpfuzz to a minimal amount $1$ Ether.
-As long as the transferred value does not exceed the sender balance, transactions remain valid (i.e., no overdraft).
-Admission decisions remain the same for two valid transactions despite their difference in nonces or values.
-    Likewise, Symbol $\mathcal{F}$ is instantiated by mpfuzz to fixed nonce $n+1$: As long as the nonce is non-consecutive, no matter what specific value it takes (e.g., $3$ or $m+1$), the mempool would deem it as a future transaction and makes the same admission decision.
-    As a side note, the nonce of Symbol $\mathcal{F}$ is fixed at a large nonce $n+1$ (we denote it by far future transactions) instead of a small nonce, say $5$ (we denote it by near future transactions), because near future transactions’ nonce could be reconnected as consecutive and lead to duplicated states.
-  • The second observation is that restricting eviction/replacement victims to only parent transactions can facilitate finding exploits quickly.
-Specifically, while any transactions can be evicted/replaced, mpfuzz evicts/replaces only parent transactions; such a strategy can cause maximal damage (i.e., the most child transactions turned) while saving the search space.
-Our symbol design reflects this idea.
-For instance, Symbol $\mathcal{R}$ covers any transaction of the same sender and nonce as a transaction in state $st$, but it is instantiated by mpfuzz to only the transaction of nonce $1$ (i.e., replacing a parent transaction).
-Appendix B Case Study: How mpfuzz Finds Exploits
-We describe how mpfuzz finds an exploit by presenting a case study on finding $XT_{6}$ in the latest Geth $\geq{}v1.11.4$.
-Mempool reduction: Recall that a Geth mempool has a capacity of $m^{\prime}=6144$ slots, and its transaction-admission policies are characterized by three essential parameters: admitting up to $py_{1}^{\prime}=1024$ future transactions and limiting up to $py_{2}^{\prime}=16$ pending transactions from any senders when more than $py_{3}^{\prime}=5120$ pending transactions are residing in the mempool.
-We set up the MUT to run the same codebase or the same admission policy but with smaller parameters: $m=3,py_{1}=1,py_{2}=2,py_{3}=2$ What follows is a description of how mpfuzz find a short exploit on this MUT.
-Fuzzing: How mpfuzz automatically finds exploits: Initially, the mempool is filled with $m=3$ normal transactions.
-That is, the initial symbolized state is $\mathcal{NNN}$.
-The seed corpus $sdb$ initially contains an empty string.
-mpfuzz retrieves the empty string and appends to it with different symbols $\mathcal{O}_{0}$, $\mathcal{C}_{0}$, $\mathcal{P}_{0}$.
-Because of the initial state $\mathcal{NNN}$, only Symbol $\mathcal{P}_{0}$ is feasible.
-It generates the mutated input $\mathcal{P}$ and instantiates it to a parent transaction of a higher GasPrice than normal transactions (as described in § 5).
-Sending the input to the mempool gets the transaction admitted, leading to transitioned state $st_{1}=\mathcal{NNP}$.
-This is a new state that is not in the corpus, and it evicts more normal transactions $\mathcal{N}$ than the previous state $st_{0}$; the input produces positive feedback, and the associated input-state pair $\langle{}P,st_{1}=\mathcal{NNP}\rangle{}$ is added to the corpus.
-Figure 6: Snapshot of the mpfuzz state-search tree when finding Exploit $XT_{6}$ on Geth $v1.11.4$.
-Next, mpfuzz retrieves from $sdb$ a seed by high energy.
-According to Table 2, the energy of state $\mathcal{NNN}$ is $\frac{1}{3*3}*0$, and the energy of state $\mathcal{NNP}$ is $\frac{1}{3*2+4}*1=1/10>0$.
-Hence, seed $\mathcal{NNP}$ is selected.
-mpfuzz tries input mutation and appends to the selected input $\mathcal{P}$ one of four new symbols, that is, $\mathcal{O}_{1}$, $\mathcal{C}_{1}$, $\mathcal{P}_{0}$ or $\mathcal{P}_{1}$.
-On state $\mathcal{NNP}$, 1) mutation transaction instantiated from symbol $\mathcal{O}_{1}$ is declined admission.
-2) Mutation transaction from $\mathcal{C}_{1}$ is admitted, transitioning the state to $\mathcal{NPC}$, which produces positive feedback and is added to $sdb$.
-3) Likewise, transaction $\mathcal{P}_{0}$ is admitted and produces state $\mathcal{NPP}$ of positive feedback; the mutated input is also added to $sdb$.
-4) Mutation $\mathcal{P}_{1}$ is admitted but produces an identical state with mutation $\mathcal{P}_{0}$; thus, the mutated input is not added to $sdb$.
-Now, there are four seeds in $sdb$: $\mathcal{NNN}(0)$, $\mathcal{NNP}(0)$, $\mathcal{NPC}(1/8)$, and $\mathcal{NPP}(1/12)$.
-In parentheses are their energy numbers.
-The seed of the highest energy $\mathcal{NPC}$ is selected ( 1 in Figure 6).
-mpfuzz then mutates $\mathcal{NPC}$ with $4$ possibilities, that is, $\mathcal{O}_{1}$, $\mathcal{C}_{1}$, $\mathcal{P}_{0}$, and $\mathcal{P}_{1}$, which produce two end states with positive-feedback, that is, $\mathcal{PPC}$ and $\mathcal{PCP}$.
-They are added to the $sdb$ with energy $\mathcal{PPC}(1/10)$ and $\mathcal{PCP}(1/10)$.
-Let’s say $\mathcal{PPC}(1/10)$ is selected ( 2).
-mpfuzz mutates input $\mathcal{P}\mathcal{C}_{1}\mathcal{P}_{0}$ with $7$ mutation transactions, that is, $\mathcal{O}_{1}$, $\mathcal{O}_{2}$, $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, $\mathcal{P}_{0}$, $\mathcal{P}_{1}$, $\mathcal{P}_{2}$, which produces one end state with positive feedback, that is, mutation $\mathcal{C}_{1}$ is admitted and transitions state to $\mathcal{FPC}$.
-After that, $\mathcal{FPC}$ is the one with the highest energy and is chosen for the next-round fuzzing ( 4).
-mpfuzz mutates input $\mathcal{P}\mathcal{C}_{1}\mathcal{P}_{0}\mathcal{C}_{1}$ with $4$ mutations and produces $2$ state transitions with positive feedback.
-Upon state $\mathcal{FPC}$, mutation $\mathcal{C}_{1}$ is admitted and leads to state $\mathcal{FEE}$, which satisfies the bug oracle of eviction attacks under $\epsilon=0$.
-The algorithm then emits the found short exploit: $\langle{}st_{0}=\mathcal{NNN},dc_{0}=\varnothing\rangle{},ops=\mathcal{P}\mathcal{C}_{1}\mathcal{P}_{0}\mathcal{C}_{1}\mathcal{C}_{1}$ (recall Definition 4.1).
-The snapshot of the state-search tree is depicted in Figure 6.
-Exploit extension: Given the short exploit automatically found on MUT (with $m=3,py_{1}=1,py_{2}=2,py_{3}=2$), the next step is to extend it to a longer exploit functional on the actual Geth mempool (with $m^{\prime}=6144,py_{1}^{\prime}=1024,py_{2}^{\prime}=16,py_{3}^{\prime}=5120$).
-Exploit extension requires manual efforts: After identifying exploit $\langle{}st_{0}=\mathcal{NNN},dc_{0}=\varnothing\rangle{},ops=\mathcal{P}\mathcal{C}_{1}\mathcal{P}_{0}\mathcal{C}_{1}\mathcal{C}_{1}$ is unique, we extend it to the longer exploit by ensuring the same admission event occurs on the actual mempool as on the smaller MUT.
-In this process, it tries the next transaction of the same sender with the previous one but with an incremented nonce.
-If it fails, it switches to the next sender.
-It also tries transaction fees/prices based on measuring the fees of actual normal transactions, which may not be fixed as $3$ in the mpfuzz setting.
-Appendix C Additional Eval. of mpfuzz Performance
-We further run our mpfuzz on different clients, including Geth $v1.10.11$, Geth $v1.11.4$, Erigon $v2.42.0$, and Nethermind $v1.18.0$.
-The tested mempool is configured at $16$ slots.
-We report the number of exploits found by mpfuzz in a $16$-hour period.
-Figure 8 presents the result that mpfuzz finds $4096$ short exploits in the first $8$ hours on Geth-$v1.10.11$.
-On Geth-$v1.11.4$, mpfuzz quickly finds one exploit (i.e., exploit $XT_{6}$ as described in § 6.1) within $2$ minutes and does not find anymore in the next $16$ hours.
-On Nethermind, it finds one exploit (exploit $XT_{1}$) within two minutes and finds the next one (exploit $XT_{4}$) near the end of $16$-th hour.
-On Erigon, it finds one exploit ($XT_{4}$) in the $16$-th hour.
-Figure 7: # exploits found
-Figure 8: # states covered
-Figure 8 shows the number of states that are explored by mpfuzz on a $16$-slot mempool in 16 hours.
-mpfuzz explores $749$ and $91428$ states on Geth $v1.10.11$ in $2$ minutes and $16$ hours respectively.
-On Geth $v1.11.4$, mpfuzz explores $372$ and $23286$ states in $2$ minutes and $16$ hours respectively.
-However, the performance of Erigon and Nethermind is much lower than that of Geth.
-mpfuzz explores $12$ and $1239$ states on Erigon in $2$ minutes and $16$ hours respectively.
-On Nethermind, mpfuzz explores $15$ and $1433$ states in $2$ minutes and $16$ hours, respectively.
-The reason mpfuzz is more performant on Geth is that we implemented an external API on Geth to initialize the mempool; in each iteration of fuzzing on Geth, mpfuzz calls the API to initialize the mempool.
-In contrast, on Nethermind and Erigon, mpfuzz restarts the client in each iteration, which is consuming.
-Appendix D Additional Attack Evaluation
-D.1 Experiment on a Single Victim Node
+In practice, the attacker can repeat sending eviction attacks to reduce the base fee before mounting the locking attack again.
+Figure 6: Gas usage of attacks $XT_{8a}$
+Figure 6 shows the Gas usage of attack transactions and normal transactions in $XT_{8a}$.
+During the locking attack from block height $45$-th to $60$-th, the attacker sends locking transactions to occupy all the empty slots.
+It makes the Gas usage of the attacker almost 100% block limit while the Gas usage of the normal transactions drops to around 00.
+When empty slots are created by arriving blocks, normal transactions compete with the attack transactions to admit into the empty slots.
+Thus, some normal transactions are included in blocks during the locking attack.
+We also manually designed another locking variant for Reth.
+Suppose the Reth node does not run block validation, and its mempool is initially synchronized with the mempool of a block validator.
+The attacker first evicts $10,000$ normal transactions from the validator’s mempool.
+Then, the Reth node’s mempool is permanently locked by the $10,000$ normal transactions evicted from the validator mempool.
+Because these 10,000 transactions will never be included in blocks, they would exist in the Reth node’s mempool forever, permanently locking it in its current state.
+Figure 7: Evaluation of locking attack $XT_{8b}$ on Reth
+Evaluation of $XT_{8b}$: To evaluate $XT_{8b}$, we initially set up the attacker node connected to the validator node running Geth, the workload generator node connected to the victim non-validator Reth node, and the Reth node connected to the validator.
+The system is run long enough to produce a certain number of blocks and the normal transactions are propagated to both validator and on-validator nodes’ mempools.
+We then mount the eviction attack ($XT_{6}$) directly to the validator Geth node that evicts the mempool transactions on the validator node and prevents them from being included in the blocks.
+By mounting eviction attacks between the $30$-th to $45$-th blocks in Figure 7, the Reth mempool is successfully locked afterward: From the figure, after the $45$-th block, the total fees of normal transactions included in blocks remain at zero Ether per block until the experiment ends on the $60$-th block.
+7-C Evaluation of Locking on Geth
+Figure 8: Experimental setup for locking attacks on Geth/OpenEthereum
+(a) Attack damage
+(b) Attack cost
+Figure 9: Mounting locking attacks on a Geth node in testnet.
+Design rational: Recall that OpenEthereum adopts an overly strict policy that determines eviction by only considering the fees of childless transactions.
+While this policy is intended to prevent turning attacks ($XT_{5}$), it is vulnerable to locking attacks.
+Because OpenEthereum has been deprecated from the operational Ethereum 2.0 networks, we re-implement its admission policy on Geth $v1.11.4$ and run the turning-hardened Geth on a victim under locking attacks.
+Settings: We set up an experimental platform for evaluating locking attacks on Goerli testnet.
+The victim node runs the turning-hardened Geth client (based on $v1.11.4$) described above.
+The attacker node runs an instrumented Geth client (based on $v1.11.4$) at the execution layer such that it can mount locking attacks.
+Specifically, the Geth instrumentation enables the attacker node to monitor the local mempool and send a crafted transaction to the victim node upon each new empty slot created (by block arrival).
+The network topology is depicted in Figure 8.
+In the experiment, we first run our three nodes for $350$ minutes and then turn on the locking attack on the attacker node.
+We collect the benign transactions received by the attacker node (the number of which is denoted by $N_{a}$), the crafted transactions the attacker sends (the number denoted by $N_{c}$), and the benign transactions received by the measurement node (the number denoted by $N_{m}$).
+We use Metric $\frac{N_{a}}{N_{m}}$ to report the (inverse) of the attack success rate – Under the same transactions received by the measurement node, the fewer transactions the attacker node receives from the victim node, the more successful the victim mempool is locked.
+Evaluation of locking attacks: We report in Figure 9a the locking attack’s success rate over the period of $350$ minutes.
+Each tick on the $X$ axis represents $1$ minute; that is, $N_{a}$/$N_{m}$ is the number of benign transactions received by the attacker/measurement node every $1$ minute.
+It is clear that the locking attack is highly successful – Right after the attack is launched at the $180$-th minute, the rate of normalized transactions received from the victim node drops essentially to zero.
+And it persists over the period where the attack lasts.
+We also calculate the attack cost by the total fees of the attacker’s transactions included in the blockchain.
+In the experiment, we wait three hours after the attack stops and consider all the blocks produced until then.
+Not a single attacker’s transaction sent in our experiment is included in the blockchain.
+This may be due to the low fee with which crafted transactions are sent.
+Understanding crafted transaction inclusion is nondeterministic and dependent on benign transactions, we estimate the worst-case attack cost, which is the total fees of crafted transactions sent by the attacker (i.e., when all the crafted transactions are included in the blockchain).
+Figure 9b reports the worst-case locking attack cost over time; on average, the attack, as successful as reported in Figure 9a, consumes $0.0004$ Ether per block.
+By comparison, it is $0.028$ Ether of transactions included per block, which is the minimal cost of the baseline transaction-spamming attack [2].
+7-D Evaluation of Found Exploits on Single Victim Node
 Figure 10: Experimental setup
@@
 There is no direct connection between the attack node and the workload node.
-The attacker node runs an instrumented Geth $v1.11.4$ client that can propagate invalid transactions.
-The victim node runs the tested client.
+The attacker node runs an instrumented Geth $v1.11.4$ client (denoted by Geth-a) that can propagate invalid transactions to its neighbors.
+The victim node runs the tested Ethereum client.
 The workload node runs a vanilla Geth $v1.11.4$ client.
@@
 The experiment platform is denoted in Figure 10.
-Among the three nodes, only the victim node is staked as a validator and would propose or produce blocks.
-In each experiment, we first run the above “attacked” setup.
+Among the three nodes, we stake Ether to the consensus-layer client on the victim node, so that only the victim node would propose or produce blocks.
+In each experiment, we first run the above “attacked” setup (i.e., with victim, attack and workload nodes).
 We then run a “regular” setup that excludes the attack node.
@@
 We compare the experiment results under the attacked setup and regular setup to show the success of the attack.
-In the “attacked” setup, we collect the blocks produced and, given a block, we report two metrics: 1) total fees of benign transactions included, 2) total fees of attack transactions included.
-We also re-run the workload and victim nodes with the “regular” setup.
-We collect the blocks produced, and, given a block, we report two metrics: 3) total fees of transactions included, and 4) the cost of a baseline spamming attack with 100% success rate.
-For the latter, given a block, we select the transaction of the highest GasPrice and report the GasPrice multiplied by the Ethereum block Gas limit.
-Table 6: Attack success rate and cost (Ether/block)
-Clients Exploit Success rate Cost Baseline
-Geth $v1.10.25$ $XT_{1}$ $99.80\%$ 00 $11.39$
-$XT_{2}$ $99.42\%$ $0.725$ $11.39$
+Evaluation of eviction/turning attacks on Single Node: We set up the experiment platform described in § 7-D.
+In each experiment, we drive benign transactions from the workload node.
+Note that the collected workload contains the timings of both benign transactions and produced blocks.
+On the $30$-th block, we start the attack.
+Recall that each attack is configured by delay $d$; the attack node observes the arrival of a produced block and waits for $d$ seconds before sending a round of crafted transactions.
+The attack phase lasts for $40$ blocks; after the $70$-th block, we stop the attack node from sending crafted transactions.
+We keep running workload and victim nodes for another $50$ blocks and stop the entire process at the $120$-th block.
+We collect the blocks produced and, given a block, we report two metrics: 1) total fees of benign transactions included, 2) total fees of attack transactions included.
+In each experiment, we also re-run the workload and victim nodes with the same setup.
+In this “no-attack”, we don’t run the attack node.
+We collect the blocks produced, and, given a block, we report two metrics: 3) total fees of transactions included (denoted as “Benign ops - no attack”), and 4) the cost of a baseline spamming attack with 100% success rate (denoted as “Spamming - analytical”).
+For the latter, given a block, we select the transaction of the highest price, and report the price multiplied by the Ethereum block Gas limit.
+(a) $XT_{6}$ w. $8$-sec. delay (single node)
+(b) $XT_{6}$ w. varying delay (single node)
+(c) $XT_{4}$ on Nethermind w. varying delay (single node)
+(d) $XT_{6}$ w. varying delay ($6$ nodes)
+(e) $XT_{4}$ on Geth w. varying delay
+                                                                                                                                    (single node)
+(f) $XT_{4}$ on Erigon w. varying delay
+                                                                                                                                   (single node)
+(g) $XT_{8a}$ on Reth w. varying delay
+                                                                                                                                    (single node)
+(h) $XT_{9}$ on openEthereum w. varying
+                                                                                                                                   delay (single node)
+Figure 11: Success rate and cost of turning-based eviction attack and locking attack.
+Figure 11a reports the metrics for $XT_{6}$ on a victim node running Geth $v1.11.4$.
+Before the attack is launched (i.e., before the $30$-th block), the two experiment runs (attacked and no attack) produce the blocks of the same benign transaction fees, i.e., around $2-3$ Ether per block.
+As soon as the attack starts from the $30$-th block, the benign transaction fees quickly drop to zero Ether except for some sporadic spikes (under $1$ Ether per block).
+The non-zero Ether cost is due to that $XT_{6}$ cannot lock the mempool.
+The adversarial transaction fees show up and remain close to zero Ether per block.
+After the attack stops on $70$-th block, the benign transactions under attack gradually recover to the level of transactions under no attack.
+We also conduct experiments with varying attack delays between 00 and $12$ seconds (note that the average time to produce a block in our collected trace is $14$ seconds).
+Figure 11c shows the results of Attack $XT_{4}$ on Nethermind.
+On shorter delay than $8$ seconds, the success rates are strictly $100\%$, and the attack cost remains at $0.025$ Ether per block.
+When the delay grows over $8$ seconds, success rates begin to drop, and attack costs increase.
+At the delay of $12$ seconds, the success rate is $80\%$, and the attack cost is $1.5$ Ether per block.
+The reason is that as the delay becomes large, the transaction sequence sent in an attack can be interrupted by block production, leading to adversarial transactions included in the blockchain and normal transactions un-evicted from the mempool.
+The results of Attack $XT_{4}$ on Geth as shown in Figure 11e are similar to those on Nethermind.
+The difference is the cost of $XT_{4}$ on Geth is higher than that of Nethermind.
+In the attack sequence of $XT_{4}$ on Geth, it needs $384$ attack senders which means $384$ attack transactions are included in blocks on shorter delay than $8$ seconds while only $1$ attack transaction is included in $XT_{4}$ on Nethermind.
+Figure 11b shows the success rates and costs of Exploit $XT_{6}$ on Geth $v1.11.4$.
+The results are similar to those in Figure 11c except that when the delay is short, the success rates of $XT_{6}$ are not $100\%$ because $XT_{6}$ cannot lock the attacked mempool and an attack sent too early risks admitting normal transactions sent after the attack.
+The results of Attack $XT_{4}$ on Erigon are shown in Figure 11f.
+The results are similar to those in Figure 11b as the $XT_{4}$ on Erigon also cannot lock the attacked mempool when the delay is short.
+Figure 11g shows the success rates and cost of $XT_{8}$ on Reth.
+When the delay is 00 second, the success rate is $99.04\%$, and the cost is $0.151$ Ether per block.
+The low-priced attack transactions sent with a 00 delay occupy all the empty slots and decline all the incoming normal transactions.
+When the delay grows, the success rate and attack cost begin to drop.
+The reason is as the delay increases, more empty slots in the mempool are occupied by normal transactions.
+As a result, more normal transactions and less attack transactions are included in blocks.
+The results of Attack $XT_{9}$ on OpenEthereum are shown in Figure 11h.
+The results are similar to those of $XT_{8}$ on Reth except when the delay is short, the cost of $XT_{9}$ on OpenEthereum is $0.177$ per block which is higher than that of Reth.
+The reason is that each attack sender in the $XT_{9}$ attack sequence sends a high-priced child transaction for declining the incoming normal transactions and the high-priced transactions are included in blocks.
+The evaluation results of all attacks on all clients are in Table IV.
+On the six Ethereum clients in the public transaction path, the attack success rates are all higher than $84.63\%$, and the attack costs are all lower than $1.172$ Ether per block, which is significantly lower than the baseline.
+On PBS clients and Ethereum-like clients, we also evaluate the attacks found by mpfuzz.
+Under the same experimental settings as § 7-D, the results show similar success rates and costs with the attacks on Geth (recall these clients are Geth forks).
+More specifically, Table IV shows that the success rates are higher than $92.60\%$, and attack costs are lower than $0.806$ Ether per block.
+TABLE IV: Attack success rate and cost
+Clients Exploit Success rate Cost (Ether/block) Baseline (Ether/block)
+Geth $XT_{1}$ $99.80\%$ 00 $11.39$
+$v1.10.25$ $XT_{2}$ $99.42\%$ $0.725$ $11.39$
 $XT_{3}$ $93.02\%$ $0.0021$ $11.39$
-Geth $v1.11.4$ $XT_{4}$ $99.42\%$ $0.806$ $11.39$
-$XT_{5}$ $92.65\%$ $0.806$ $11.39$
+Geth $XT_{4}$ $99.42\%$ $0.806$ $11.39$
+$v1.11.4$ $XT_{5}$ $92.65\%$ $0.806$ $11.39$
 $XT_{6}$ $94.74\%$ $0.0022$ $11.39$
 Erigon $v2.42.0$ $XT_{4}$ $92.53\%$ $1.172$ $17.7$
-Nethermind $v1.18.0$ $XT_{4}$ $99.60\%$ $0.0021$ $10.75$
-$XT_{7}$ $84.63\%$ $0.20$ $10.75$
-Besu $v22.7.4$ $XT_{2}$ $99.63\%$ $1.04$ $17.7$
-$XT_{4}$ $99.60\%$ $1.06$ $17.7$
+Nethermind $XT_{4}$ $99.60\%$ $0.0021$ $10.75$
+$v1.18.0$ $XT_{7}$ $84.63\%$ $0.20$ $10.75$
+Besu $XT_{2}$ $99.63\%$ $1.04$ $17.7$
+$v22.7.4$ $XT_{4}$ $99.60\%$ $1.06$ $17.7$
 […]
@@
 bloXroute builder-ws $XT_{6}$ $94.74\%$ $0.0022$ $11.39$
-go-opera $v1.1.3$ $XT_{2}$ $99.12\%$ $0.201$ $11.39$
-$XT_{3}$ $92.60\%$ $0.0021$ $11.39$
+go-opera $XT_{2}$ $99.12\%$ $0.201$ $11.39$
+$v1.1.3$ $XT_{3}$ $92.60\%$ $0.0021$ $11.39$
 $XT_{4}$ $99.13\%$ $0.221$ $11.39$
@@
 […]
-OpenEthereum $v3.3.5$ $XT_{4}$ $99.56\%$ $0.233$ $11.39$
-$XT_{9}$ $98.36\%$ $0.177$ $18.35$
-The evaluation results of all attacks on all clients are in Table 6.
-On the six Ethereum clients in the public transaction path, the attack success rates are all higher than $84.63\%$, and the attack costs are all lower than $1.172$ Ether per block, which is significantly lower than the baseline.
-We also evaluate the attacks found on PBS and Ethereum-like clients.
-Under the same experimental settings as § D.1, the results show similar success rates and costs with the attacks on Geth (recall these clients are Geth forks).
-More specifically, Table 6 shows that the success rates are higher than $92.60\%$, and attack costs are lower than $0.806$ Ether per block.
+OpenEthereum $XT_{4}$ $99.56\%$ $0.233$ $11.39$
+$v3.3.5$ $XT_{9}$ $98.36\%$ $0.177$ $18.35$
+7-E Evaluation of Found Exploits on Multi Nodes
+Evaluation settings: We extend our experiment platform to support a network of victim nodes, instead of a single victim node.
+Specifically, the attack node is connected to the first victim non-validator node $V_{1}$.
+Node $V_{1}$ is connected to the second victim non-validator node $V_{2}$, which is further connected to node $V_{3}$.
+The chain of victims continues until it reaches the $6$-th non-validator victim $V_{6}$, which is connected to the validator victim node $V_{0}$.
+The workload node is connected to the validator node $V_{0}$.
+We run the victim network among a set of geo-distributed cloud instances on the Internet.
+We place victim nodes $V_{1},V_{3},V_{5}$ in one geographic area in Amazon AWS cloud (a.k.a., one Availability Zone or AZ).
+We place the other victim nodes $V_{2},V_{4},V_{6},V_{0}$ in a different geographic area, so that each message sent between $V_{i}$ and $V_{i+1}$ has to travel across continents on the Internet.
+The experiment architecture is illustrated in Figure 10.
+Experiment results: On the multi-victim platform, we conduct similar experiments as in § 7-D.
+The results are presented in Figure 11d.
+Compared with the results on the single node, the attack cost increases faster and is much higher.
+For instance, when the delay is $10$/$12$ seconds, the attack cost is $8$/$8.5$ Ether per block, which is higher than the attack damage (i.e., the victim transaction fees).
+Under the delay, the attacks incur higher attack cost than damage and are no longer ADAMS.
+By contrast, from Figure 11b, the attack costs on the single node with $10$/$12$-second delays are $1$/$1.4$ Ether per block, both of which are below the attack damage.
+The cause of the higher attack cost is the longer time to propagate the transaction sequence of an ADAMS attack in the multi-node setting than in the single node; the longer time implies the higher chance the transaction sequence is interrupted by the production of the next block and that the transactions are included to the block without being turned invalid.
+8 Evaluation of mpfuzz
+8-A Ablation Study
+In this section, we describe $4$ baseline fuzzers to demonstrate the performance efficacy of the techniques proposed in mpfuzz as an ablation study.
+Baseline B1 - stateless: We implement a stateless fuzzer in the GoFuzz framework.
+1) Given a bit string generated by GoFuzz, the fuzzer code parses it into a sequence of transactions; the length of the sequence depends on the length of the bit string.
+It skips a bitstring that is too short.
+Specifically, when parsing the bitstring into transaction sequence, every 3 bits are parsed into one transaction, corresponding to 6 adversarial symbolized transactions proposed in mpfuzz.
+2) The fuzzer then sets up and initializes a Geth txpool instance with $m$ normal transactions generated from the symbolized transaction $\mathcal{N}$.
+3) The fuzzer sends the transaction sequence to the initialized mempool for execution.
+4) It checks the mempool state after execution: If the eviction-based test oracle (Equations 1,  2 and  3) is met, it emits the current transaction sequence as an exploit.
+5) As a stateless fuzzer, it cannot use the same feedback technique as that of mpfuzz.
+In contrast, it takes code coverage as feedback, specifically, if this run increases the code coverage in Geth client, the fuzzer adds the current bit-string to the seeds corpus.
+Baseline B2a - non-symbolized state: Compared to mpfuzz, the only difference is that the baseline B2a takes concrete state in the state coverage, while mpfuzz takes symbolized state in the state coverage.
+Specifically, after sending the current transaction sequence to the mempool, it sorts the transactions in the mempool by senders and nonces.
+It then hashes the sorted transactions in the mempool and checks the presence of the hash digest in the seeds.
+If no hash exists, the current transaction sequence would increase the concrete-state coverage and be inserted into the corpus.
+However, the energy of a state is determined the same way as mpfuzz.
+Specifically, given the concrete state that an input reaches, the energy is determined by its symbolized state.
+Baseline B2b - non-symbolized input: We implement the third baseline fuzzer, which is also similar to mpfuzz but without symbolized input mutation.
+In each iteration, the baseline fuzzer B2b appends to the current transaction sequence a new transaction.
+Given the m-slot mempool, the fuzzer tries $m$ values for senders, $m$ values for nonces, $m$ values for Gas price, and $m$ values for Ether amount.
+After sending to the mempool the current transaction sequence, it uses the same way to determine the feedback and energy as mpfuzz.
+Baseline B3 - non-promising feedback & energy We build the last baseline fuzzer, which is the same with mpfuzz except that it removes the state promising-ness from its feedback and energy mechanism proposed in mpfuzz from the seed selection.
+Specifically, the feedback formula in B3 includes only state coverage ($st\_coverage$) and excludes state promisingness ($st\_promising$ as in Equation 7).
+In addition, B3 uses as the energy the number of invalid transactions on a mempool state.
+The seed with more invalid transactions in the mempool has higher energy or priority to be selected for the next round of fuzzing.
+Experimental settings of fuzzing: We run the four baselines and mpfuzz against a MUT in a Geth client reconfigured under two settings: A small setting where the mempool is resized to $6$ slots $m=6$ and all fuzzers run for two hours (if test oracle is not triggered), and a medium setting where $m=16$ and fuzzers run for $16$ hours.
+The fuzzing experiments are run on a local machine with an Intel i7-7700k CPU of $4$ cores and $64$ GB RAM.
+Results: Table V reports the time used that the first exploit is found under the small and medium MUT settings.
+For the small MUT of $6$ slots, baseline B1 cannot find any attack in two hours, B2a, B2b and B3 can find the Exploit $XT_{3}$ in $0.10$, $34.61$ and $1.66$ minutes respectively.
+By contrast, mpfuzz finds the same exploit ($XT_{3}$) in $0.03$ minutes.
+For the medium setting, with $16$ slots, baselines B1, B2b, and B3 cannot find any exploit in $16$ hours, while B2a can find Exploit $XT_{3}$ in $0.26$ minutes.
+By contrast, mpfuzz finds the same exploit in $0.06$ minutes.
+The result that B1 has the poorest performance shows a stateless fuzzer would be ineffective in discovering ADAMS exploits.
+As B2a and B3 are more performant compared to B2b, it shows the symbolized input technique improves efficiency more compared to that of the promising feedback and symbolized state coverage.
+The comparison of results between B2a and B3 indicates that the state promisingness in feedback is more beneficial for performance than symbolized state coverage.
+TABLE V: Fuzzing Geth $v1.11.3$’s mempool (in minutes) by different approaches to detect Exploit $XT_{3}$. OT means overtime.
+Settings B1 B2a B2b B3 mpfuzz
+$6$slot-$2$h OT $0.10$ $34.61$ $1.66$ $0.03$
+$16$slot-$16$h OT $0.26$ OT OT $0.06$
+The other experiments, including true/false positive rates and additional performance evaluation evaluation on different clients can be found in Section 7.2 and Appendix C, available in the online supplemental material, and are identical to the original conference paper [37].
+9 Case Study: How mpfuzz Finds Exploits in MUT and Extends to a Large Mempool
+We describe how mpfuzz finds an exploit that evades the defense.
+We present a case study on finding $XT_{6}$ in the latest Geth $\geq{}v1.11.4$.
+Recall that a Geth mempool has a capacity of $m^{\prime}=6144$ slots, and its transaction-admission policies are characterized by three essential parameters: admitting up to $py_{1}^{\prime}=1024$ future transactions and limiting up to $py_{2}^{\prime}=16$ pending transactions from any senders when more than $py_{3}^{\prime}=5120$ pending transactions are residing in the mempool.
+For ease of description, we set up the MUT to run the same codebase or the same admission policy but with different, smaller parameters: $m=4,py_{1}=1,py_{2}=2,py_{3}=2$.
+We first describe how mpfuzz finds a short exploit of $XT_{6}$ on the MUT.
+We then describe how mpfuzz extends the short exploit into a longer version suitable for a medium-sized mempool.
+Fuzzing: How mpfuzz automatically finds exploits: Initially, the mempool is filled with $m=4$ normal transactions.
+That is, the initial symbolized state is $\mathcal{NNNN}$.
+The seed corpus $sdb$ initially contains an empty string.
+To simplify the description, we only show Symbols $\mathcal{C}$ and $\mathcal{P}$ in the input mutation for ease of illustration.
+mpfuzz retrieves the empty string and appends to it with different symbols $\mathcal{C}_{0}$ and $\mathcal{P}_{0}$.
+Because of the initial state $\mathcal{NNNN}$, only Symbol $\mathcal{P}_{0}$ is feasible.
+It generates the mutated input $\mathcal{P}$ and instantiates it to a parent transaction of a higher fee than normal transactions (as described in § 5).
+Sending the input to the mempool gets the transaction admitted, leading to transitioned state $st_{1}=\mathcal{NNNP}$.
+This is a new state that is not in the corpus, and it evicts more normal transactions $\mathcal{N}$ than the previous state $st_{0}$; the input produces positive feedback, and the associated input-state pair $\langle{}P,st_{1}=\mathcal{NNNP}\rangle{}$ is added to the corpus.
+Next, mpfuzz retrieves from $sdb$ a seed by high energy.
+According to Table II, the energy of state $\mathcal{NNNN}$ is $\frac{1}{3*4}*0$, and the energy of state $\mathcal{NNNP}$ is $\frac{1}{3*3+4}*1=1/13>0$.
+Hence, seed $\mathcal{NNNP}$ is selected.
+mpfuzz tries input mutation and appends to the selected input $\mathcal{P}$ one of three new symbols, that is, $\mathcal{C}_{1}$, $\mathcal{P}_{0}$ or $\mathcal{P}_{1}$.
+On state $\mathcal{NNNP}$, 1) Mutation transaction from $\mathcal{C}_{1}$ is admitted, transitioning the state to $\mathcal{NNPC}$, which produces positive feedback and is added to $sdb$.
+2) Likewise, transaction $\mathcal{P}_{0}$ is admitted and produces state $\mathcal{NNPP}$ of positive feedback; the mutated input is also added to $sdb$.
+3) Mutation $\mathcal{P}_{1}$ is admitted but produces an identical state with mutation $\mathcal{P}_{0}$; thus, the mutated input is not added to $sdb$.
+Figure 12: Snapshot of the mpfuzz state-search tree when finding Exploit $XT_{6}$ on Geth $v1.11.4$.
+Now, there are four seeds in $sdb$: $\mathcal{NNNN}(0)$, $\mathcal{NNNP}(0)$, $\mathcal{NNPC}(1/11)$, and $\mathcal{NNPP}(1/15)$.
+In parentheses are their energy numbers.
+The seed of the highest energy $\mathcal{NNPC}(1/11)$ is selected (i.e., Step 1 in Figure 12).
+mpfuzz then mutates $\mathcal{NNPC}$ with three possibilities, that is, $\mathcal{C}_{1}$, $\mathcal{P}_{0}$, and $\mathcal{P}_{1}$, which produce two end states with positive-feedback, that is, $\mathcal{NPPC}$ and $\mathcal{NPCP}$.
+They are added to the $sdb$ with energy $\mathcal{NPPC}(1/13)$ and $\mathcal{NPCP}(1/13)$.
+[…]
+[…]
+Figure 13: Exploit $XT_{6}$ extension on Geth $v1.11.4$ with a medium-size mempool.
+We then construct the attack transaction sequence on the larger-size mempool by aligning the patterns identified from the small exploit with corresponding admission events.
+Specifically, this entails verifying that the admission event pattern of a given event $f(st_{0},tx_{i})\Rightarrow{}st_{n}$ on the larger-size mempool is present within the $aep$.
+Initially, the mempool is filled with $m=8$ normal transactions and the symbolized state is $\mathcal{NNNNNNNN}$.
+There are two patterns in $aep$ whose initial state $st_{0}^{\prime}$ matches the mempool state, that is $\langle{}\mathcal{N},\mathcal{P},\mathcal{P}\rangle{}$ and $\langle{}\mathcal{N},\mathcal{C},\mathcal{C}\rangle{}$ ( 1a).
+Between these two patterns, only sending transaction $\mathcal{P}$ is feasible.
+Therefore, mpfuzz sends $\mathcal{P}$ upon the initial state which leads to transitioned state $\mathcal{NNNNNNNP}$.
+In this admission event, the submission of transaction $\mathcal{P}$ results in its admission in the mempool by evicting a $\mathcal{N}$ in the initial state.
+Thus, it aligns the pattern of $\langle{}\mathcal{N},\mathcal{P},\mathcal{P}\rangle{}$.
+This positive outcome prompts mpfuzz to proceed with the extension process accordingly.
+[…]
+[…]
+10 Discussions
+[…]
+Ethical concerns: When evaluating attacks, we only mounted attacks on the testnet and did so with minimal impacts on the tested network.
+For instance, our attack lasts a short period of time, say no more than $4$ blocks produced.
+We did not test our attack on the Ethereum mainnet.
+Additionally, we obscure the first few digits of the block number in the attack screenshot, i.e., Figure 3.
+When reporting bugs, we disclose to the developers the mitigation design tradeoff and the risk of fixing one attacks by enabling other attacks (described above).
+To prevent introducing new bugs, we did not suggest fixes against turning-based locking ($XT_{5}$, $XT_{6}$) and locking ($XT_{8}$).
+11 Conclusion and Future Works
+Conclusion: This paper presents mpfuzz, the first mempool fuzzer to find asymmetric DoS bugs by exploring symbolized mempool states and optimistically estimating the promisingness of an intermediate state in reaching bug oracles.
+Running mpfuzz on popular Ethereum clients discovers new mempool-DoS bugs, which exhibit various sophisticated patterns, including stealthy mempool eviction and mempool locking.
+Limitations and future works: The exploit generation in this work is not fully automated.
+Manual tasks include mempool reduction, mpfuzz setup (configuring $\epsilon$ and symbols), etc.
+Automating these tasks is the future work.
+mpfuzz’s bug oracles neither captures complete dependencies among concrete transactions nor guarantees completeness in finding vulnerabilities.
+Certified mempool security with completeness is also an open problem.
+mpfuzz targets a victim mempool of limited size: the vulnerabilities found in this work are related to transaction eviction, which does not occur in a mempool of infinite capacity.
+Thus, the mpfuzz workflow cannot find exploits in the mempool of infinite capacity.
+It is an open problem whether a mempool of large or infinite capacity has DoS bugs.
+12 Acknowledgments
+The authors thank the anonymous reviewers in USENIX security’24.
+The authors appreciate the discussion with Kangjie Lu in the early stage of this work.
+All authors but Kai Li are partially supported by two Ethereum Foundation academic grants and NSF grants CNS-2139801, CNS-1815814, DGE2104532.
+Kai Li is supported by NSF grant CNS2347486 and one Ethereum Academic Grant.